### v1

# SageAttention2 Technical Report: Accurate 4 Bit Attention for Plug-and-play Inference Acceleration

Abstract
Although quantization for linear layers has been widely used, its application to accelerate the attention process remains limited. SageAttention utilizes 8-bit matrix multiplication, 16-bit matrix multiplication with 16-bit accumulator, and precision-enhancing methods, implementing an accurate and 2x speedup kernel compared to FlashAttention2. To further enhance the efficiency of attention computation while maintaining precision, we propose SageAttention2, which utilizes significantly faster 4-bit matrix multiplication (Matmul) alongside additional precision-enhancing techniques. First, we propose to quantize matrixes $(Q,K)$ to INT4 in a warp-level granularity and quantize matrixes $(\widetilde{P},V)$ to FP8. Second, we propose a method to smooth $Q$ and $V$, enhancing the accuracy of attention with INT4 $QK$ and FP8 $PV$. Third, we analyze the quantization accuracy across timesteps and layers, then propose an adaptive quantization method to ensure the end-to-end metrics over various models. The operations per second (OPS) of SageAttention2 surpass FlashAttention2 and xformers by about 3x and 5x on RTX4090, respectively. Comprehensive experiments confirm that our approach incurs negligible end-to-end metrics loss across diverse models—including those for large language processing, image generation, and video generation. The codes are available at https://github.com/thu-ml/SageAttention.
footnotetext: * Equal Contribution.
Figure 1: The left figure shows the kernel speedup on RTX4090 GPU. The right figure shows the end-to-end inference speedup of generating the first token and performance metrics for the needle-in-a-haystack task (gkamradt, 2023) with a sequence length of 100K on Llama3.1 on L20 GPU.
1 Introduction
Due to the quadratic complexity of attention, efficient implementation becomes increasingly crucial as sequences lengthen in real-world applications (Jiang et al., 2024). Several strategies have been developed to mitigate computational demands of attention —such as (1) Linear attention methods (Wang et al., 2020; Choromanski et al., 2020; Yu et al., 2022; Katharopoulos et al., 2020) that reduce complexity to $O(N)$ and (2) Sparse attention methods (Liu et al., 2021; Chu et al., 2021; Li et al., ; Xiao et al., 2023b, 2024; Chen et al., 2023; Jiang et al., 2024; Venkataramanan et al., 2023; Gao et al., 2024; Fu et al., 2024) that selectively process parts of the context — these methods are only suitable for a limited range of models and tasks. Widely used attention methods optimize attention by exploiting hardware capacities to enhance the computation speed, such as FlashAttention (Dao et al., 2022), FlashAttention2 (Dao, 2023), FlashAttention3 (Shah et al., 2024), xformers (Lefaudeux et al., 2022), and SageAttention (Zhang et al., 2024). Those works do not omit the computation of attention over parts of the context and achieve impressive speed and precision performance across various models and tasks.
For the two matrix multiplication (Matmul) operations in attention: $QK^{\top}$ and $PV$, SageAttention accelerates them by quantizing the $QK^{\top}$ to INT8 and uses FP16 Matmul with FP16 accumulators for $PV$. Moreover, to keep the accuracy of attention, SageAttention proposes smoothing $K$ by eliminating its channel-wise outliers. SageAttention achieves 2 $\times$ and 2.7 $\times$ speedup than FlashAttention2 and xformers and is the first quantized attention that incurs negligible end-to-end metrics loss across language, image, and video generation models. However, SageAttention has two weaknesses. (W1) INT8 Matmul achieves only half the speed of INT4. (W2) FP16 Matmul with FP16 accumulators is only compatible with RTX4090 and RTX3090 GPUs. To leverage the faster INT4 tensor cores for $QK^{\top}$ and using a method that can generally accelerate $PV$, we propose SageAttention2, which quantizes $Q,K$ to INT4 and $P,V$ to FP8.
Challenges. Quantizing $Q,K$ to INT4 and $P,V$ to FP8 presents significant challenges. For example, when only per-tensor quantizing $Q,K$ to INT4, the text-to-video model CogvideoX will generate a completely blurry video (See Figure 2), and Llama2 only achieves a random-guessing-level accuracy of 25% on MMLU. After investigating deeply, we identified three primary challenges: (C1) The numerical range of INT4, which in quantization typically spans from $-7$ to $7$ and totals 15 numbers (Lin et al., 2024), leads to significant quantization errors when $Q$ and $K$ have some abnormal values. (C2) In specific layers and timesteps (for text-to-image/video) of some models, quantizing $Q$ and $K$ to INT4 and quantizing $P$ and $V$ to FP8 introduces noticeable errors in attention computation. These errors in worst-case layers/timesteps significantly impact the precision of the end-to-end output. (C3) We discover that the FP32 accumulator designed for FP8 matrix multiplication in the tensor core (mma.f32.f8.f8.f32) is actually FP22, specifically with 1 sign bit, 8 exponent bits, and 13 mantissa bits. This will lead to an accuracy loss of $PV$.
Figure 2: An example of quantizing Q, K to INT4 from CogvideoX.
Our approach. To address these challenges, we detailedly analyze the reasons and propose two methods. First, for the significant channel-wise outliers in matrices $Q$ and $K$, we adopt smoothing $K$ in SageAttention and propose an effective method to remove these outliers in $Q$. Specifically, we propose subtracting the average value of the channel dimension of $Q$, referred to as $\overrightarrow{Q}_{m}$. Then, we add $\overrightarrow{Q}_{m}K$ after the $QK$ Matmul to ensure the correctness of the attention computation. Second, we observe that specific layers and timesteps consistently present quantization challenges across different inputs. To maintain accuracy, we apply an adaptive mix precision method. Specifically, we employ 8-bit (INT8+FP8) attention for these problematic layers and timesteps and 4-bit (INT4+FP8) attention for the others. Third, to mitigate the accuracy loss associated with using the 22-bit accumulator for FP8 Matmul of $PV$, we propose a method to smooth $V$ to enhance the accuracy performance.
Performance. Importantly, we offer a high-performance implementation of SageAttention2 on RTX4090 and L20 GPUs. This implementation achieves a peak performance of 485 TOPS on the RTX4090, outperforming FlashAttention2 and xformers by approximately 3.1x and 5.4x, respectively. We extensively evaluate the end-to-end metrics of state-of-the-art text, image, and video generation models using SageAttention2. Across all models and tasks, SageAttention2 can be directly adopted in a plug-and-play manner with negligible loss in model performance.
2 Preliminary
2.1 FlashAttention
The computation of self-attention can be formulated as follows: $S=QK^{\top}/\sqrt{d},~{}P=\sigma(S),~{}O=PV$, where $\sigma(S)_{ij}=\exp(S_{ij})/\sum_{k}\exp(S_{ik})$ is the softmax operation. The matrices $Q$, $K$, and $V$ each have dimensions $N\times d$, while the matrix $S$, $P$ are $N\times N$. While $d$ is typically small, e.g., 64 or 128, $N$ can be thousands if not millions. Therefore, the $N\times N$ matrices $(S,P)$ are much larger than $(Q,K,V)$, and a naive implementation suffers from the huge amount of global memory IO for $(S,P)$ reads/writes. FlashAttention (Dao, 2023) proposes to tile $Q$, $K$, and $V$ from the token dimension into blocks $\{Q_{i}\},\{K_{i}\},\{V_{i}\}$ with block size of $b_{q}$, $b_{kv}$, $b_{kv}$, respectively. Then, to avoid the memory IO for $(S,P)$, it uses online softmax (Milakov & Gimelshein, 2018) to progressively compute each block of $O$, i.e., $O_{i}$:
First, for each block of $\{K_{i}\},\{V_{i}\}$, it computes the following equations iteratively:
$S^{j}_{i}=Q_{i}K_{j}^{\top}/\sqrt{d},~{}~{}(m^{j}_{i},\widetilde{P}^{j}_{i})=\tilde{\sigma}(m^{j-1}_{i},S^{j}_{i}),~{}~{}l_{i}^{j}=$ $\exp(m^{j}_{i}-m^{j-1}_{i})l_{i}^{j-1}+\mathrm{rowsum}(\widetilde{P}^{j}_{i}),$
$O^{j}_{i}=\mathrm{diag}\left(\exp(m^{j}_{i}-m^{j-1}_{i})\right)O^{j-1}_{i}$ $+\widetilde{P}^{j}_{i}V_{j}$
Where $m_{i}^{j}$ and $l_{i}^{j}$ are $b_{q}\times 1$ vectors, which are initialized to $\infty$- ∞ and 00 respectively. The $\tilde{\sigma}()$ is an online softmax operator: $m^{j}_{i}=\max\{m^{j-1}_{i},\mathrm{rowmax}(S^{j}_{i})\},~{}\widetilde{P}_{j}^{i}=\exp(S^{j}_{i}-m^{j}_{i})$.
Finally, the output $O_{i}$ can be computed by $O_{i}=\mathrm{diag}(l_{i}^{j})^{-1}O_{i}^{j}$.
2.2 Quantization
A matrix multiplication $C=AB$ can be accelerated with quantization as:
$(\delta_{A},\hat{A})=\psi(A),~{}~{}(\delta_{B},\hat{B})=\psi(B),~{}~{}\hat{C}=\hat{A}\hat{B},$ $~{}~{}C=\psi^{-1}_{\delta_{A}\delta_{B}}(\hat{C})$ (1)
$\psi$ is a quantizer which converts a high-precision (e.g., FP32 or FP16) matrix $A$ to a low-precision format $\hat{A}$ (e.g., INT4 or FP8) with a scale $\delta_{A}$, and $\psi^{-1}$ is a dequantizer to convert back to high-precision. We should have $\psi^{-1}_{\delta_{A}}(\hat{A})\approx A$. The actual matrix multiplication $\hat{A}\hat{B}$ is carried in low-precision. In modern GPUs, low-precision matrix multiplication is usually multiple times faster than higher-precision ones. Many quantizers depend on the numerical format and granularity, e.g., how many elements share a common scale factor. For example, an INT4 per-tensor quantizer first computes the scale as the maximum absolute value of the entire tensor, scales the elements to the maximum representable range of INT4 [-7, +7], and then casts to INT4 with rounding: $\hat{A}=\lceil A/\delta_{A}\rfloor,\delta_{A}=\max(|{A}|)/7$. Likewise, per-token quantizer assigns a scale factor for each token of a tensor: $\hat{A}[i,:]=\lceil A[i,:]/\delta_{A}[i]\rfloor,\delta_{A}[i]=\max(|{A[i,:]}|)/7$. Also, per-channel quantizer assigns a scale factor for each channel of the tensor, i.e., along the channel dimension: $A[:,i]=\lceil A[:,i]/\delta_{A}[i]\rfloor,\delta_{A}[i]=\max(|A[:,i]|)/{7}$. Dequantization process is a element-wise scaling: $\psi_{\delta_{A}\delta_{B}}^{-1}(\hat{A}\hat{B})=\hat{A}\hat{B}*\delta_{A}*\delta_{B}$.
2.3 SageAttention
Based on the block tiling in FlashAttention2 (Dao, 2023), SageAttention (Zhang et al., 2024) quantize $Q,K$ to INT8 in a per-block granularity. Moreover, to keep the quantization accuracy, SageAttention proposes to smooth $K$ first:
$K=K$ $-\text{mean}(K)$
$\hat{Q}_{i}=\lceil Q_{i}/\delta_{Q}\rfloor,~{}~{}\delta_{Q}=\max(|{Q_{i}|})/127,$ $~{}~{}\hat{K}_{j}=\lceil K_{j}/\delta_{K}\rfloor,~{}~{}\delta_{K}=\max(|{K_{j}|})/127$
$S_{ij}=$ $Q_{i}K_{j}^{\top}*\delta_{Q}*\delta_{K}$
Where $Q_{i},K_{j}$ are the tilled block in FlashAttention, and $\text{mean}(K)$ is the average value of the channel dimension of $K$. SageAttention keeps $\widetilde{P}V$ as FP16 and uses FP16 Matmul with an FP16 accumulator for $\widetilde{P}V$. However, FP16 Matmul with an FP16 accumulator only has a speedup effect on RTX4090 and RTX3090 GPUs.
3 SageAttention-2
Figure 3: Workflow of SageAttention2. 1 Smooth Q,K,V. 2 A GEMV to obtain $\Delta S$. 3 Per-warp quantize Q,K and per-channel quantize V. 4 Perform the SageAttention2 kernel. 5 Correct the output.
3.1 Formulation
Based on the introduction of FlashAttention and quantization presented in Section 2, we now describe the quantized attention approach we developed.
Quantization: ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{Q},\hat{Q},~{}~{}\overrightarrow{Q}_{m})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\phi_{Q}}}(Q),~{}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{K},\hat{K})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\phi_{K}}}(K),~{}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{V},\hat{V},\overrightarrow{V}_{m})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\phi_{V}}}(V)$
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\Delta S=\overrightarrow{Q_{m}}K^{\top}}}$ (2)
Attention: $S={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi^{-1}_{\delta_{Q}\delta_{K}}(\hat{Q}\hat{K}^{\top})+\Delta S}},~{}~{}(m^{\prime},\widetilde{P})=\tilde{\sigma}(m,S),~{}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{P},\hat{P})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{P}}}(\widetilde{P})$
$O=\mathrm{diag}\left(\exp(m^{\prime}-m)\right)O+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi^{-1}_{\delta_{P}\delta_{V}}(\hat{P}\hat{V})}}$ (3)
$\phi_{Q}$, $\phi_{K}$, $\phi_{V}$ are three transformations to obtain quantized $Q$, $K$, and $V$, which we will discuss in subsequent sections. For simplicity, we omit all superscripts and subscripts, but the matrices used in attention are still tiles, and the computation is still organized as FlashAttention described in Section 2.1. Compared to the original full-precision version, as shown in Eq. 23, SageAttention2 adds quantizers to $Q,K,P,V$ and dequantizers to the product to accelerate both Matmuls of $QK$ and $\widetilde{P}V$.
Figure 4: Typical examples of tensors’ data distribution in attention.
Table 1: End-to-end metrics comparison of different quantization methods, where Q,K are quantized into INT4, while P,V stay in full precision.
Q, K Smoothing (Q+K) Llama3.1 (Lambda) $\uparrow$ Llama3.1 (WikeText) $\downarrow$ CogVideo (vqa-a) $\uparrow$ CogVideo (vqa-t) $\uparrow$
Full-Precision - 81.5% 6.013 77.605 75.360
INT4 Quantization 72.6% 11.698 27.114 24.670
80.8% 6.219 77.276 75.147
Table 2: Average accuracy across all layers using different quantization granularities.
Method Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
Per-token 99.45% 0.0649 0.0335
Per-warp 99.45% 0.0648 0.0334
Per-block 98.03% 0.1492 0.0744
Per-tensor 97.15% 0.1800 0.0865
Table 3: Worst accuracy across all layers using different quantization granularities.
Method Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
Per-token 96.76% 0.1916 0.0775
Per-warp 96.71% 0.1956 0.0779
Per-block 90.68% 0.3615 0.1490
Per-tensor 85.85% 0.4687 0.2261
3.2 Per-warp INT4 Quantization
SageAttention uses per-block quantization, which quantizes each block of $Q$ and $K$ for a GPU streaming processor. Such a quantization strategy could achieve an accuracy performance close to per-token quantization and avoid the overhead of the dot product of the quantization scale vectors $\delta_{Q}$ and $\delta_{K}$. However, quantizing $Q$ and $K$ to INT4 demands a more accurate quantization granularity. We propose per-warp quantization, a more precise and granular quantization approach than the per-block quantizer, also with no additional overhead of vectors dot product.
Specifically, each block of $Q$ in SageAttention will be split into $c_{w}$ segments, each processed by one of $c_{w}$ GPU warps in a GPU streaming processor (SM). Subsequently, each warp performs Matmul for its assigned segment. Per-warp INT4 quantization assigns a scale factor for every $b_{q}/{wc}$ tokens for $Q$:
$\hat{Q}[\frac{i*b_{q}}{c_{w}}$ $:\frac{b_{q}*(i+1)}{c_{w}},:]=\left\lceil\frac{Q[\frac{i*b_{q}}{c_{w}}:\frac{b_{q}*(i+1)}{c_{w}},:]}{\delta_{Q}[i]}\right\rfloor$ (4)
$\delta_{Q}[i]=\frac{\max(\left|\,Q[\frac{i*b_{q}}{c_{w}}:\frac{b_{q}*(i+1)}{c_{w}},:]\,\right|)}{7}$ (5)
Per-warp quantization will provide an accuracy performance superior to per-block quantization used in SageAttention, which we will discuss in Section 3.5.
Figure 5: An example of quantized value distribution of $Q$ before and after smoothing $Q$.
3.3 Smooth Q
The representative range for quantization of INT4 is notably restrictive, i.e., only $152^{4}-1=$ 15 values. This limitation significantly degrades performance when quantizing $Q,K$ into INT4 in attention. For example, using INT4 to quantize $Q,K$ can cause the perplexity of Llama3.1 on WikiText to increase by more than 90%, and the quality of videos generated by CogvideoX decrease about 3x (See Table 1). We analyze the data distribution of $Q,K$ in real models. For example, we find that Llama3.1 and CogvideoX show distinct channel-wised outliers in $Q,K$ (See Figure 4). While per-channel quantization could mitigate the quantization error caused by such outliers, this method is impracticable for $Q,K$ because quantization must be conducted along the outer axis (token dimension) of $QK^{\top}$.
To address this challenge, we propose a method to eliminate the quantization error brought by the outliers in $Q,K$:
$\overrightarrow{Q_{m}}=\mathrm{mean}(Q),~{}~{}\gamma(Q)=$ $Q-\overrightarrow{Q_{m}},~{}~{}\gamma(K)=K-\mathrm{mean}(K),$
$S=QK=$ $\gamma(Q)\gamma(K)+\overrightarrow{Q_{m}}\gamma(K)$
$\phi_{Q}(Q)=\psi_{Q}\circ\gamma(Q),~{}\overrightarrow{Q_{m}}$ $,~{}~{}~{}~{}~{}\phi_{K}(K)=\psi_{K}\circ\gamma(K)$ (6)
Where $\mathrm{mean}(K)$ and $\gamma(K)$ have been discussed in SageAttention (Zhang et al., 2024), and the transformation for $K$ does not change the attention score $\widetilde{P}$. $\mathrm{mean}(Q)=\frac{1}{N}\sum_{t=1}^{N}Q[t,:]$ is a vector with a shape $1\times d$. For the transformation of $Q$, we will perform a GEMV (general matrix-vector multiplication) between $\overrightarrow{Q_{m}}$ and $K^{\top}$, i.e., $\overrightarrow{Q_{m}}K^{\top}$ as $\Delta S$. This $\Delta S$ will be added to $S$ during the attention computation. Finally, the transformation from full-precision $Q,K$ to quantized $\hat{Q},\hat{K}$ can be expressed as shown in Eqution 6, where $\psi_{Q}$, $\psi_{K}$ is two quantizers for $Q$ and $K$. In other words, full-precision $Q,K$ are substracted with the means of the channel dimension before being quantized.
Figure 5 shows an example from CogvideoX of the distribution of quantized $Q$ with and without smoothing Q. We can find that with smoothing Q, the range of INT4 is utilized more uniformly and fully. Table 1 presents end-to-end metrics for different quantization methods with and without smoothing Q+K on Llama3.1 (Dubey et al., 2024) and CogvideoX (Yang et al., 2024). The results demonstrate that smoothing Q+K offers significant accuracy benefits. Also, Table 10 and Table 10 show that the order of effectiveness is smoothing Q+K $>$ smoothing Q $>$ smoothing K $>$ other baselines.
Algorithm 1 Implementation of SageAttention2.
  Input: Matrices $Q(\text{FP16}),K(\text{FP16}),V(\text{FP16})\in\mathbb{R}^{N\times d}$, block size $b_{q},b_{kv}$, warp count $c_{w}$.
  Preprocessing: $K=K-\mathrm{mean}(K)$, $\overrightarrow{Q}_{m}=\mathrm{mean}(Q)$, $Q=Q-\overrightarrow{Q}_{m}$, $\Delta S=\text{GEMV}(\overrightarrow{Q}_{m},K^{\top})$, $\overrightarrow{V}_{m}=\mathrm{mean}(V),V=V-\overrightarrow{V}_{m}$.
  Quantization: ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{Q},\hat{Q})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{Q}}}(Q)$ // INT4 Per-warp. ${}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{K},\hat{K})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{K}}}(K)$. ${}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{V},\hat{V})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{V}}}(V)$. // FP8 Per-channel.
  Divide $\hat{Q}$ to $T_{m}={N}/{b_{q}}$ blocks $\{\hat{Q}_{i}\}$; divide $\hat{K}$, $\hat{V}$, and $\Delta S$ to $T_{n}={N}/{b_{kv}}$ blocks $\{\hat{K}_{i}\}$, $\{\hat{V}_{i}\}$, and $\{\Delta S_{i}\}$;
  for $i=1$ to $T_{m}$ do
     Load $\hat{Q}_{i}$ and $\delta_{Q}[i*c_{w}:(i+1)*c_{w}]$ into an SM ;
     for j in [1, $T_{n}$do
        Load $\hat{K}_{j}$, $\hat{V}_{j}$, and $\delta_{K}[j]$ into the SM ;
        $w=\text{range}(c_{w}),st=w*c_{w}$
        $S_{i}^{j}[st:st+c_{w}]=\mathrm{Matmul}(\hat{Q}_{i}[st:st+c_{w}],\hat{K}_{j}^{\top})\times\delta_{Q}[st+w]\times\delta_{K}[j]$ + $\Delta S_{j}$; // Paralleled by $c_{w}$ warps.
        $m_{i}^{j}=\mathrm{max}(m_{i}^{j-1},\mathrm{rowmax}(S_{i}^{j}))$, $\widetilde{P}_{i}^{j}=\mathrm{exp}(S_{i}^{j}-m_{i}^{j})$, $l_{i}^{j}=e^{m_{i}^{j-1}-m_{i}^{j}}+\mathrm{rowsum}(\widetilde{P}_{i}^{j})$ ;
        $O_{i}^{j}=\mathrm{diag}(e^{m_{i}^{j-1}-m_{i}^{j}})^{-1}O_{i}^{j-1}+$ $\mathrm{Matmul}((\widetilde{P}_{i}^{j}*448)\text{.to(FP8.e4m3)},V_{j})$ ;
     end for
     Load $\delta_{V}$ into an SM;
     $O_{i}=\mathrm{diag}(l_{i}^{T_{n}})^{-1}O_{i}^{T_{n}}/448*\delta_{V}$ ;
     Write $O_{i}$ ;
  end for
  $O=\{O_{i}\}$, $O=O+\overrightarrow{V}_{m}$.
  return $O$
Table 4: Average accuracy using different data types of $(\widetilde{P},V)$ across all layers of a CogvideoX model, where $(Q,K)$ are smoothed.
$Q,K$ $\widetilde{P},V$ Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
INT4 INT8 77.05% 0.5618 0.5044
E5M2 99.20% 0.0905 0.0903
E4M3 99.44% 0.0683 0.0347
FP16 99.45% 0.0649 0.0335
Table 5: Worst accuracy using different data types of $(\widetilde{P},V)$ across all layers of a CogvideoX model, where $(Q,K)$ are smoothed.
$Q,K$ $\widetilde{P},V$ Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
INT4 INT8 19.52% 0.9579 1.4483
E5M2 94.94% 0.2327 0.2361
E4M3 96.70% 0.1956 0.0779
FP16 96.76% 0.1916 0.0775
Table 6: An accuracy example on real tensors of CogvideoX model with or without smoothing $V$.
Smooth V Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
98.25% 0.1980 0.2387
99.75% 0.0406 0.0773
Figure 6: An example of dot product precison a row of $\widetilde{P}$ and a column of $V$ presented by FP22 data type.
Table 7: Error of the FP8 Matmul instruction of mma(f8f8f32).
Precision of Accumulated Value E8M13 E8M23
Error compared to FP32 0 ||||mma(f32.f16.f16.f32) - mma(f32.f8.f8.f32)||||
3.4 Smooth V
To avoid the drawback of SageAttention where FP16 Matmul with an FP16 accumulator for $PV$ is only effective on GPUs such as the RTX4090, we adopt the approach of quantizing $P$ and $V$ to FP8 to leverage the universal acceleration of FP8 Tensor cores. However, we discover that the accumulator for the mma(f32f8f8f32) instruction on the Ada architecture is actually FP22, specifically with 1 sign bit, 8 exponent bits, and 13 mantissa bits. Specifically, for mma(f32f8f8f32) instruction $C=AB+D$, where $A,B$ are tensors in FP8 data type and $C,D$ are tensors with FP32 data type, we initialize the $A,B$ to zero and vary $D$ to test the data type of the accumulator. As shown in Table 7, when $D$ is initialized with 1 sign bit, 8 exponent bits, and 13 mantissa bits, the value of $C$ matches the result of the mma(f16f16f32) instruction. However, when $D$ is initialized with more than 13 mantissa bits, the error of $C$ corresponds to the difference between the results of mma(f32f16f16f32) and mma(f32f8f8f32). Consequently, matrix multiplication of matrices $\widetilde{P}$ and $V$, quantized to FP8, incurs a certain degree of precision loss compared to using an FP32 accumulator. To mitigate this precision loss as much as possible, we propose to smooth $V$:
$\gamma(V)=V-\mathrm{mean}(V),$ $~{}~{}~{}~{}\mathrm{mean}(V)=\overrightarrow{V_{m}}$
$\phi_{V}(V)=$ $\psi_{V}\circ\gamma,~{}\overrightarrow{V_{m}}$ (7)
As shown in Figure 6, this strategy enhances the accuracy of FP22 for values in $\widetilde{P}V$ for the following reasons: Each row of $\widetilde{P}$ spans a value range from 0 to 1, and each column of $V$ consistently features channel-wise outliers that are exclusively positive or negative, typically ranging between 8 and 9. Consequently, the values of $\widetilde{P}V$ could be quite large. However, the floating-point number representation range is not uniform—it is denser near zero. Therefore, by subtracting the mean along the channel dimension from $V$, the values of $\widetilde{P}V$ will be closer to zero, resulting in higher representational precision. We believe such an explanation and strategy could address lots of issues [^1] https://github.com/triton-lang/triton/issues/4476, https://github.com/triton-lang/triton/issues/5065 reported by the community. Additionally, to maintain the correctness of the attention computation, it is only necessary to add $\overrightarrow{V_{m}}$ to the final calculation of $O$: $O=O+\overrightarrow{V_{m}}$. This is because the sum of each row of the $\widetilde{P}$ matrix equals 1, so $\widetilde{P}\overrightarrow{V_{m}}=\overrightarrow{V_{m}}$. In other words, this method decomposes $V$ into two parts: $\overrightarrow{V_{m}}$ and $V$. For $V$, it centers the values of each column around zero, which leads to the dot product result of a row from the quantized $\widetilde{P}$ matrix with a column from the quantized $V$ matrix being closer to zero. This makes the representation of FP22 more accurate. Meanwhile, $\overrightarrow{V_{m}}$ is retained in FP16 and added to $O$ at the end without causing a loss of precision for the $\overrightarrow{V_{m}}$ part.
Table 6 shows the attention accuracy on real tensors sampled from CogvideoX with and without smoothing $V$. It demonstrates that smoothing V could improve the accuracy of SageAttention2 when quantizing $Q,K$ to INT4 and $\widetilde{P},V$ to FP8.
3.5 Quantization for Q, K, P, V
Quantization for $Q,K$. We propose $\psi_{Q}(Q)$ and $\psi_{K}(K)$ in granularity of per-warp. This is first because per-channel quantization is not feasible since the scale factors of the inner axis of $QK^{\top}$ cannot used to do dequantization (Xiao et al., 2023a). Second, as shown in Table 3 and Table 3, we compare the average and worst-case accuracies of INT4 quantization at per-token, per-warp, per-block, and per-tensor granularity using real $Q,K,V$ across all layers of CogvideoX. Results indicate that the accuracy of per-warp quantization is very close to per-token and outperforms a lot than per-block and per-tensor. Furthermore, per-warp quantization incurs less dequantization overhead than per-token as discussed in Section 3.2.
Quantization for $\widetilde{P},V$. We choose FP8, specifically the E4M3 data type, for $\psi_{P}(\widetilde{P})$ and $\psi_{V}(V)$ for two reasons. First, most GPUs have tensor cores that support FP8 Matmul operations, which are twice as fast as those using FP16. Second, Table 5 and Table 5 show the average and worst accuracy of different data types used for $\widetilde{P},V$ using real $Q,K,V$ across all layers of CogvideoX. We can see that the accuracy of using E4M3 is very close to that of using FP16 and superior to that of E5M2 and INT8. We propose to use $\psi_{P}(\widetilde{P})$ in per-block and $\psi_{V}(V)$ in per-channel for three reasons. First, per-channel quantization for $\widetilde{P}$ and per-token quantization for $V$ are not viable because dequantization requires scale factors of outer axis. Second, $\widetilde{P}=\mathrm{exp}(S_{i}-\mathrm{rowmax}(S_{i}))$, where $S_{i}$ is the Matmul result of a block of $Q$ and $K^{\top}$, the max value in each row of $\widetilde{P}$ is 1. Hence, we can assign a single static scale $s=\frac{1}{448}$ to a block $\widetilde{P}$, whose accuracy equals per-token quantization. Third, per-channel quantization can address the channel-wised outlier of $V$.
Accuracy metrics. We use three metrics to assess the accuracy of quantized attention output $O^{\prime}$ compared to attention output in full-precision $O$: First, we flatten $O^{\prime}$ and $O$ into vectors in the shape of $1\times n$. Then, Cosine similarity: $CosSim=\sum OO^{\prime}/\smash{\sqrt{\sum O^{2}}}\smash{\sqrt{\sum O^{\prime 2}}}$, Relative L1 distance: $L1=\sum|O-O^{\prime}|/\sum|O|$, Root mean square error: $RMSE=\sqrt{(1/n)\sum(O-O^{\prime})^{2}}$.
Table 8: Two kernel implementations of SageAttention2.
     Kernel      $\psi_{Q}(Q)$, $\psi_{K}(K)$      $\psi_{P}(\widetilde{P})$, $\psi_{V}(V)$
     SageAttn2-4b      INT4, per-warp      FP8, per-block;  FP8, per-channel
     SageAttn2-8b      INT8, per-warp      FP8, per-block;  FP8, per-channel
Figure 7: Mean and standard deviation of $cossim*(1-L1)$) of SageAttn-4b in different layers and timesteps for different inputs in Llama3.1 and CogvideoX.
3.6 Adaptive Quantization over Layer and Timestep
Based on the discussion in Section 3.5, we implement two attention kernels (See Table 8) based on a choice: Using INT4 or INT8 quantization for $\psi_{Q}(Q)$ and $\psi_{K}(K)$. The speed of these kernels is in the order (SageAttn2-4b$>$SageAttn2-8b), but the accuracy order is opposite. Although SageAttn2-4b could perform adequately in lots of tasks, for instance, it shows very slight end-to-end metric loss of Lambda benchmark on Llama3.1, SageAttn2-4b falls short in some hard scenarios. Consequently, we first conduct a detailed analysis of the accuracy performance on various models using SageAttn2-4b and SageAttn2-8b. Then, we propose an effective and easy-to-use adaptive method.
Accuracy across layers and timesteps. Figure 7 displays the mean and standard deviation of the accuracy, calculated using $CosSim*(1-L1)$, for SageAttn-4b across different layers and timesteps of different inputs in Llama3.1 and CogvideoX. It can be observed that specific layers and timesteps exhibit specific ranges of errors.
Our method. Our adaptive strategy is detailed as follows: First, we assess the average accuracy of SageAttn2-4b in comparison to full precision attention for each (layer, timestep) combination within a model, using several random inputs. For models without a timestep dimension, such as language models, the evaluation is confined to the layer dimension. Second, we will sort the (layer, timestep) combinations in descending order based on their $CosSim*(1-L1)$ values. We then identify the smallest $\xi$of these combinations. Finally, for the identified (layer, timestep) combinations, we consistently apply SageAttn2-8b across all other prompts. We call such a strategy SageAttn-mix and will report the values of $\xi$we used for each model in Section 4.1.
Figure 8: Speed comparison between SageAttention2 and baselines (RTX4090, headdim=64).
Figure 9: Speed comparison between SageAttention2 and baselines (RTX4090, headdim=128).
Figure 10: Speed comparison between SageAttention2 and baselines (RTX4090, headdim=256).
Figure 11: Speed comparison between SageAttention2 and baselines (L20, headdim=64).
Figure 12: Speed comparison between SageAttention2 and baselines (L20, headdim=128).
Figure 13: Speed comparison between SageAttention2 and baselines (L20, headdim=256).
4 Experiment
4.1 Setup
Models. We validate the effectiveness of SageAttention2 across a diverse set of representative models from language, image, and video generation. Specifically, we conduct experiments on five models: Llama2 (7B) (Touvron et al., 2023), Llama3.1 (8B) (Dubey et al., 2024), and GLM4 (9B) (GLM et al., 2024) for text2text, CogvideoX (2B) (Yang et al., 2024) and Open-Sora (Zheng et al., 2024) for text2video, Flux (schnell) (Black Forest Labs, 2023) for text2image, and TIMM (Wightman, 2019) for image classification.
Datasets. Text-to-text models are evaluated on four zero-shot tasks: WikiText (Merity et al., 2022) to assess the model’s prediction confidence, LAMBADA (Paperno et al., 2016) evaluate contextual understanding, MMLU (Hendrycks et al., 2020) for measuring knowledge across various subjects, and Longbench (Bai et al., 2023) for evaluating bilingual, multitask, and comprehensive assessment of long context understanding capabilities. Text-to-video models are evaluated using the open-sora (Zheng et al., 2024) prompt sets. Flux is assessed on MJHQ-30K (Li et al., 2024). TIMM is evaluated on on three image datasets: ImageNet (Deng et al., 2009), ImageNet-Sketch (Sketch) (Wang et al., 2019), and ImageNet-Rendition (ImageNet-r) (Hendrycks et al., 2021).
Metrics. For text-to-text models, we use perplexity (ppl.) (Jelinek et al., 1977) for WikiText, Accuracy (Acc.) for LAMBADA and MMLU, and Longbench score (Bai et al., 2023). For text-to-video models, following Zhao et al. (2024), we evaluate the quality of generated videos on five metrics: CLIPSIM and CLIP-Temp (CLIP-T) (Liu et al., 2024) to measure the text-video alignment; (VQA-a) and (VQA-t) to assess the video aesthetic and technical quality, respectively; and Flow-score (FScore) for temporal consistency (Wu et al., 2023). For text-to-image models, generated images are compared with the images in MJHQ-30K dataset in three aspects: FID (Heusel et al., 2017) and sFID (Salimans et al., 2016) for fidelity evaluation, Clipscore (CLIP) (Hessel et al., 2021) for text-image alignment, and ImageReward (IR) (Xu et al., 2024) for human preference. For TIMM, we use Accuracy.
Implemetation details. We implement our attention kernels using CUDA and conduct experiments on Ubuntu 22.04 servers with RTX4090 and L20 GPUs.
Baselines. (1) SmoothAttn. Following Qserve (Lin et al., 2024), we apply smooth quant for matrixes $Q,K$ with smoothing factor $\alpha=0.5$. (2) HadmdAttn. Following Quarot (Ashkboos et al., 2024), we apply random Hadamard transformation for matrixes $Q,K$ before INT4 quantization.
Hyper-parameters of SageAttn-mix. We use $\xi$ of 50% for Llama2, 60% for CogvideoX, 25% for Open-Sora, 30% for Llama3.1 and GLM4, 0% for Flux, and TIMM.
Figure 14: Comparison examples from CogvideoX, prompts are sampled from open-sora prompt sets.
Table 9: Average accuracy across all layers using different smooth methods.
Method CosSim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
None 80.04% 0.3906 0.2223
HadmdAttn 79.77% 0.3782 0.2180
SmoothAttn 90.21% 0.3383 0.1952
Smooth K (Ours) 98.07% 0.1493 0.0743
Smooth Q (Ours) 98.30% 0.1250 0.0712
SageAttn2-4b 99.46% 0.0648 0.0334
Table 10: Worst accuracy across all layers using different smooth methods.
Method CosSim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
None 4.83% 0.9979 0.7784
HadmdAttn 4.85% 0.9978 0.7785
SmoothAttn 64.49% 0.9262 0.7204
Smooth K (Ours) 90.86% 0.3565 0.1464
Smooth Q (Ours) 93.10% 0.2989 0.2195
SageAttn2-4b 96.71% 0.1956 0.0779
Table 11: End-to-end metrics loss across text, image, and video generation models.
Model Attention WikiText (Ppl.) $\downarrow$ Lambda (Acc.) $\uparrow$ MMLU (Acc.) $\uparrow$ Longbench $\uparrow$
Llama2 Full-Precision 5.823 0.886 0.439 -
HadmdAttn 6.706 0.865 0.355 -
SmoothAttn 6.690 0.871 0.395 -
SageAttn2-4b 6.018 0.886 0.436 -
SageAttn2-mix 5.883 0.883 0.431 -
Llama3.1 Full-Precision 6.013 0.815 0.635 49.40
HadmdAttn 7.661 0.756 0.502 44.62
SmoothAttn 7.087 0.788 0.551 43.76
SageAttn2-4b 6.219 0.808 0.617 48.61
SageAttn2-mix 6.131 0.816 0.629 49.01
GLM4 Full-Precision 7.241 0.432 0.743 49.78
HadmdAttn 7.932 0.435 0.676 46.27
SmoothAttn 8.841 0.442 0.599 43.10
SageAttn2-4b 7.341 0.435 0.732 49.06
SageAttn2-mix 7.303 0.434 0.737 49.77
Model Attention CLIPSIM $\uparrow$ CLIP-T $\uparrow$ VQA-a $\uparrow$ VQA-t $\uparrow$ FScore $\uparrow$
CogvideoX Full-Precision 0.1836 0.9975 77.605 75.360 3.006
HadmdAttn 0.1742 0.9877 36.028 23.786 0.550
SmoothAttn 0.1763 0.9870 37.444 42.184 0.792
SageAttn2-4b 0.1813 0.9969 77.276 75.147 2.070
SageAttn2-mix 0.1816 0.9976 75.686 78.600 2.884
Open-Sora Full-Precision 0.1831 0.9996 46.713 59.553 0.368
SageAttn2-4b 0.1821 0.9994 42.270 55.965 0.364
SageAttn2-mix 0.1814 0.9994 44.509 59.097 0.383
        Model         Attention        FID $\downarrow$        sFID $\downarrow$        CLIP $\uparrow$        IR $\uparrow$
        Flux         Full-Precision        11.303        17.603        32.603        0.9745
        HadmdAttn        11.163        17.693        32.592        0.9638
        SmoothAttn        10.941        18.098        32.582        0.9613
        SageAttn2-4b        10.563        17.052        32.631        0.9747
Model Attention ImageNet (Acc.) $\uparrow$ Sketch (Acc.) $\uparrow$ ImageNet-r (Acc.) $\uparrow$
TIMM Full-Precision 84.79% 45.32% 59.55%
SageAttn2-4b 84.67% 45.07% 59.11%
4.2 Speed and Accuracy of Kernels
Speed. We conduct experiments to compare the Speed of SageAttention2 against baselines using configurations with headdim=64, headdim=128, and headdim=256, both with and without Causal Mask (Vaswani, 2017). Specifically, Figure 8, Figure 9, and Figure 10 show the speed of SageAttention2 and baselines across varying sequence lengths on RTX4090. These results indicate that SageAttention2 achieves a peak of 485 TOPS and is 3.1x faster than FlashAttention2 and 5.4x faster than xformers. Figure 11, Figure 12, and Figure 13 illustrate the results on L20 GPU, indicating that SageAttention2 achieves a peak of 288 TOPS and is 2.7x faster than FlashAttention2 and 4.6x faster than xformers.
Accuracy. Table 10 and Table 10 show that the average and worst accuracy of different methods with INT4 $Q,K$ and FP8 $P,V$ across all layers of CogvideoX. The results indicate that the accuracy of SageAttn-4b is superior to other baselines.
Figure 15: End-to-end performance using different ratio of SageAttn-8b.
4.3 End-to-end Performance
Metrics loss. We assessed the end-to-end metrics of various models using SageAttention2 compared to using attention in full precision. Detailed evaluation results are presented in Table 11 for Llama2, Llama3.1, GLM4, CogvideoX, Open-Sora, Flux, and TIMM, respectively. The results indicate that SageAttn-4b outperforms all baselines and maintains most of the end-to-end accuracy across all models. Additionally, using adaptive quantization techniques, SageAttn-mix achieves performance comparable to full-precision attention.
4.4 Ablation Study
Adaptive quaitzation. To analyze the impact of using different ratio of SageAttn-8b in adaptive quantization, Table 15 shows the changes in the perplexity of Llama3.1 under various ratio of SageAttn-8b. It can be observed that even if SageAttn-4b is used exclusively, the overall end-to-end representation is sufficiently good, and the higher the ratio of SageAttn-8b used, the closer the accuracy approaches that of using full-precision attention.
Overhead of smoothing Q, K. Note that the overhead of smoothing $Q,K$ only contains $Q-\overrightarrow{Q_{m}}$, $K-\mathrm{mean}(K)$, and $\overrightarrow{Q_{m}}K^{\top}$. The speed overhead of them accounts for about 3.5% of the attention kernel.
5 Conclusion and Future Work
We introduce SageAttention2, an efficient and accurate 4-bit quantization method for attention. First, we propose to quantize $(Q,K)$ to INT4 in a warp-level granularity and quantize $(\widetilde{P},V)$ to FP8. Second, we propose a method to smooth matrixes $Q$ and $V$, enhancing the accuracy of attention with INT4 $QK$ and FP8 $PV$. Third, we analyze the quantization accuracy across timesteps and layers, then propose an adaptive quantization method to ensure the end-to-end metrics over various models. Our implementation is faster than FlashAttention2 and xformers by approximately 3.1x and 5.4x on RTX4090, respectively. Extensive testing confirms that our approach maintains end-to-end metrics across various models, including language, image, and video generation.
Future Work. We leave the implementation of FP8 MatMul with an FP16 accumulator for $\widetilde{P}V$ and SageAttention2 on the Hopper architecture for future work.

### v3

@@
-SageAttention2 Technical Report: Accurate 4 Bit Attention for Plug-and-play Inference Acceleration
+SageAttention2: Efficient Attention with Thorough Outlier Smoothing and Per-thread INT4 Quantization
 Abstract
 Although quantization for linear layers has been widely used, its application to accelerate the attention process remains limited.
-SageAttention utilizes 8-bit matrix multiplication, 16-bit matrix multiplication with 16-bit accumulator, and precision-enhancing methods, implementing an accurate and 2x speedup kernel compared to FlashAttention2.
-To further enhance the efficiency of attention computation while maintaining precision, we propose SageAttention2, which utilizes significantly faster 4-bit matrix multiplication (Matmul) alongside additional precision-enhancing techniques.
-First, we propose to quantize matrixes $(Q,K)$ to INT4 in a warp-level granularity and quantize matrixes $(\widetilde{P},V)$ to FP8.
-Second, we propose a method to smooth $Q$ and $V$, enhancing the accuracy of attention with INT4 $QK$ and FP8 $PV$.
-Third, we analyze the quantization accuracy across timesteps and layers, then propose an adaptive quantization method to ensure the end-to-end metrics over various models.
+To further enhance the efficiency of attention computation compared to SageAttention while maintaining precision, we propose SageAttention2, which utilizes significantly faster 4-bit matrix multiplication (Matmul) alongside additional precision-enhancing techniques.
+First, we propose to quantize matrixes $(Q,K)$ to INT4 in a hardware-friendly thread-level granularity and quantize matrixes $(\widetilde{P},V)$ to FP8.
+Second, we propose a method to smooth $Q$, enhancing the accuracy of INT4 $QK^{\top}$.
+Third, we propose to use an FP32 Matmul buffer for $PV$ to enhance the accuracy of FP8 $\widetilde{P}V$.
 The operations per second (OPS) of SageAttention2 surpass FlashAttention2 and xformers by about 3x and 5x on RTX4090, respectively.
@@
 footnotetext: * Equal Contribution.
-Figure 1: The left figure shows the kernel speedup on RTX4090 GPU.
-The right figure shows the end-to-end inference speedup of generating the first token and performance metrics for the needle-in-a-haystack task (gkamradt, 2023) with a sequence length of 100K on Llama3.1 on L20 GPU.
+Figure 1: The upper left figure shows the kernel speedup on RTX4090 GPU.
+The upper right figure shows the end-to-end inference speedup of generating the first token and performance metrics for the needle-in-a-haystack task (gkamradt, 2023) with a sequence length of 100K on Llama3.1 on L20 GPU.
+The figure below shows two videos generated by CogvideoX (1.5-5B) using FlashAttention2 and SageAttention2 on RTX4090, showing that SageAttention2 accelerates generation 1.8x with no video quality loss.
 1 Introduction
-Due to the quadratic complexity of attention, efficient implementation becomes increasingly crucial as sequences lengthen in real-world applications (Jiang et al., 2024).
+Due to the quadratic time complexity of attention, efficiently implementing it becomes crucial as sequence lengths increase in real-world applications (Jiang et al., 2024).
 Several strategies have been developed to mitigate computational demands of attention —such as (1) Linear attention methods (Wang et al., 2020; Choromanski et al., 2020; Yu et al., 2022; Katharopoulos et al., 2020) that reduce complexity to $O(N)$ and (2) Sparse attention methods (Liu et al., 2021; Chu et al., 2021; Li et al., ; Xiao et al., 2023b, 2024; Chen et al., 2023; Jiang et al., 2024; Venkataramanan et al., 2023; Gao et al., 2024; Fu et al., 2024) that selectively process parts of the context — these methods are only suitable for a limited range of models and tasks.
@@
 Those works do not omit the computation of attention over parts of the context and achieve impressive speed and precision performance across various models and tasks.
-For the two matrix multiplication (Matmul) operations in attention: $QK^{\top}$ and $PV$, SageAttention accelerates them by quantizing the $QK^{\top}$ to INT8 and uses FP16 Matmul with FP16 accumulators for $PV$.
+Motivation.
+For the two matrix multiplication (Matmul) operations in attention: $QK^{\top}$ and $\widetilde{P}V$, SageAttention accelerates them by quantizing the $QK^{\top}$ to INT8 and uses FP16 Matmul with FP16 accumulators for $\widetilde{P}V$.
 Moreover, to keep the accuracy of attention, SageAttention proposes smoothing $K$ by eliminating its channel-wise outliers.
@@
 (W2) FP16 Matmul with FP16 accumulators is only compatible with RTX4090 and RTX3090 GPUs.
-To leverage the faster INT4 tensor cores for $QK^{\top}$ and using a method that can generally accelerate $PV$, we propose SageAttention2, which quantizes $Q,K$ to INT4 and $P,V$ to FP8.
+To leverage the faster INT4 tensor cores for $QK^{\top}$ and using a method that can accelerate $\widetilde{P}V$ on a broader range of GPUs, we propose SageAttention2, which quantizes $Q,K$ to INT4 and $\widetilde{P},V$ to FP8.
 Challenges.
-Quantizing $Q,K$ to INT4 and $P,V$ to FP8 presents significant challenges.
+Quantizing $Q,K$ to INT4 and $\widetilde{P},V$ to FP8 presents significant challenges.
 For example, when only per-tensor quantizing $Q,K$ to INT4, the text-to-video model CogvideoX will generate a completely blurry video (See Figure 2), and Llama2 only achieves a random-guessing-level accuracy of 25% on MMLU.
-After investigating deeply, we identified three primary challenges: (C1) The numerical range of INT4, which in quantization typically spans from $-7$ to $7$ and totals 15 numbers (Lin et al., 2024), leads to significant quantization errors when $Q$ and $K$ have some abnormal values.
-(C2) In specific layers and timesteps (for text-to-image/video) of some models, quantizing $Q$ and $K$ to INT4 and quantizing $P$ and $V$ to FP8 introduces noticeable errors in attention computation.
-These errors in worst-case layers/timesteps significantly impact the precision of the end-to-end output.
+After investigating deeply, we identified three primary challenges: (C1) The numerical range of INT4 is quite limited compared to FP16 or INT8, leading to significant quantization errors when $Q$ and $K$ have some abnormal values.
 (C2) We discover that the FP32 accumulator designed for FP8 matrix multiplication in the tensor core (mma.f32.f8.f8.f32) is actually FP22, specifically with 1 sign bit, 8 exponent bits, and 13 mantissa bits.
-This will lead to an accuracy loss of $PV$.
+This will lead to an accuracy loss of $\widetilde{P}V$.
 Figure 2: An example of quantizing Q, K to INT4 from CogvideoX.
 Our approach.
-To address these challenges, we detailedly analyze the reasons and propose two methods.
-First, for the significant channel-wise outliers in matrices $Q$ and $K$, we adopt smoothing $K$ in SageAttention and propose an effective method to remove these outliers in $Q$.
+To address (C1), we first discovered that the per-block quantization of $Q,K$ in SageAttention does not achieve sufficient accuracy for INT4 quantization.
+Simultaneously, to avoid the extra latency caused by per-token dequantization, where each GPU thread corresponds to multiple quantization scales, we propose an adequately precise quantization method that incurs no additional dequantization overhead.
+Specifically, we introduce a per-thread quantization approach based on the mapping between the GPU thread and memory layout of matrices as dictated by the PTX mma instructions.
+This method groups tokens corresponding to the same thread for quantization and dequantization, ensuring that each thread is associated with a single quantization scale.
+This approach achieves much better accuracy performance than per-block quantization with no additional dequantization overhead.
+Second, for the significant channel-wise outliers in matrices $Q$ and $K$, we adopt smoothing $K$ in SageAttention and propose an effective method to remove these outliers in $Q$.
 Specifically, we propose subtracting the average value of the channel dimension of $Q$, referred to as $\overrightarrow{Q}_{m}$.
-Then, we add $\overrightarrow{Q}_{m}K$ after the $QK$ Matmul to ensure the correctness of the attention computation.
-Second, we observe that specific layers and timesteps consistently present quantization challenges across different inputs.
-To maintain accuracy, we apply an adaptive mix precision method.
-Specifically, we employ 8-bit (INT8+FP8) attention for these problematic layers and timesteps and 4-bit (INT4+FP8) attention for the others.
-Third, to mitigate the accuracy loss associated with using the 22-bit accumulator for FP8 Matmul of $PV$, we propose a method to smooth $V$ to enhance the accuracy performance.
+Then, we add $\overrightarrow{Q}_{m}K$ after the $QK^{\top}$ Matmul to ensure the correctness of the attention computation.
+To address (C2)—the accuracy loss from using a 22-bit accumulator for FP8 Matmul of $\widetilde{P}V$—, we propose to use an FP32 buffer to accumulate the values from the 22-bit accumulator after each block Matmul of $\widetilde{P}V$.
+This method confines the errors to the block range.
+Additionally, we design an optional technique to enhance the accuracy of the 22-bit accumulator.
+Specifically, we could smooth $V$ by subtracting the average value of its channel dimension and adding the subtracted item to the attention output to maintain the correctness.
 Performance.
 Importantly, we offer a high-performance implementation of SageAttention2 on RTX4090 and L20 GPUs.
-This implementation achieves a peak performance of 485 TOPS on the RTX4090, outperforming FlashAttention2 and xformers by approximately 3.1x and 5.4x, respectively.
+This implementation achieves a peak performance of 481 TOPS on the RTX4090, outperforming FlashAttention2 and xformers by approximately 3x and 5x, respectively.
+Note that FlashAttention3 is tailored to and can only be used with the Nvidia Hopper architecture, i.e., H100 and H800.
 We extensively evaluate the end-to-end metrics of state-of-the-art text, image, and video generation models using SageAttention2.
@@
 The matrices $Q$, $K$, and $V$ each have dimensions $N\times d$, while the matrix $S$, $P$ are $N\times N$.
-While $d$ is typically small, e.g., 64 or 128, $N$ can be thousands if not millions.
+While $d$ is typically small, e.g., 64 or 128, $N$ can be millions.
 Therefore, the $N\times N$ matrices $(S,P)$ are much larger than $(Q,K,V)$, and a naive implementation suffers from the huge amount of global memory IO for $(S,P)$ reads/writes.
-FlashAttention (Dao, 2023) proposes to tile $Q$, $K$, and $V$ from the token dimension into blocks $\{Q_{i}\},\{K_{i}\},\{V_{i}\}$ with block size of $b_{q}$, $b_{kv}$, $b_{kv}$, respectively.
+FlashAttention (Dao, 2023) proposes to tile $Q$, $K$, and $V$ from the token dimension into blocks $\{Q_{i}\},\{K_{i}\},\{V_{i}\}$ with block size of $b_{q}$, $b_{k}$, $b_{v}$, respectively, where $b_{k}$ = $b_{v}$.
 Then, to avoid the memory IO for $(S,P)$, it uses online softmax (Milakov & Gimelshein, 2018) to progressively compute each block of $O$, i.e., $O_{i}$:
 First, for each block of $\{K_{i}\},\{V_{i}\}$, it computes the following equations iteratively:
-$S^{j}_{i}=Q_{i}K_{j}^{\top}/\sqrt{d},~{}~{}(m^{j}_{i},\widetilde{P}^{j}_{i})=\tilde{\sigma}(m^{j-1}_{i},S^{j}_{i}),~{}~{}l_{i}^{j}=$ $\exp(m^{j}_{i}-m^{j-1}_{i})l_{i}^{j-1}+\mathrm{rowsum}(\widetilde{P}^{j}_{i}),$
-$O^{j}_{i}=\mathrm{diag}\left(\exp(m^{j}_{i}-m^{j-1}_{i})\right)O^{j-1}_{i}$ $+\widetilde{P}^{j}_{i}V_{j}$
+$S^{j}_{i}=Q_{i}K_{j}^{\top}/\sqrt{d},~{}~{}(m^{j}_{i},\widetilde{P}^{j}_{i})=\tilde{\sigma}(m^{j-1}_{i},S^{j}_{i}),~{}~{}l_{i}^{j}=$ $\exp(m^{j-1}_{i}-m^{j}_{i})l_{i}^{j-1}+\mathrm{rowsum}(\widetilde{P}^{j}_{i}),$
+$O^{j}_{i}=\mathrm{diag}\left(\exp(m^{j-1}_{i}-m^{j}_{i})\right)O^{j-1}_{i}$ $+\widetilde{P}^{j}_{i}V_{j}$
 Where $m_{i}^{j}$ and $l_{i}^{j}$ are $b_{q}\times 1$ vectors, which are initialized to $\infty$- ∞ and 00 respectively.
@@
 We should have $\psi^{-1}_{\delta_{A}}(\hat{A})\approx A$.
-The actual matrix multiplication $\hat{A}\hat{B}$ is carried in low-precision.
+The actual matrix multiplication $\hat{A}\hat{B}$ is carried in low precision.
 In modern GPUs, low-precision matrix multiplication is usually multiple times faster than higher-precision ones.
@@
 Also, per-channel quantizer assigns a scale factor for each channel of the tensor, i.e., along the channel dimension: $A[:,i]=\lceil A[:,i]/\delta_{A}[i]\rfloor,\delta_{A}[i]=\max(|A[:,i]|)/{7}$.
-Dequantization process is a element-wise scaling: $\psi_{\delta_{A}\delta_{B}}^{-1}(\hat{A}\hat{B})=\hat{A}\hat{B}*\delta_{A}*\delta_{B}$.
+The dequantization process is an element-wise scaling.
+For example, for per-tensor dequantization, $\psi_{\delta_{A}\delta_{B}}^{-1}(\hat{A}\hat{B})=\hat{A}\hat{B}*\delta_{A}*\delta_{B}$.
 2.3 SageAttention
@@
 Where $Q_{i},K_{j}$ are the tilled block in FlashAttention, and $\text{mean}(K)$ is the average value of the channel dimension of $K$.
-SageAttention keeps $\widetilde{P}V$ as FP16 and uses FP16 Matmul with an FP16 accumulator for $\widetilde{P}V$.
+Additionally, SageAttention keeps $\widetilde{P}V$ as FP16 and uses FP16 Matmul with an FP16 accumulator for $\widetilde{P}V$.
 However, FP16 Matmul with an FP16 accumulator only has a speedup effect on RTX4090 and RTX3090 GPUs.
 3 SageAttention-2
-Figure 3: Workflow of SageAttention2. 1 Smooth Q,K,V. 2 A GEMV to obtain $\Delta S$. 3 Per-warp quantize Q,K and per-channel quantize V. 4 Perform the SageAttention2 kernel. 5 Correct the output.
+Figure 3: Workflow of SageAttention2. 1 Smooth Q,K,V. 2 A GEMV to obtain $\Delta S$. 3 Per-thread quantize Q,K and per-channel quantize V. 4 Perform the SageAttention2 kernel. 5 Correct the output.
 3.1 Formulation
@@
 Figure 4: Typical examples of tensors’ data distribution in attention.
-Table 1: End-to-end metrics comparison of different quantization methods, where Q,K are quantized into INT4, while P,V stay in full precision.
-Q, K Smoothing (Q+K) Llama3.1 (Lambda) $\uparrow$ Llama3.1 (WikeText) $\downarrow$ CogVideo (vqa-a) $\uparrow$ CogVideo (vqa-t) $\uparrow$
+Table 1: End-to-end metrics comparison of different quantization methods, where $Q,K$ are quantized into INT4, while $\widetilde{P},V$ stay in full precision.
+Q, K Smoothing (Q+K) Llama3.1 (Lambda) $\uparrow$ Llama3.1 (WikeText) $\downarrow$ CogVideoX (vqa-a) $\uparrow$ CogVideoX (vqa-t) $\uparrow$
 Full-Precision - 81.5% 6.013 77.605 75.360
@@
 80.8% 6.219 77.276 75.147
-Table 2: Average accuracy across all layers using different quantization granularities.
-Method Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
-Per-token 99.45% 0.0649 0.0335
-Per-warp 99.45% 0.0648 0.0334
-Per-block 98.03% 0.1492 0.0744
-Per-tensor 97.15% 0.1800 0.0865
-Table 3: Worst accuracy across all layers using different quantization granularities.
-Method Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
-Per-token 96.76% 0.1916 0.0775
-Per-warp 96.71% 0.1956 0.0779
-Per-block 90.68% 0.3615 0.1490
-Per-tensor 85.85% 0.4687 0.2261
-3.2 Per-warp INT4 Quantization
-SageAttention uses per-block quantization, which quantizes each block of $Q$ and $K$ for a GPU streaming processor.
-Such a quantization strategy could achieve an accuracy performance close to per-token quantization and avoid the overhead of the dot product of the quantization scale vectors $\delta_{Q}$ and $\delta_{K}$.
-However, quantizing $Q$ and $K$ to INT4 demands a more accurate quantization granularity.
-We propose per-warp quantization, a more precise and granular quantization approach than the per-block quantizer, also with no additional overhead of vectors dot product.
-Specifically, each block of $Q$ in SageAttention will be split into $c_{w}$ segments, each processed by one of $c_{w}$ GPU warps in a GPU streaming processor (SM).
-Subsequently, each warp performs Matmul for its assigned segment.
-Per-warp INT4 quantization assigns a scale factor for every $b_{q}/{wc}$ tokens for $Q$:
-$\hat{Q}[\frac{i*b_{q}}{c_{w}}$ $:\frac{b_{q}*(i+1)}{c_{w}},:]=\left\lceil\frac{Q[\frac{i*b_{q}}{c_{w}}:\frac{b_{q}*(i+1)}{c_{w}},:]}{\delta_{Q}[i]}\right\rfloor$ (4)
-$\delta_{Q}[i]=\frac{\max(\left|\,Q[\frac{i*b_{q}}{c_{w}}:\frac{b_{q}*(i+1)}{c_{w}},:]\,\right|)}{7}$ (5)
-Per-warp quantization will provide an accuracy performance superior to per-block quantization used in SageAttention, which we will discuss in Section 3.5.
 Figure 5: An example of quantized value distribution of $Q$ before and after smoothing $Q$.
@@
 $\overrightarrow{Q_{m}}=\mathrm{mean}(Q),~{}~{}\gamma(Q)=$ $Q-\overrightarrow{Q_{m}},~{}~{}\gamma(K)=K-\mathrm{mean}(K),$
-$S=QK=$ $\gamma(Q)\gamma(K)+\overrightarrow{Q_{m}}\gamma(K)$
+$S=QK^{\top}=$ $\gamma(Q)\gamma(K)^{\top}+\overrightarrow{Q_{m}}\gamma(K)^{\top}$
 $\phi_{Q}(Q)=\psi_{Q}\circ\gamma(Q),~{}\overrightarrow{Q_{m}}$ $,~{}~{}~{}~{}~{}\phi_{K}(K)=\psi_{K}\circ\gamma(K)$ (4)
@@
 This $\Delta S$ will be added to $S$ during the attention computation.
-Finally, the transformation from full-precision $Q,K$ to quantized $\hat{Q},\hat{K}$ can be expressed as shown in Eqution 6, where $\psi_{Q}$, $\psi_{K}$ is two quantizers for $Q$ and $K$.
+Finally, the transformation from full-precision $Q,K$ to quantized $\hat{Q},\hat{K}$ can be expressed as shown in Eqution 4, where $\psi_{Q}$, $\psi_{K}$ is two quantizers for $Q$ and $K$.
 In other words, full-precision $Q,K$ are substracted with the means of the channel dimension before being quantized.
@@
 Also, Table 9 and Table 9 show that the order of effectiveness is smoothing Q+K $>$ smoothing Q $>$ smoothing K $>$ other baselines.
+Figure 6: An example of per-thread quantization.
+The left figure shows the correspondence between the quantization scales and the tokens in each GPU warp.
+The right figure shows the correspondence between quantization tokens and GPU threads in a MMA.m16n8k64 instruction, showing that each GPU thread only corresponds to one quantization scale in $\delta_{Q}$ and $\delta_{K}$ in dequantization.
+Table 2: Average accuracy across all layers of CogvideoX using different quantization granularities.
+Method Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
+Per-token 99.45% 0.0649 0.0335
+Per-thread 99.45% 0.0622 0.0313
+Per-block 98.03% 0.1492 0.0744
+Per-tensor 97.15% 0.1800 0.0865
+Table 3: Worst accuracy across all layers of CogvideoX using different quantization granularities.
+Method Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
+Per-token 96.76% 0.1916 0.0775
+Per-thread 96.72% 0.1932 0.0776
+Per-block 90.68% 0.3615 0.1490
+Per-tensor 85.85% 0.4687 0.2261
+3.3 INT4 Per-thread Quantization
+SageAttention uses per-block quantization, which quantizes each block of $Q$ and $K$ for a GPU streaming processor.
+Such a quantization strategy could achieve an accuracy performance close to per-token quantization and avoid the overhead of the dot product of the quantization scale vectors $\delta_{Q}$ and $\delta_{K}$.
+However, quantizing $Q$ and $K$ to INT4 demands a more accurate quantization granularity.
+We propose per-thread quantization, a more precise and granular quantization approach than the per-block quantizer, also without the additional overhead of vectors dot product required for per-token quantization.
+Specifically, each block of $Q$, i.e., $Q_{i}$, in SageAttention will be split into $c_{w}$ segments and processed by $c_{w}$ GPU warps in a GPU streaming processor (SM).
+We call each segment of $Q_{i}$ as $Q_{w}$, and $k_{w}=K_{j}$ because a $K_{j}$ only corresponds to one warp.
+Then, each warp containing 32 threads uses the mma.m16n8k64 PTX instruction (NVIDIA, ) for the $Q_{w}K_{j}^{\top}$.
+According to the layout requirement of this instruction, we find that the $Q_{w}[8\times(n\%8)]$ could share one quantization scale, and $K_{j}[8\times(n\%8)]$ and $K_{j}[8\times(n\%8+1)]$ could share one quantization scale, where $n$ is the token index.
+Such a quantization method is more fine-grained and with no additional overhead.
+This is because it assigns different GPU threads to distinct quantization groups based on the MMA instruction layout, with each thread performing dequantization only using a single quantization scale value.
+The per-thread quantization process can be formulated as follows.
+$i_{\delta q}=$ $\lfloor(n*8*c_{w}/b_{q})\rfloor$
+$q_{i}[i_{\delta q}]=$$\{8\times(n\%8)+\lfloor(n*c_{w}/b_{q})\rfloor*b_{q}/c_{w}\},~{}~{}n\in[0,N]$
+$\delta_{Q}[i_{\delta q}]=$ $\frac{\max(\left|\,Q[q_{i}[i_{\delta q}]]\,\right|)}{7}$
+$\hat{Q}[q_{i}[i_{\delta q}]]=$ $\left\lceil\frac{Q[q_{i}[i_{\delta q}]]}{\delta_{Q}[i_{\delta q}]}\right\rfloor$ (5)
+$i_{\delta k}=$ $\lfloor(n*4/b_{k})\rfloor$
+$k_{n}[i_{\delta k}]=$$\{8\times(n\%8)+\lfloor n/b_{k}\rfloor*b_{k}\}\cup\{8\times(n\%8)+1+\lfloor n/b_{k}\rfloor*b_{k}\},~{}~{}n\in[0,N]$
+$\delta_{K}[i_{\delta k}]=$ $\frac{\max(\left|\,K[k_{n}[i_{\delta k}]]\,\right|)}{7}$
+$\hat{K}[k_{n}[i_{\delta k}]]=$ $\left\lceil\frac{K[k_{n}[i_{\delta k}]]}{\delta_{K}[i_{\delta k}]}\right\rfloor$ (6)
+We show an example of per-thread quantization in Figure 6.
 Algorithm 1 Implementation of SageAttention2.
   Input: Matrices $Q(\text{FP16}),K(\text{FP16}),V(\text{FP16})\in\mathbb{R}^{N\times d}$, block size $b_{q},b_{kv}$, warp count $c_{w}$.
-  Preprocessing: $K=K-\mathrm{mean}(K)$, $\overrightarrow{Q}_{m}=\mathrm{mean}(Q)$, $Q=Q-\overrightarrow{Q}_{m}$, $\Delta S=\text{GEMV}(\overrightarrow{Q}_{m},K^{\top})$, $\overrightarrow{V}_{m}=\mathrm{mean}(V),V=V-\overrightarrow{V}_{m}$.
-  Quantization: ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{Q},\hat{Q})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{Q}}}(Q)$ // INT4 Per-warp.
-${}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{K},\hat{K})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{K}}}(K)$.
+  Preprocessing: $K=K-\mathrm{mean}(K)$, $\overrightarrow{Q}_{m}=\mathrm{mean}(Q)$, $Q=Q-\overrightarrow{Q}_{m}$, $\Delta S=\text{GEMV}(\overrightarrow{Q}_{m},K^{\top})$.
+  Quantization: ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{Q},\hat{Q})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{Q}}}(Q)$, ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{K},\hat{K})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{K}}}(K)$.
+//per-thread.
 ${}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{V},\hat{V})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{V}}}(V)$.
-// FP8 Per-channel.
+//per-channel.
   Divide $\hat{Q}$ to $T_{m}={N}/{b_{q}}$ blocks $\{\hat{Q}_{i}\}$; divide $\hat{K}$, $\hat{V}$, and $\Delta S$ to $T_{n}={N}/{b_{kv}}$ blocks $\{\hat{K}_{i}\}$, $\{\hat{V}_{i}\}$, and $\{\Delta S_{i}\}$;
   for $i=1$ to $T_{m}$ do
-     Load $\hat{Q}_{i}$ and $\delta_{Q}[i*c_{w}:(i+1)*c_{w}]$ into an SM ;
+     Load $\hat{Q}_{i}$ and $\delta_{Q}=\delta_{Q}[c_{w}*8i:c_{w}*8(i+1)]$ into an SM ;
      for j in [1, $T_{n}$do
-        Load $\hat{K}_{j}$, $\hat{V}_{j}$, and $\delta_{K}[j]$ into the SM ;
+        Load $\hat{K}_{j}$, $\hat{V}_{j}$, and $\delta_{K}=\delta_{K}[4j:4(j+1)]$ into the SM ;
         $w=\text{range}(c_{w}),st=w*c_{w}$
-        $S_{i}^{j}[st:st+c_{w}]=\mathrm{Matmul}(\hat{Q}_{i}[st:st+c_{w}],\hat{K}_{j}^{\top})\times\delta_{Q}[st+w]\times\delta_{K}[j]$ + $\Delta S_{j}$; // Paralleled by $c_{w}$ warps.
+        $S_{i}^{j}[st:st+c_{w}]=\psi^{-1}_{\delta_{Q}\delta_{K}}(\mathrm{Matmul}(\hat{Q}_{i}[st:st+c_{w}],\hat{K}_{j}^{\top}))+\Delta S_{j}$; // Paralleled by $c_{w}$ warps.
+The dequantization $\psi^{-1}_{\delta_{Q}\delta_{K}}$ is illustrated in Figure 6.
         $m_{i}^{j}=\mathrm{max}(m_{i}^{j-1},\mathrm{rowmax}(S_{i}^{j}))$, ${}\widetilde{P}_{i}^{j}=\mathrm{exp}(S_{i}^{j}-m_{i}^{j})$, ${}l_{i}^{j}=e^{m_{i}^{j-1}-m_{i}^{j}}+\mathrm{rowsum}(\widetilde{P}_{i}^{j})$ ;
-        $O_{i}^{j}=\mathrm{diag}(e^{m_{i}^{j-1}-m_{i}^{j}})^{-1}O_{i}^{j-1}+$ $\mathrm{Matmul}((\widetilde{P}_{i}^{j}*448)\text{.to(FP8.e4m3)},V_{j})$ ;
+        $O_{i}^{j}$(FP22) = $\mathrm{Matmul}((\widetilde{P}_{i}^{j}*448)\text{.to(FP8.e4m3)},V_{j})$ ;
+        $O_{i}^{j}$(FP32) = $\mathrm{diag}(e^{m_{i}^{j-1}-m_{i}^{j}})^{-1}O_{i}^{j-1}(FP32)+O_{i}^{j}$(FP22) ;
      end for
      Load $\delta_{V}$ into an SM;
-     $O_{i}=\mathrm{diag}(l_{i}^{T_{n}})^{-1}O_{i}^{T_{n}}/448*\delta_{V}$ ;
+[…]
      Write $O_{i}$ ;
   end for
-  $O=\{O_{i}\}$, $O=O+\overrightarrow{V}_{m}$.
+  $O=\{O_{i}\}$
   return $O$
+Table 4: Error of the FP8 Matmul instruction of mma(f8f8f32).
+Precision of Accumulated Value E8M13 E8M23
+Error compared to FP32 0 ||||mma(f32.f16.f16.f32) - mma(f32.f8.f8.f32)||||
+3.4 FP32 MMA Buffer for FP22 Accumulator
+To avoid the drawback of SageAttention where FP16 Matmul with an FP16 accumulator for $PV$ is only effective on GPUs such as the RTX4090, we adopt the approach of quantizing $P$ and $V$ to FP8 to leverage the universal acceleration of FP8 Tensor cores.
+However, we discover that the accumulator for the mma(f32f8f8f32) instruction on the Ada architecture is actually FP22, specifically with 1 sign bit, 8 exponent bits, and 13 mantissa bits.
+Specifically, for mma(f32f8f8f32) instruction $C=AB+D$, where $A,B$ are tensors in FP8 data type and $C,D$ are tensors with FP32 data type, we initialize the $A,B$ to zero and vary $D$ to test the data type of the accumulator.
+As shown in Table 4, when $D$ is initialized with 1 sign bit, 8 exponent bits, and 13 mantissa bits, the value of $C$ matches the result of the mma(f16f16f32) instruction.
+However, when $D$ is initialized with more than 13 mantissa bits, the error of $C$ corresponds to the difference between the results of mma(f32f16f16f32) and mma(f32f8f8f32).
+Consequently, the matrix multiplication of $\widetilde{P}V$, quantized to FP8, incurs a certain degree of accuracy loss compared to using an FP32 accumulator.
+To mitigate this accuracy loss as much as possible, we propose to use an FP32 buffer to accumulate the values of $\widetilde{P}_{i}^{j}V_{j}$ in FP22.
+Specifically, as shown in Algorithm 1, after each Matmul between the tiled $\widetilde{P}$ and $V$, we store the Matmul result in the FP32 buffer.
+This approach prevents the FP22 accumulator from accumulating errors across the entire sequence, confining errors to the tiled range instead.
 Table 5: Average accuracy using different data types of $(\widetilde{P},V)$ across all layers of a CogvideoX model, where $(Q,K)$ are smoothed.
@@
 FP16 96.76% 0.1916 0.0775
-Table 6: An accuracy example on real tensors of CogvideoX model with or without smoothing $V$.
-Smooth V Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
-98.25% 0.1980 0.2387
-99.75% 0.0406 0.0773
-Figure 6: An example of dot product precison a row of $\widetilde{P}$ and a column of $V$ presented by FP22 data type.
-Table 7: Error of the FP8 Matmul instruction of mma(f8f8f32).
-Precision of Accumulated Value E8M13 E8M23
-Error compared to FP32 0 ||||mma(f32.f16.f16.f32) - mma(f32.f8.f8.f32)||||
-3.4 Smooth V
-To avoid the drawback of SageAttention where FP16 Matmul with an FP16 accumulator for $PV$ is only effective on GPUs such as the RTX4090, we adopt the approach of quantizing $P$ and $V$ to FP8 to leverage the universal acceleration of FP8 Tensor cores.
-However, we discover that the accumulator for the mma(f32f8f8f32) instruction on the Ada architecture is actually FP22, specifically with 1 sign bit, 8 exponent bits, and 13 mantissa bits.
-Specifically, for mma(f32f8f8f32) instruction $C=AB+D$, where $A,B$ are tensors in FP8 data type and $C,D$ are tensors with FP32 data type, we initialize the $A,B$ to zero and vary $D$ to test the data type of the accumulator.
-As shown in Table 7, when $D$ is initialized with 1 sign bit, 8 exponent bits, and 13 mantissa bits, the value of $C$ matches the result of the mma(f16f16f32) instruction.
-However, when $D$ is initialized with more than 13 mantissa bits, the error of $C$ corresponds to the difference between the results of mma(f32f16f16f32) and mma(f32f8f8f32).
-Consequently, matrix multiplication of matrices $\widetilde{P}$ and $V$, quantized to FP8, incurs a certain degree of precision loss compared to using an FP32 accumulator.
-To mitigate this precision loss as much as possible, we propose to smooth $V$:
-$\gamma(V)=V-\mathrm{mean}(V),$ $~{}~{}~{}~{}\mathrm{mean}(V)=\overrightarrow{V_{m}}$
-$\phi_{V}(V)=$ $\psi_{V}\circ\gamma,~{}\overrightarrow{V_{m}}$ (7)
-As shown in Figure 6, this strategy enhances the accuracy of FP22 for values in $\widetilde{P}V$ for the following reasons: Each row of $\widetilde{P}$ spans a value range from 0 to 1, and each column of $V$ consistently features channel-wise outliers that are exclusively positive or negative, typically ranging between 8 and 9.
-Consequently, the values of $\widetilde{P}V$ could be quite large.
-However, the floating-point number representation range is not uniform—it is denser near zero.
-Therefore, by subtracting the mean along the channel dimension from $V$, the values of $\widetilde{P}V$ will be closer to zero, resulting in higher representational precision.
-We believe such an explanation and strategy could address lots of issues [^1] https://github.com/triton-lang/triton/issues/4476, https://github.com/triton-lang/triton/issues/5065 reported by the community.
-Additionally, to maintain the correctness of the attention computation, it is only necessary to add $\overrightarrow{V_{m}}$ to the final calculation of $O$: $O=O+\overrightarrow{V_{m}}$.
-This is because the sum of each row of the $\widetilde{P}$ matrix equals 1, so $\widetilde{P}\overrightarrow{V_{m}}=\overrightarrow{V_{m}}$.
-In other words, this method decomposes $V$ into two parts: $\overrightarrow{V_{m}}$ and $V$.
-For $V$, it centers the values of each column around zero, which leads to the dot product result of a row from the quantized $\widetilde{P}$ matrix with a column from the quantized $V$ matrix being closer to zero.
-This makes the representation of FP22 more accurate.
-Meanwhile, $\overrightarrow{V_{m}}$ is retained in FP16 and added to $O$ at the end without causing a loss of precision for the $\overrightarrow{V_{m}}$ part.
-Table 6 shows the attention accuracy on real tensors sampled from CogvideoX with and without smoothing $V$.
-It demonstrates that smoothing V could improve the accuracy of SageAttention2 when quantizing $Q,K$ to INT4 and $\widetilde{P},V$ to FP8.
+Table 7: Two kernel implementations of SageAttention2.
+      Kernel       $\psi_{Q}(Q)$, $\psi_{K}(K)$       $\psi_{P}(\widetilde{P})$, $\psi_{V}(V)$
+      SageAttn2-4b       INT4, per-thread       FP8, per-block;  FP8, per-channel
+      SageAttn2-8b       INT8, per-thread       FP8, per-block;  FP8, per-channel
 3.5 Quantization for Q, K, P, V
 Quantization for $Q,K$.
-We propose $\psi_{Q}(Q)$ and $\psi_{K}(K)$ in granularity of per-warp.
+We propose $\psi_{Q}(Q)$ and $\psi_{K}(K)$ in granularity of per-thread.
 This is first because per-channel quantization is not feasible since the scale factors of the inner axis of $QK^{\top}$ cannot used to do dequantization (Xiao et al., 2023a).
-Second, as shown in Table 3 and Table 3, we compare the average and worst-case accuracies of INT4 quantization at per-token, per-warp, per-block, and per-tensor granularity using real $Q,K,V$ across all layers of CogvideoX.
-Results indicate that the accuracy of per-warp quantization is very close to per-token and outperforms a lot than per-block and per-tensor.
-Furthermore, per-warp quantization incurs less dequantization overhead than per-token as discussed in Section 3.2.
+Second, as shown in Table 3 and Table 3, we compare the average and worst-case accuracies of INT4 quantization at per-token, per-thread, per-block, and per-tensor granularity using real $Q,K,V$ across all layers of CogvideoX.
+Results indicate that the accuracy of per-thread quantization is very close to per-token and outperforms a lot more than per-block and per-tensor.
+Furthermore, per-thread quantization incurs less dequantization overhead than per-token, as discussed in Section 3.3.
 Quantization for $\widetilde{P},V$.
@@
 Second, $\widetilde{P}=\mathrm{exp}(S_{i}-\mathrm{rowmax}(S_{i}))$, where $S_{i}$ is the Matmul result of a block of $Q$ and $K^{\top}$, the max value in each row of $\widetilde{P}$ is 1.
-Hence, we can assign a single static scale $s=\frac{1}{448}$ to a block $\widetilde{P}$, whose accuracy equals per-token quantization.
+Hence, we can assign a single static scale $s=\frac{1}{448}$ to a block $\widetilde{P}$.
 Third, per-channel quantization can address the channel-wised outlier of $V$.
@@
 Then, Cosine similarity: $CosSim=\sum OO^{\prime}/\smash{\sqrt{\sum O^{2}}}\smash{\sqrt{\sum O^{\prime 2}}}$, Relative L1 distance: $L1=\sum|O-O^{\prime}|/\sum|O|$, Root mean square error: $RMSE=\sqrt{(1/n)\sum(O-O^{\prime})^{2}}$.
-Table 8: Two kernel implementations of SageAttention2.
-     Kernel      $\psi_{Q}(Q)$, $\psi_{K}(K)$      $\psi_{P}(\widetilde{P})$, $\psi_{V}(V)$
-     SageAttn2-4b      INT4, per-warp      FP8, per-block;  FP8, per-channel
-     SageAttn2-8b      INT8, per-warp      FP8, per-block;  FP8, per-channel
-Figure 7: Mean and standard deviation of $cossim*(1-L1)$) of SageAttn-4b in different layers and timesteps for different inputs in Llama3.1 and CogvideoX.
-3.6 Adaptive Quantization over Layer and Timestep
-Based on the discussion in Section 3.5, we implement two attention kernels (See Table 8) based on a choice: Using INT4 or INT8 quantization for $\psi_{Q}(Q)$ and $\psi_{K}(K)$.
-The speed of these kernels is in the order (SageAttn2-4b$>$SageAttn2-8b), but the accuracy order is opposite.
-Although SageAttn2-4b could perform adequately in lots of tasks, for instance, it shows very slight end-to-end metric loss of Lambda benchmark on Llama3.1, SageAttn2-4b falls short in some hard scenarios.
-Consequently, we first conduct a detailed analysis of the accuracy performance on various models using SageAttn2-4b and SageAttn2-8b.
-Then, we propose an effective and easy-to-use adaptive method.
-Accuracy across layers and timesteps.
-Figure 7 displays the mean and standard deviation of the accuracy, calculated using $CosSim*(1-L1)$, for SageAttn-4b across different layers and timesteps of different inputs in Llama3.1 and CogvideoX.
-It can be observed that specific layers and timesteps exhibit specific ranges of errors.
-Our method.
-Our adaptive strategy is detailed as follows: First, we assess the average accuracy of SageAttn2-4b in comparison to full precision attention for each (layer, timestep) combination within a model, using several random inputs.
-For models without a timestep dimension, such as language models, the evaluation is confined to the layer dimension.
-Second, we will sort the (layer, timestep) combinations in descending order based on their $CosSim*(1-L1)$ values.
-We then identify the smallest $\xi$of these combinations.
-Finally, for the identified (layer, timestep) combinations, we consistently apply SageAttn2-8b across all other prompts.
-We call such a strategy SageAttn-mix and will report the values of $\xi$we used for each model in Section 4.1.
+Figure 7: An example of dot product precison a row of $\widetilde{P}$ and a column of $V$ presented by FP22 data type.
+3.6 An Optional Technique: Smooth V
+We also figure out another way to mitigate the accuracy loss due to the FP22 accumulator when $V$ possesses channel-wise biases:
+$\gamma(V)=V-\mathrm{mean}(V),$ $~{}~{}~{}~{}\mathrm{mean}(V)=\overrightarrow{V_{m}}$
+$\phi_{V}(V)=$ $\psi_{V}\circ\gamma,~{}\overrightarrow{V_{m}}$ (7)
+As shown in Figure 7, this strategy could enhance the accuracy of FP22 for values in $\widetilde{P}V$ for the following reasons: Each row of $\widetilde{P}$ spans a value range from 0 to 1, and each column of $V$ consistently features channel-wise biases that are exclusively positive or negative, for instance, ranging between 8 and 9.
+Consequently, the values of $\widetilde{P}V$ could be quite large.
+However, the floating-point number representation range is not uniform—it is denser near zero.
+Therefore, by subtracting the mean along the channel dimension from $V$, the values of $\widetilde{P}V$ will be closer to zero, resulting in higher representational precision.
+We believe such an explanation and strategy could address lots of issues [^1] https://github.com/triton-lang/triton/issues/4476, https://github.com/triton-lang/triton/issues/5065 reported by the community.
+Additionally, to maintain the correctness of the attention computation, it is only necessary to add $\overrightarrow{V_{m}}$ to the final calculation of $O$: $O=O+\overrightarrow{V_{m}}$.
+This is because the sum of each row of the $\widetilde{P}$ matrix equals 1, so $\widetilde{P}\overrightarrow{V_{m}}=\overrightarrow{V_{m}}$.
+In other words, this method decomposes $V$ into two parts: $\overrightarrow{V_{m}}$ and $V$.
+For $V$, it centers the values of each column around zero, which leads to the dot product result between a row from the quantized $\widetilde{P}$ matrix and a column from the quantized $V$ matrix being closer to zero.
+This makes the representation of FP22 more accurate.
+Meanwhile, $\overrightarrow{V_{m}}$ is retained in FP16 and added to $O$ at the end without causing a loss of precision for the $\overrightarrow{V_{m}}$ part.
+Remark.
+This technique is optional and not employed in our main experiments.
+This is because its benefits are marginal, only significantly enhancing attention accuracy primarily when $V$ exhibits channel-wise outliers, which are absent in some models, such as $V$ in Llama3.1 (see Figure 4).
 Figure 8: Speed comparison between SageAttention2 and baselines (RTX4090, headdim=64).
 Figure 9: Speed comparison between SageAttention2 and baselines (RTX4090, headdim=128).
-Figure 10: Speed comparison between SageAttention2 and baselines (RTX4090, headdim=256).
 Figure 10: Speed comparison between SageAttention2 and baselines (L20, headdim=64).
 Figure 11: Speed comparison between SageAttention2 and baselines (L20, headdim=128).
-Figure 13: Speed comparison between SageAttention2 and baselines (L20, headdim=256).
 4 Experiment
@@
 We validate the effectiveness of SageAttention2 across a diverse set of representative models from language, image, and video generation.
-Specifically, we conduct experiments on five models: Llama2 (7B) (Touvron et al., 2023), Llama3.1 (8B) (Dubey et al., 2024), and GLM4 (9B) (GLM et al., 2024) for text2text, CogvideoX (2B) (Yang et al., 2024) and Open-Sora (Zheng et al., 2024) for text2video, Flux (schnell) (Black Forest Labs, 2023) for text2image, and TIMM (Wightman, 2019) for image classification.
+Specifically, we conduct experiments on ten models: Llama2 (7B) (Touvron et al., 2023), Llama3.1 (8B) (Dubey et al., 2024), and GLM4 (9B) (GLM et al., 2024) for text2text, CogvideoX (2B), CogvideoX (1.5-5B) (Yang et al., 2024), HunyuanVideo (Weijie Kong & Jie Jiang, 2024), and Mochi (Team, 2024) for text2video, Flux (schnell) (Black Forest Labs, 2023) and Stable-Diffusion3.5 (turbo) (Stability AI, 2023) for text2image, and TIMM (Wightman, 2019) for image classification.
 Datasets.
@@
 Text-to-video models are evaluated using the open-sora (Zheng et al., 2024) prompt sets.
-Flux is assessed on MJHQ-30K (Li et al., 2024).
+Text-to-image models are assessed on MJHQ-30K (Li et al., 2024).
 TIMM is evaluated on on three image datasets: ImageNet (Deng et al., 2009), ImageNet-Sketch (Sketch) (Wang et al., 2019), and ImageNet-Rendition (ImageNet-r) (Hendrycks et al., 2021).
@@
 For text-to-image models, generated images are compared with the images in MJHQ-30K dataset in three aspects: FID (Heusel et al., 2017) and sFID (Salimans et al., 2016) for fidelity evaluation, Clipscore (CLIP) (Hessel et al., 2021) for text-image alignment, and ImageReward (IR) (Xu et al., 2024) for human preference.
-For TIMM, we use Accuracy.
+For TIMM, we use classification accuracy.
 Implemetation details. We implement our attention kernels using CUDA and conduct experiments on Ubuntu 22.04 servers with RTX4090 and L20 GPUs.
@@
 Following Quarot (Ashkboos et al., 2024), we apply random Hadamard transformation for matrixes $Q,K$ before INT4 quantization.
-Hyper-parameters of SageAttn-mix. We use $\xi$ of 50% for Llama2, 60% for CogvideoX, 25% for Open-Sora, 30% for Llama3.1 and GLM4, 0% for Flux, and TIMM.
-Figure 14: Comparison examples from CogvideoX, prompts are sampled from open-sora prompt sets.
-Table 9: Average accuracy across all layers using different smooth methods.
+(3) SageAttention (Zhang et al., 2024) uses smoothing $K$, INT8 per-block quantization for $Q,K$, and FP16 for $\widetilde{P},V$.
+Figure 12: Comparison examples from HunyuanVideo, prompts are sampled from open-sora prompt sets.
+Figure 13: Comparison examples from CogvideoX (2B), prompts are sampled from open-sora prompt sets.
+Table 8: Average accuracy across all layers of CogvideoX using different smooth methods.
 Method CosSim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
@@
 SageAttn2-4b 99.46% 0.0648 0.0334
-Table 10: Worst accuracy across all layers using different smooth methods.
+Table 9: Worst accuracy across all layers of CogvideoX using different smooth methods.
 Method CosSim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
@@
 SageAttn2-4b 96.71% 0.1956 0.0779
-Table 11: End-to-end metrics loss across text, image, and video generation models.
+Table 10: End-to-end metrics across text and video generation models. - means the model does not support this dataset evaluation. ✗ indicates an inability to generate results for evaluation.
 Model Attention WikiText (Ppl.) $\downarrow$ Lambda (Acc.) $\uparrow$ MMLU (Acc.) $\uparrow$ Longbench $\uparrow$
 Llama2 Full-Precision 5.823 0.886 0.439 -
-HadmdAttn 6.706 0.865 0.355 -
-SmoothAttn 6.690 0.871 0.395 -
-SageAttn2-4b 6.018 0.886 0.436 -
-SageAttn2-mix 5.883 0.883 0.431 -
+HadmdAttn 6.771 0.860 0.360 -
+SmoothAttn 6.717 0.867 0.392 -
+SageAttn2-4b 5.912 0.881 0.428 -
+SageAttn2-8b 5.828 0.886 0.438 -
+SageAttention 5.824 0.887 0.439 -
 Llama3.1 Full-Precision 6.013 0.815 0.635 49.40
-HadmdAttn 7.661 0.756 0.502 44.62
-SmoothAttn 7.087 0.788 0.551 43.76
-SageAttn2-4b 6.219 0.808 0.617 48.61
-SageAttn2-mix 6.131 0.816 0.629 49.01
+HadmdAttn 7.872 0.762 0.500 44.07
+SmoothAttn 7.180 0.783 0.541 44.69
+SageAttn2-4b 6.256 0.798 0.607 48.79
+SageAttn2-8b 6.019 0.811 0.634 49.59
+SageAttention 6.017 0.812 0.634 49.55
 GLM4 Full-Precision 7.241 0.432 0.743 49.78
-HadmdAttn 7.932 0.435 0.676 46.27
-SmoothAttn 8.841 0.442 0.599 43.10
-SageAttn2-4b 7.341 0.435 0.732 49.06
-SageAttn2-mix 7.303 0.434 0.737 49.77
+HadmdAttn 7.989 0.435 0.669 45.97
+SmoothAttn 8.943 0.449 0.592 42.20
+SageAttn2-4b 7.352 0.433 0.725 49.23
+SageAttn2-8b 7.242 0.432 0.745 49.60
+SageAttention 7.243 0.433 0.744 49.79
 Model Attention CLIPSIM $\uparrow$ CLIP-T $\uparrow$ VQA-a $\uparrow$ VQA-t $\uparrow$ FScore $\uparrow$
-CogvideoX Full-Precision 0.1836 0.9975 77.605 75.360 3.006
-HadmdAttn 0.1742 0.9877 36.028 23.786 0.550
-SmoothAttn 0.1763 0.9870 37.444 42.184 0.792
-SageAttn2-4b 0.1813 0.9969 77.276 75.147 2.070
-SageAttn2-mix 0.1816 0.9976 75.686 78.600 2.884
-Open-Sora Full-Precision 0.1831 0.9996 46.713 59.553 0.368
-SageAttn2-4b 0.1821 0.9994 42.270 55.965 0.364
-SageAttn2-mix 0.1814 0.9994 44.509 59.097 0.383
+CogvideoX (2B) Full-Precision 0.1836 0.9975 77.605 75.360 3.006
+HadmdAttn 0.1742 0.9877 29.780 23.985 0.499
+SmoothAttn 0.1741 0.9870 41.703 47.043 0.624
+SageAttn2-4b 0.1821 0.9973 77.368 74.906 2.603
+SageAttn2-8b 0.1829 0.9977 76.532 74.281 2.941
+SageAttention 0.1833 0.9976 76.997 71.360 2.988
+CogvideoX (1.5-5B) Full-Precision 0.1778 0.9979 70.231 70.928 2.507
+HadmdAttn 0.1576 0.9933 8.990 2.299
+SmoothAttn 0.1559 0.9950 8.812 2.277
+SageAttn2-4b 0.1721 0.9978 57.729 52.989 2.884
+SageAttn2-8b 0.1775 0.9980 69.492 74.415 2.487
+SageAttention
+HunyuanVideo Full-Precision 0.1783 0.9995 82.516 75.934 0.604
+HadmdAttn 0.1727 0.9989 7.514 0.762 0.175
+SmoothAttn 0.1739 0.9988 6.987 0.609 0.148
+SageAttn2-4b 0.1751 0.9995 81.478 65.371 0.610
+SageAttn2-8b 0.1782 0.9996 81.786 75.354 0.586
+SageAttention 0.1786 0.9995 82.496 79.843 0.597
+Mochi Full-Precision 0.1798 0.9986 45.549 65.416 1.266
+HadmdAttn 0.1733 0.9980 9.053 25.133 0.704
+SmoothAttn 0.1687 0.9978 3.383 3.480 0.241
+SageAttn2-4b 0.1783 0.9986 35.955 43.735 1.137
+SageAttn2-8b 0.1797 0.9986 46.760 64.901 1.255
+SageAttention 0.1800 0.9987 48.707 63.763 1.269
+Table 11: Table 10 continued. End-to-end metrics across image generation and classification models.
         Model         Attention        FID $\downarrow$        sFID $\downarrow$        CLIP $\uparrow$        IR $\uparrow$
-        Flux         Full-Precision        11.303        17.603        32.603        0.9745
-        HadmdAttn        11.163        17.693        32.592        0.9638
-        SmoothAttn        10.941        18.098        32.582        0.9613
-        SageAttn2-4b        10.563        17.052        32.631        0.9747
+        Flux         Full-Precision        10.960        16.648        26.180        1.009
+        HadmdAttn        11.353        18.495        26.123        0.965
+        SmoothAttn        11.149        19.017        26.109        0.959
+        SageAttn2-4b        10.577        17.497        26.141        0.998
+        SageAttn2-8b        10.927        16.723        26.175        1.009
+        SageAttention        10.944        16.641        26.171        1.008
+               Stable-Dif        fusion3.5         Full-Precision        14.105        15.646        25.505        0.902
+        HadmdAttn        14.259        15.909        25.513        0.886
+        SmoothAttn        14.161        15.649        25.510        0.887
+        SageAttn2-4b        14.097        15.397        25.487        0.895
+        SageAttn2-8b        14.106        15.647        25.499        0.901
+        SageAttention        14.140        15.678        25.503        0.902
 Model Attention ImageNet (Acc.) $\uparrow$ Sketch (Acc.) $\uparrow$ ImageNet-r (Acc.) $\uparrow$
 TIMM Full-Precision 84.79% 45.32% 59.55%
-SageAttn2-4b 84.67% 45.07% 59.11%
+HadmdAttn 84.50% 44.89% 58.80%
+SmoothAttn 84.40% 44.68% 58.73%
+SageAttn2-4b 86.67% 45.24% 59.29%
+SageAttn2-8b 84.79% 45.39% 59.57%
+SageAttention 84.74% 45.38% 59.95%
 4.2 Speed and Accuracy of Kernels
 Speed.
-We conduct experiments to compare the Speed of SageAttention2 against baselines using configurations with headdim=64, headdim=128, and headdim=256, both with and without Causal Mask (Vaswani, 2017).
-Specifically, Figure 8, Figure 9, and Figure 10 show the speed of SageAttention2 and baselines across varying sequence lengths on RTX4090.
-These results indicate that SageAttention2 achieves a peak of 485 TOPS and is 3.1x faster than FlashAttention2 and 5.4x faster than xformers.
-Figure 11, Figure 12, and Figure 13 illustrate the results on L20 GPU, indicating that SageAttention2 achieves a peak of 288 TOPS and is 2.7x faster than FlashAttention2 and 4.6x faster than xformers.
+We conduct experiments to compare the Speed of SageAttention2 against baselines using configurations with headdim=64 and headdim=128, both with and without Causal Mask (Vaswani, 2017).
+Specifically, Figure 8 and Figure 9, show the speed of SageAttention2 and baselines across varying sequence lengths on RTX4090.
+These results indicate that SageAttn2-4b and SageAttn2-8b are approximately 3x and 2.7x faster than FlashAttention2, and about 4.5x and 4x faster than xformers, respectively.
 Accuracy.
 Table 9 and Table 9 show that the average and worst accuracy of different methods with INT4 $Q,K$ and FP8 $P,V$ across all layers of CogvideoX.
-The results indicate that the accuracy of SageAttn-4b is superior to other baselines.
-Figure 15: End-to-end performance using different ratio of SageAttn-8b.
+The results indicate that the accuracy of SageAttn2-4b is superior to other baselines.
+Table 12: End-to-end generation latency using SageAttention2 (The latency of Llama3.1 is the time to first token generation using different sequence lengths).
+Model GPU Original SageAttn2-8b SageAttn2-4b
+CogvideoX (2B) RTX4090 86 s 54 s 52 s
+CogvideoX (1.5-5B) RTX4090 1040 s 577 s 555 s
+HunyuanVideo L20 2221 s 1486 s 1435 s
+Mochi L20 2336 s 1316 s 1190 s
+Llama3.1 (48K tokens) RTX4090 9.2 s 5.7 s 5.6 s
+Llama3.1 (100K tokens) L20 39.9 s 25.4 s 23.2 s
 4.3 End-to-end Performance
@@
 We assessed the end-to-end metrics of various models using SageAttention2 compared to using attention in full precision.
-Detailed evaluation results are presented in Table 11 for Llama2, Llama3.1, GLM4, CogvideoX, Open-Sora, Flux, and TIMM, respectively.
-The results indicate that SageAttn-4b outperforms all baselines and maintains most of the end-to-end accuracy across all models.
-Additionally, using adaptive quantization techniques, SageAttn-mix achieves performance comparable to full-precision attention.
+Detailed evaluation results are presented in Table 10 for Llama2, Llama3.1, GLM4, CogvideoX (2B), CogvideoX (1.5-5B), HunyuanVideo, Mochi, Flux, Stable-Diffusion3.5, and TIMM, respectively.
+The results indicate that SageAttn2-4b outperforms all baselines and maintains most of the end-to-end accuracy across all models.
+Additionally, SageAttn2-8b incurs almost no metrics loss across various models.
+Visible examples.
+Figure12 and Figure 13 show some visible comparision examples from HunyuanVideo and CogvideoX (1.5-5B).
+We can observe that SageAttn2-8b does not introduce any visible differences compared to full-precision attention, whereas SageAttn2-4b has minor differences compared to full-precision attention but is much better than the baselines.
+End-to-end speedup.
+We compared the original generation latency and the latency using SageAttention2 for models with long sequence lengths in Table 12, observing significant speedup effects.
+For instance, SageAttention2 achieved a 1.8x speedup in CogvideoX (1.5-5B) without any metrics loss (SageAttn2-8b).
+SageAttn2-4b further accelerated these models but with a little metrics loss.
+Table 13: An accuracy example on real tensors of CogvideoX model with or without smoothing $V$.
+Smooth V Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
+98.25% 0.1980 0.2387
+99.75% 0.0406 0.0773
 4.4 Ablation Study
-Adaptive quaitzation.
-To analyze the impact of using different ratio of SageAttn-8b in adaptive quantization, Table 15 shows the changes in the perplexity of Llama3.1 under various ratio of SageAttn-8b.
-It can be observed that even if SageAttn-4b is used exclusively, the overall end-to-end representation is sufficiently good, and the higher the ratio of SageAttn-8b used, the closer the accuracy approaches that of using full-precision attention.
 Overhead of smoothing Q, K.
 Note that the overhead of smoothing $Q,K$ only contains $Q-\overrightarrow{Q_{m}}$, $K-\mathrm{mean}(K)$, and $\overrightarrow{Q_{m}}K^{\top}$.
-The speed overhead of them accounts for about 3.5% of the attention kernel.
+The speed overhead of smoothing $Q,K$ accounts for about 3.5% of the attention kernel.
+Benefit of smoothing V.
+Table 13 shows the attention accuracy on real tensors sampled from CogvideoX with and without smoothing $V$.
+It demonstrates that smoothing V could improve the accuracy of SageAttention2 when quantizing $Q,K$ to INT4 and $\widetilde{P},V$ to FP8.
 5 Conclusion and Future Work
-We introduce SageAttention2, an efficient and accurate 4-bit quantization method for attention.
-First, we propose to quantize $(Q,K)$ to INT4 in a warp-level granularity and quantize $(\widetilde{P},V)$ to FP8.
-Second, we propose a method to smooth matrixes $Q$ and $V$, enhancing the accuracy of attention with INT4 $QK$ and FP8 $PV$.
-Third, we analyze the quantization accuracy across timesteps and layers, then propose an adaptive quantization method to ensure the end-to-end metrics over various models.
-Our implementation is faster than FlashAttention2 and xformers by approximately 3.1x and 5.4x on RTX4090, respectively.
+We introduce SageAttention2, an efficient and accurate quantization method for attention.
+First, we propose to quantize matrixes $(Q,K)$ to INT4 in a thread-level granularity and quantize matrixes $(\widetilde{P},V)$ to FP8.
+Second, we propose a method to smooth $Q$, enhancing the accuracy of INT4 $QK^{\top}$.
+Third, we propose to use an FP32 Matmul buffer for $PV$ to enhance the accuracy of FP8 $\widetilde{P}V$.
+Our implementation is faster than FlashAttention2 and xformers by approximately 3x and 5x on RTX4090, respectively.
 Extensive testing confirms that our approach maintains end-to-end metrics across various models, including language, image, and video generation.

### v4

@@
 To further enhance the efficiency of attention computation compared to SageAttention while maintaining precision, we propose SageAttention2, which utilizes significantly faster 4-bit matrix multiplication (Matmul) alongside additional precision-enhancing techniques.
-First, we propose to quantize matrixes $(Q,K)$ to INT4 in a hardware-friendly thread-level granularity and quantize matrixes $(\widetilde{P},V)$ to FP8.
+First, we propose to quantize matrices $(Q,K)$ to INT4 in a hardware-friendly thread-level granularity and quantize matrices $(\widetilde{P},V)$ to FP8.
 Second, we propose a method to smooth $Q$, enhancing the accuracy of INT4 $QK^{\top}$.
-Third, we propose to use an FP32 Matmul buffer for $PV$ to enhance the accuracy of FP8 $\widetilde{P}V$.
-The operations per second (OPS) of SageAttention2 surpass FlashAttention2 and xformers by about 3x and 5x on RTX4090, respectively.
-Comprehensive experiments confirm that our approach incurs negligible end-to-end metrics loss across diverse models—including those for large language processing, image generation, and video generation.
-The codes are available at https://github.com/thu-ml/SageAttention.
-footnotetext: * Equal Contribution.
+Third, we propose a two-level accumulation strategy for $\widetilde{P}V$ to enhance the accuracy of FP8 $\widetilde{P}V$.
+The operations per second (OPS) of SageAttention2 surpass FlashAttention2 and xformers by about 3x and 4.5x on RTX4090, respectively.
+Moreover, SageAttention2 matches the speed of FlashAttention3(fp8) on the Hopper GPUs, while delivering much higher accuracy.
+Comprehensive experiments confirm that our approach incurs negligible end-to-end metrics loss across diverse models—including those for language, image, and video generation.
+The code is available at https://github.com/thu-ml/SageAttention
+Machine Learning, ICML
 Figure 1: The upper left figure shows the kernel speedup on RTX4090 GPU.
-The upper right figure shows the end-to-end inference speedup of generating the first token and performance metrics for the needle-in-a-haystack task (gkamradt, 2023) with a sequence length of 100K on Llama3.1 on L20 GPU.
-The figure below shows two videos generated by CogvideoX (1.5-5B) using FlashAttention2 and SageAttention2 on RTX4090, showing that SageAttention2 accelerates generation 1.8x with no video quality loss.
+The upper right figure shows the end-to-end inference speedup of generating the first token and performance metrics for the Needle-in-a-Haystack task (Kamradt, 2023) with a sequence length of 100K on Llama3.1 on L20 GPU.
+The figure below shows two videos generated by CogvideoX (1.5-5B) using FlashAttention2 and SageAttention2 on RTX4090, showing that SageAttention2 accelerates generation by 1.8x with no video quality loss.
 1 Introduction
-Due to the quadratic time complexity of attention, efficiently implementing it becomes crucial as sequence lengths increase in real-world applications (Jiang et al., 2024).
-Several strategies have been developed to mitigate computational demands of attention —such as (1) Linear attention methods (Wang et al., 2020; Choromanski et al., 2020; Yu et al., 2022; Katharopoulos et al., 2020) that reduce complexity to $O(N)$ and (2) Sparse attention methods (Liu et al., 2021; Chu et al., 2021; Li et al., ; Xiao et al., 2023b, 2024; Chen et al., 2023; Jiang et al., 2024; Venkataramanan et al., 2023; Gao et al., 2024; Fu et al., 2024) that selectively process parts of the context — these methods are only suitable for a limited range of models and tasks.
-Widely used attention methods optimize attention by exploiting hardware capacities to enhance the computation speed, such as FlashAttention (Dao et al., 2022), FlashAttention2 (Dao, 2023), FlashAttention3 (Shah et al., 2024), xformers (Lefaudeux et al., 2022), and SageAttention (Zhang et al., 2024).
-Those works do not omit the computation of attention over parts of the context and achieve impressive speed and precision performance across various models and tasks.
+Due to the quadratic time complexity of attention, its efficient implementation becomes crucial as sequence lengths increase in real-world applications (Jiang et al., 2024).
+Several strategies have been developed to mitigate computational demands of attention —such as (1) Linear attention methods (Wang et al., 2020; Choromanski et al., 2021; Yu et al., 2022; Katharopoulos et al., 2020) that reduce complexity to $O(N)$ and (2) Sparse attention methods (Liu et al., 2021; Chu et al., 2021; Li et al., 2022; Xiao et al., 2024b, a; Chen et al., 2024; Jiang et al., 2024; Venkataramanan et al., 2024; Gao et al., 2024; Fu et al., 2024) that selectively process parts of the context — these methods are only suitable for a limited range of models and tasks.
+The widely used attention methods optimize attention by exploiting hardware capacities, such as FlashAttention (Dao et al., 2022), FlashAttention2 (Dao, 2024), FlashAttention3 (Shah et al., 2024), xformers (Lefaudeux et al., 2022), and SageAttention (Zhang et al., 2025).
+These works do not omit computations across the entire sequence and achieve impressive speed and accuracy performance across various tasks.
 Motivation.
@@
 Moreover, to keep the accuracy of attention, SageAttention proposes smoothing $K$ by eliminating its channel-wise outliers.
-SageAttention achieves 2 $\times$ and 2.7 $\times$ speedup than FlashAttention2 and xformers and is the first quantized attention that incurs negligible end-to-end metrics loss across language, image, and video generation models.
+SageAttention achieves 2 $\times$ speedup than FlashAttention2 and is the first quantized attention that incurs negligible end-to-end metrics loss across language, image, and video generation models.
 However, SageAttention has two weaknesses.
 (W1) INT8 Matmul achieves only half the speed of INT4.
-(W2) FP16 Matmul with FP16 accumulators is only compatible with RTX4090 and RTX3090 GPUs.
-To leverage the faster INT4 tensor cores for $QK^{\top}$ and using a method that can accelerate $\widetilde{P}V$ on a broader range of GPUs, we propose SageAttention2, which quantizes $Q,K$ to INT4 and $\widetilde{P},V$ to FP8.
+(W2) FP16 Matmul with FP16 accumulators provides a speedup only on RTX 4090 and RTX 3090 GPUs.
+To leverage the faster INT4 tensor cores for $QK^{\top}$ and using a method that can accelerate $\widetilde{P}V$ on a broader range of GPUs, we propose to quantize $Q,K$ to INT4 and $\widetilde{P},V$ to FP8.
 Challenges.
 Quantizing $Q,K$ to INT4 and $\widetilde{P},V$ to FP8 presents significant challenges.
-For example, when only per-tensor quantizing $Q,K$ to INT4, the text-to-video model CogvideoX will generate a completely blurry video (See Figure 2), and Llama2 only achieves a random-guessing-level accuracy of 25% on MMLU.
+For example, when only per-tensor quantizing $Q,K$ to INT4, the text-to-video model CogvideoX will generate a completely blurry video, and Llama3 only achieves a random-guessing-level accuracy of 25% on MMLU.
 After investigating deeply, we identified three primary challenges: (C1) The numerical range of INT4 is quite limited compared to FP16 or INT8, leading to significant quantization errors when $Q$ and $K$ have some abnormal values.
@@
 This will lead to an accuracy loss of $\widetilde{P}V$.
-Figure 2: An example of quantizing Q, K to INT4 from CogvideoX.
 Our approach.
-To address (C1), we first discovered that the per-block quantization of $Q,K$ in SageAttention does not achieve sufficient accuracy for INT4 quantization.
+To address (C1), we first discover that the per-block quantization of $Q,K$ in SageAttention does not achieve sufficient accuracy for INT4 quantization.
 Simultaneously, to avoid the extra latency caused by per-token dequantization, where each GPU thread corresponds to multiple quantization scales, we propose an adequately precise quantization method that incurs no additional dequantization overhead.
-Specifically, we introduce a per-thread quantization approach based on the mapping between the GPU thread and memory layout of matrices as dictated by the PTX mma instructions.
+Specifically, we introduce a per-thread quantization approach based on the mapping between the GPU threads and memory layout of matrices as dictated by the PTX mma instructions.
 This method groups tokens corresponding to the same thread for quantization and dequantization, ensuring that each thread is associated with a single quantization scale.
 This approach achieves much better accuracy performance than per-block quantization with no additional dequantization overhead.
-Second, for the significant channel-wise outliers in matrices $Q$ and $K$, we adopt smoothing $K$ in SageAttention and propose an effective method to remove these outliers in $Q$.
+Second, for the significant channel-wise outliers in matrices $Q$ and $K$, we adopt smoothing $K$ in SageAttention and further propose an effective method to remove these outliers in $Q$.
 Specifically, we propose subtracting the average value of the channel dimension of $Q$, referred to as $\overrightarrow{Q}_{m}$.
 Then, we add $\overrightarrow{Q}_{m}K$ after the $QK^{\top}$ Matmul to ensure the correctness of the attention computation.
-To address (C2)—the accuracy loss from using a 22-bit accumulator for FP8 Matmul of $\widetilde{P}V$—, we propose to use an FP32 buffer to accumulate the values from the 22-bit accumulator after each block Matmul of $\widetilde{P}V$.
+To address (C2), the accuracy loss from the 22-bit accumulator for FP8 Matmul of $\widetilde{P}V$, we propose a two-level accumulation strategy that uses an FP32 buffer to accumulate the values from the 22-bit accumulator after each block Matmul of $\widetilde{P}V$.
 This method confines the errors to the block range.
@@
 Importantly, we offer a high-performance implementation of SageAttention2 on RTX4090 and L20 GPUs.
-This implementation achieves a peak performance of 481 TOPS on the RTX4090, outperforming FlashAttention2 and xformers by approximately 3x and 5x, respectively.
-Note that FlashAttention3 is tailored to and can only be used with the Nvidia Hopper architecture, i.e., H100 and H800.
+This implementation achieves a peak performance of 481 TOPS on the RTX4090, outperforming FlashAttention2 and xformers by approximately 3x and 4.5x, respectively.
+Note that FlashAttention3 is tailored to and can only be used with the Nvidia Hopper architecture.
+Moreover, SageAttention2 matches the speed of FlashAttention3(fp8) on the Hopper GPUs, while delivering much better accuracy.
+For example, on CogVideoX-1.5 (Yang et al., 2025), Mochi (Team, 2024) and HunyuanVideo (Kong et al., 2024), our method does not compromise end-to-end accuracy, whereas videos generated using FlashAttention3(fp8) often suffer noticable degradation, as visualized in Fig.
+8.
 We extensively evaluate the end-to-end metrics of state-of-the-art text, image, and video generation models using SageAttention2.
-Across all models and tasks, SageAttention2 can be directly adopted in a plug-and-play manner with negligible loss in model performance.
+SageAttention2 can accelerate models in a plug-and-play way with negligible loss in end-to-end metrics.
 2 Preliminary
 2.1 FlashAttention
-The computation of self-attention can be formulated as follows: $S=QK^{\top}/\sqrt{d},~{}P=\sigma(S),~{}O=PV$, where $\sigma(S)_{ij}=\exp(S_{ij})/\sum_{k}\exp(S_{ik})$ is the softmax operation.
-The matrices $Q$, $K$, and $V$ each have dimensions $N\times d$, while the matrix $S$, $P$ are $N\times N$.
-While $d$ is typically small, e.g., 64 or 128, $N$ can be millions.
-Therefore, the $N\times N$ matrices $(S,P)$ are much larger than $(Q,K,V)$, and a naive implementation suffers from the huge amount of global memory IO for $(S,P)$ reads/writes.
-FlashAttention (Dao, 2023) proposes to tile $Q$, $K$, and $V$ from the token dimension into blocks $\{Q_{i}\},\{K_{i}\},\{V_{i}\}$ with block size of $b_{q}$, $b_{k}$, $b_{v}$, respectively, where $b_{k}$ = $b_{v}$.
-Then, to avoid the memory IO for $(S,P)$, it uses online softmax (Milakov & Gimelshein, 2018) to progressively compute each block of $O$, i.e., $O_{i}$:
-First, for each block of $\{K_{i}\},\{V_{i}\}$, it computes the following equations iteratively:
-$S^{j}_{i}=Q_{i}K_{j}^{\top}/\sqrt{d},~{}~{}(m^{j}_{i},\widetilde{P}^{j}_{i})=\tilde{\sigma}(m^{j-1}_{i},S^{j}_{i}),~{}~{}l_{i}^{j}=$ $\exp(m^{j-1}_{i}-m^{j}_{i})l_{i}^{j-1}+\mathrm{rowsum}(\widetilde{P}^{j}_{i}),$
-$O^{j}_{i}=\mathrm{diag}\left(\exp(m^{j-1}_{i}-m^{j}_{i})\right)O^{j-1}_{i}$ $+\widetilde{P}^{j}_{i}V_{j}$
-Where $m_{i}^{j}$ and $l_{i}^{j}$ are $b_{q}\times 1$ vectors, which are initialized to $\infty$- ∞ and 00 respectively.
-The $\tilde{\sigma}()$ is an online softmax operator: $m^{j}_{i}=\max\{m^{j-1}_{i},\mathrm{rowmax}(S^{j}_{i})\},~{}\widetilde{P}_{j}^{i}=\exp(S^{j}_{i}-m^{j}_{i})$.
-Finally, the output $O_{i}$ can be computed by $O_{i}=\mathrm{diag}(l_{i}^{j})^{-1}O_{i}^{j}$.
+The attention computation can be formulated as: $S=QK^{\top}/\sqrt{d},~{}P=\sigma(S),~{}O=PV$, where $\sigma(S)_{ij}=\exp(S_{ij})/\sum_{k}\exp(S_{ik})$.
+The matrices $Q$, $K$, and $V$ each has dimensionality $N\times d$, and $S$, $P$ are $N\times N$.
+$d$ is typically small, e.g., 64 or 128, and $N$ can be thousands or millions.
+The time complexity of attention is $O(N^{2})$, primarily due to two matrix multiplications ($QK^{\top}$ and $PV$), both with complexities of $O(N^{2}d)$.
+FlashAttention (Dao, 2024) is a GPU-friendly attention implementation, which tiles $Q$, $K$, and $V$ from the token dimension into blocks $\{Q_{i}\}_{i=1}^{n_{q}},\{K_{i}\}_{i=q}^{n_{k}},\{V_{i}\}_{i=1}^{n_{v}}$ with block sizes of $b_{q}$, $b_{k}$, $b_{v}$ tokens, respectively, where $n_{q},n_{k},n_{v}$ are the number of tiles, and $b_{k}$ = $b_{v}$.
+FlashAttention computes the output matrix $O$ in parallel in tiles.
+Each streaming multiprocessor (SM) computes a block $O_{i}$ (corresponds to a $Q_{i}$) by iteratively loads $K_{j},V_{j}$ for each $j$, and update the output with online softmax (Milakov & Gimelshein, 2018):
+$S_{ij}=$ $Q_{i}K_{j}^{\top}/\sqrt{d},~{}~{}(m_{ij},\widetilde{P}_{ij})=\tilde{\sigma}(m_{i,j-1},S_{ij}),$ (1)
+$l_{ij}=$ $\exp(m_{i,j-1}-m_{ij})l_{i,j-1}+\mathrm{rowsum}(\widetilde{P}_{ij}),$
+$O_{ij}=$ $\mathrm{diag}\left(\exp(m_{i,j-1}-m_{ij})\right)O_{i,j-1}+\widetilde{P}_{ij}V_{j},$
+where $m_{ij}$ and $l_{ij}$ are $b_{q}$-dimenalional vectors, initialized with $\infty$- ∞ and 00 respectively.
+$\tilde{\sigma}(\cdot)$ is an online softmax operator: $m_{ij}=\max\{m_{i,j-1},\mathrm{rowmax}(S_{ij})\},~{}\widetilde{P}_{ij}=\exp(S_{ij}-m_{ij})$.
+Finally, the output is computed as $O_{i}=\mathrm{diag}(l_{i,n_{q}})^{-1}O_{i,n_{q}}$.
 2.2 Quantization
 A matrix multiplication $C=AB$ can be accelerated with quantization as:
-$(\delta_{A},\hat{A})=\psi(A),~{}~{}(\delta_{B},\hat{B})=\psi(B),~{}~{}\hat{C}=\hat{A}\hat{B},$ $~{}~{}C=\psi^{-1}_{\delta_{A}\delta_{B}}(\hat{C})$ (1)
+$(\delta_{A},\hat{A})=\psi(A),~{}~{}(\delta_{B},\hat{B})=\psi(B),~{}~{}C=\psi^{-1}_{\delta_{A}\delta_{B}}(\hat{A}\hat{B})$
 $\psi$ is a quantizer which converts a high-precision (e.g., FP32 or FP16) matrix $A$ to a low-precision format $\hat{A}$ (e.g., INT4 or FP8) with a scale $\delta_{A}$, and $\psi^{-1}$ is a dequantizer to convert back to high-precision.
 We should have $\psi^{-1}_{\delta_{A}}(\hat{A})\approx A$.
-The actual matrix multiplication $\hat{A}\hat{B}$ is carried in low precision.
+The actual matrix multiplication $\hat{A}\hat{B}$ is carried out with low precision.
 In modern GPUs, low-precision matrix multiplication is usually multiple times faster than higher-precision ones.
-Many quantizers depend on the numerical format and granularity, e.g., how many elements share a common scale factor.
-For example, an INT4 per-tensor quantizer first computes the scale as the maximum absolute value of the entire tensor, scales the elements to the maximum representable range of INT4 [-7, +7], and then casts to INT4 with rounding: $\hat{A}=\lceil A/\delta_{A}\rfloor,\delta_{A}=\max(|{A}|)/7$.
-Likewise, per-token quantizer assigns a scale factor for each token of a tensor: $\hat{A}[i,:]=\lceil A[i,:]/\delta_{A}[i]\rfloor,\delta_{A}[i]=\max(|{A[i,:]}|)/7$.
-Also, per-channel quantizer assigns a scale factor for each channel of the tensor, i.e., along the channel dimension: $A[:,i]=\lceil A[:,i]/\delta_{A}[i]\rfloor,\delta_{A}[i]=\max(|A[:,i]|)/{7}$.
+Quantizers differ in numerical format and granularity, e.g., how many elements (“quantization group”) share a common scale factor.
+For example, an INT4, per-tensor quantizer first computes the scale as the maximum absolute value of the entire tensor, scales the elements to the maximum representable range of INT4 [-7, +7], and then casts to INT4 with rounding: $\hat{A}=\lceil A/\delta_{A}\rfloor,\delta_{A}=\max(|{A}|)/7$.
 The dequantization process is an element-wise scaling.
-For example, for per-tensor dequantization, $\psi_{\delta_{A}\delta_{B}}^{-1}(\hat{A}\hat{B})=\hat{A}\hat{B}*\delta_{A}*\delta_{B}$.
+For example, for per-tensor dequantization, $\psi_{\delta_{A}\delta_{B}}^{-1}(\hat{A}\hat{B})=\hat{A}\hat{B}\times\delta_{A}\delta_{B}$.
+Table 1: Speedup compared to matrix multiplication in FP16 with an FP32 accumulator.
+GPU MM Input MM Accumulator Speedup
+RTX4090 FP16 FP16 2x
+FP8 FP32 2x
+L40, L20 H100 FP16 FP16 1x
+FP8 FP32 2x
 2.3 SageAttention
-Based on the block tiling in FlashAttention2 (Dao, 2023), SageAttention (Zhang et al., 2024) quantize $Q,K$ to INT8 in a per-block granularity.
-Moreover, to keep the quantization accuracy, SageAttention proposes to smooth $K$ first:
-$K=K$ $-\text{mean}(K)$
-$\hat{Q}_{i}=\lceil Q_{i}/\delta_{Q}\rfloor,~{}~{}\delta_{Q}=\max(|{Q_{i}|})/127,$ $~{}~{}\hat{K}_{j}=\lceil K_{j}/\delta_{K}\rfloor,~{}~{}\delta_{K}=\max(|{K_{j}|})/127$
-$S_{ij}=$ $Q_{i}K_{j}^{\top}*\delta_{Q}*\delta_{K}$
-Where $Q_{i},K_{j}$ are the tilled block in FlashAttention, and $\text{mean}(K)$ is the average value of the channel dimension of $K$.
-Additionally, SageAttention keeps $\widetilde{P}V$ as FP16 and uses FP16 Matmul with an FP16 accumulator for $\widetilde{P}V$.
-However, FP16 Matmul with an FP16 accumulator only has a speedup effect on RTX4090 and RTX3090 GPUs.
-3 SageAttention-2
+Based on the block tiling in FlashAttention (Dao et al., 2022), SageAttention (Zhang et al., 2025) quantizes $Q,K$ to INT8 in a per-block granularity, i.e., each $Q_{i},K_{i}$ has a separate scalar scale: $\delta_{Q_{i}}=\max(|{Q_{i}|})/127,\delta_{K_{j}}=\max(|{K_{j}|})/127$.
+In this way, the product $S_{ij}$ in Eq.
+(1) can be approximated as $S_{ij}\approx\hat{Q}_{i}\hat{K}_{j}^{\top}\times(\delta_{Q_{i}}\delta_{K_{j}}/\sqrt{d})$.
+To maintain the accuracy, SageAttention proposes a preprocessing technique to by subtracting the token-wise mean from $K$.
+Additionally, SageAttention keeps $\widetilde{P}_{ij}$ and $V_{j}$ in FP16, but utilizes an FP16 accumulator (rather than FP32) for computing the product $\widetilde{P}_{ij}V_{j}$.
+Reducing the accumulator precision can accelerate the matrix multiplication (MM) on RTX4090 GPU.
+However, other GPUs, such as L20, L40, or H100, do not exhibit this behavior, as shown in Table 1.
+3 SageAttention2
+In this section, we introduce SageAttention2, an efficient and accurate quantized attention.
+The workflow of SageAttention2 is shown in Fig.
+2.
+We quantize $Q,K$ to INT4 and $\tilde{P},V$ to FP8 to maximize the efficiency and propose several techniques, including $QK$-smoothing, per-thread quantization, and two-level accumulation to preserve the accuracy, which we shall discuss in subsequent subsections.
 Figure 2: Workflow of SageAttention2. 1 Smooth Q,K,V. 2 A GEMV to obtain $\Delta S$. 3 Per-thread quantize Q,K and per-channel quantize V. 4 Perform the SageAttention2 kernel. 5 Correct the output.
-3.1 Formulation
-Based on the introduction of FlashAttention and quantization presented in Section 2, we now describe the quantized attention approach we developed.
-Quantization: ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{Q},\hat{Q},~{}~{}\overrightarrow{Q}_{m})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\phi_{Q}}}(Q),~{}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{K},\hat{K})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\phi_{K}}}(K),~{}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{V},\hat{V},\overrightarrow{V}_{m})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\phi_{V}}}(V)$
-${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\Delta S=\overrightarrow{Q_{m}}K^{\top}}}$ (2)
-Attention: $S={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi^{-1}_{\delta_{Q}\delta_{K}}(\hat{Q}\hat{K}^{\top})+\Delta S}},~{}~{}(m^{\prime},\widetilde{P})=\tilde{\sigma}(m,S),~{}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{P},\hat{P})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{P}}}(\widetilde{P})$
-$O=\mathrm{diag}\left(\exp(m^{\prime}-m)\right)O+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi^{-1}_{\delta_{P}\delta_{V}}(\hat{P}\hat{V})}}$ (3)
-$\phi_{Q}$, $\phi_{K}$, $\phi_{V}$ are three transformations to obtain quantized $Q$, $K$, and $V$, which we will discuss in subsequent sections.
-For simplicity, we omit all superscripts and subscripts, but the matrices used in attention are still tiles, and the computation is still organized as FlashAttention described in Section 2.1.
-Compared to the original full-precision version, as shown in Eq.
-23, SageAttention2 adds quantizers to $Q,K,P,V$ and dequantizers to the product to accelerate both Matmuls of $QK$ and $\widetilde{P}V$.
 Figure 3: Typical examples of tensors’ data distribution in attention.
-Table 1: End-to-end metrics comparison of different quantization methods, where $Q,K$ are quantized into INT4, while $\widetilde{P},V$ stay in full precision.
-Q, K Smoothing (Q+K) Llama3.1 (Lambda) $\uparrow$ Llama3.1 (WikeText) $\downarrow$ CogVideoX (vqa-a) $\uparrow$ CogVideoX (vqa-t) $\uparrow$
-Full-Precision - 81.5% 6.013 77.605 75.360
-INT4 Quantization 72.6% 11.698 27.114 24.670
-80.8% 6.219 77.276 75.147
-Figure 5: An example of quantized value distribution of $Q$ before and after smoothing $Q$.
 3.1 Smooth Q
-The representative range for quantization of INT4 is notably restrictive, i.e., only $152^{4}-1=$ 15 values.
-This limitation significantly degrades performance when quantizing $Q,K$ into INT4 in attention.
-For example, using INT4 to quantize $Q,K$ can cause the perplexity of Llama3.1 on WikiText to increase by more than 90%, and the quality of videos generated by CogvideoX decrease about 3x (See Table 1).
-We analyze the data distribution of $Q,K$ in real models.
-For example, we find that Llama3.1 and CogvideoX show distinct channel-wised outliers in $Q,K$ (See Figure 4).
-While per-channel quantization could mitigate the quantization error caused by such outliers, this method is impracticable for $Q,K$ because quantization must be conducted along the outer axis (token dimension) of $QK^{\top}$.
-To address this challenge, we propose a method to eliminate the quantization error brought by the outliers in $Q,K$:
-$\overrightarrow{Q_{m}}=\mathrm{mean}(Q),~{}~{}\gamma(Q)=$ $Q-\overrightarrow{Q_{m}},~{}~{}\gamma(K)=K-\mathrm{mean}(K),$
-$S=QK^{\top}=$ $\gamma(Q)\gamma(K)^{\top}+\overrightarrow{Q_{m}}\gamma(K)^{\top}$
-$\phi_{Q}(Q)=\psi_{Q}\circ\gamma(Q),~{}\overrightarrow{Q_{m}}$ $,~{}~{}~{}~{}~{}\phi_{K}(K)=\psi_{K}\circ\gamma(K)$ (4)
-Where $\mathrm{mean}(K)$ and $\gamma(K)$ have been discussed in SageAttention (Zhang et al., 2024), and the transformation for $K$ does not change the attention score $\widetilde{P}$.
-$\mathrm{mean}(Q)=\frac{1}{N}\sum_{t=1}^{N}Q[t,:]$ is a vector with a shape $1\times d$.
-For the transformation of $Q$, we will perform a GEMV (general matrix-vector multiplication) between $\overrightarrow{Q_{m}}$ and $K^{\top}$, i.e., $\overrightarrow{Q_{m}}K^{\top}$ as $\Delta S$.
-This $\Delta S$ will be added to $S$ during the attention computation.
-Finally, the transformation from full-precision $Q,K$ to quantized $\hat{Q},\hat{K}$ can be expressed as shown in Eqution 4, where $\psi_{Q}$, $\psi_{K}$ is two quantizers for $Q$ and $K$.
-In other words, full-precision $Q,K$ are substracted with the means of the channel dimension before being quantized.
-Figure 5 shows an example from CogvideoX of the distribution of quantized $Q$ with and without smoothing Q.
+First, we discuss how to accurately compute $QK^{\top}$ with INT4.
+The numerical range of INT4 is notably restrictive.
+This affects quantization due to the presence of outliers (Lin et al., 2024).
+Given the INT4 range [-7, +7], any element will be quantized to zero if it is more than 14 times (0.5 vs 7) smaller than the largest element in the group.
+Since outliers are much larger than other elements, it is quite likely that most non-outlier elements are quantized to zero, resulting in significant accuracy degradation.
+Therefore, to keep the quantization accurate, we need to keep the largest element small, making the magnitude of elements as uniform as possible.
+Such technique is called smoothing.
+Here, we propose a smoothing technique inspired by SageAttention (Zhang et al., 2025).
+SageAttention observed that $Q,K$ for all tokens are actually highly similar, with only small variations between different tokens (see Fig.
+3).
+We propose to smooth $K$ as SageAttention does and further smooth $Q$ by subtracting a common mean of each block:
+$\gamma(Q_{i})=Q_{i}-\bar{q}_{i},~{}~{}~{}\gamma(K_{j})=K_{j}-\bar{k},$ (2)
+where $\bar{q}_{i}=\mathrm{mean}(Q_{i}),\bar{k}=\mathrm{mean}(K)$ are $1\times D$ vectors, the mean is conducted along the token axis, and $\bar{q}_{i},\bar{k}$ are broadcasted to tokens in a block and a tensor for subtraction, respectively.
+With the decomposion, we have $S_{ij}=Q_{i}K_{j}^{\top}=(\bar{q}_{i}+\gamma(Q_{i}))(\bar{k}+\gamma(K_{j}))^{\top}=\bar{q}_{i}\bar{k}^{\top}+\bar{q}_{i}\gamma(K_{j})^{\top}+\gamma(Q_{i})\bar{k}^{\top}+\gamma(Q_{i})\gamma(K_{j})^{\top}=\gamma(Q_{i})\gamma(K_{j})^{\top}+\Delta{S_{ij}}+b$.
+Here, $\Delta{S_{ij}}=\bar{q}_{i}\gamma(K_{j})^{\top}$ is an $1\times N$ vector, and $b=\bar{q}_{i}\bar{k}^{\top}+\gamma(Q_{i})\bar{k}^{\top}$ is an $N\times 1$ vector.
+We do not need to compute $b$ since adding a common bias to an entire row of $S$ does not affect the result after softmax.
+Therefore, we can accelerate $Q_{i}K_{j}^{\top}$ with INT4 by the following two stages:
+(1) preprocessing: smooth $Q,K$ according to Eq.
+(2), apply quantization $(\delta_{Q_{i}},\hat{Q}_{i})=\psi_{Q}(\gamma(Q_{i})),(\delta_{K_{j}},\hat{K}_{j})=\psi_{K}(\gamma(K_{j}))$, and compute $\Delta{S_{ij}}=\bar{q}_{i}\gamma(K_{j})^{\top}$.
+The smoothing, quantization, and the GEMV (general matrix-vector multiplication) for computing $\Delta S$ can be fused into a single kernel, which scans the off-chip $Q$ and $K$ only once.
+(2) attention: execute the low-precision GEMM, dequantize, and add back the vector $\Delta S$: $S_{ij}=\psi^{-1}_{\delta_{Q_{i}}\delta_{K_{j}}}(\hat{Q}_{i}\hat{K}_{j}^{\top})+\Delta S_{ij}$.
+These operations are all done on chip, and the dequantization and vector addition only adds a marginal overhead compared to the expensive mma operation for MM.
+Importantly, $\gamma(Q_{i}),\gamma(K_{j})$ are quantized rather than $Q_{i},K_{j}$.
+Since the smoothed matrices are much smaller in magnitude and contain less outliers, the quantization accuracy can be significantly improved.
+Remark.
+Classical techniques to improve the activation-weight MM, such as per-channel quantization, or SmoothQuant (Xiao et al., 2023) are not applicable here for the query-key MM in attention.
+Per-channel quantization is impracticable for $Q,K$ because quantization must be conducted along the outer axis (token dimension) of $QK^{\top}$ (Xiao et al., 2023).
+On the other hand, both $Q$ and $K$ have significant outliers, so trading the quantization accuracy between them with SmoothQuant cannot work effectively, as shown in Sec.
+4.1.
+Here, we utilize the unique token-similarity pattern in attention to derive an accurate dedicated quantization method for $Q$ and $K$.
+The previous work SageAttention only smooths $K$, so it is less accurate than our method.
+Empirical results.
+Fig.
+19 in Appendix A.7 shows an example from CogvideoX of the distribution of $\hat{Q}$ with and without smoothing $Q$.
 We can find that with smoothing $Q$, the range of INT4 is utilized more uniformly and fully.
-Table 1 presents end-to-end metrics for different quantization methods with and without smoothing Q+K on Llama3.1 (Dubey et al., 2024) and CogvideoX (Yang et al., 2024).
+Table 5 presents end-to-end metrics for different quantization methods with and without smoothing Q+K on Llama3.1 (Dubey et al., 2024) and CogvideoX (2b) (Yang et al., 2025).
 The results demonstrate that smoothing Q+K offers significant accuracy benefits.
-Also, Table 9 and Table 9 show that the order of effectiveness is smoothing Q+K $>$ smoothing Q $>$ smoothing K $>$ other baselines.
+Also, Table 4 and 17 show that the order of effectiveness is smoothing Q+K $>$ smoothing Q $>$ smoothing K $>$ other baselines.
 Figure 4: An example of per-thread quantization.
@@
 The right figure shows the correspondence between quantization tokens and GPU threads in a MMA.m16n8k64 instruction, showing that each GPU thread only corresponds to one quantization scale in $\delta_{Q}$ and $\delta_{K}$ in dequantization.
+3.2 INT4 Per-thread Quantization
+Orthogonal to smoothing, we can mitigate the problem of outliers by refining the quantization granularity so the number of affected elements by outliers becomes smaller.
+Although per-token offers a detailed level of granularity, it results in significant overhead during dequantization.
+Specifically, each GPU thread in per-token quantization must handle multiple quantization scales, leading to a high latency of the dot product of the quantization scale vectors $\delta_{Q}$ and $\delta_{K}$.
+SageAttention uses per-block quantization, where each block $Q_{i}$ ($b_{q}$ tokens) and $K_{i}$ ($b_{k}$ tokens) have a single quantization scale.
+Such a quantization strategy could achieve an accuracy performance close to per-token quantization and avoid the high dequantization overhead.
+However, quantizing $Q$ and $K$ to INT4 demands a finer quantization granularity.
+To address this, we propose per-thread quantization, a more precise and granular approach than the per-block quantizer, also without the additional overhead of vectors dot product between $\delta_{Q}$ and $\delta_{K}$.
+Specifically, each block of $Q$, i.e., $Q_{i}$, in SageAttention will be split into $c_{w}$ segments and processed by $c_{w}$ GPU warps in a GPU streaming processor (SM).
+We call each segment of $Q_{i}$ as $Q_{w}$, and $k_{w}=K_{j}$ since $K_{j}$ is shared among warps.
+Then, each warp containing 32 threads uses the mma.m16n8k64 PTX instruction (NVIDIA, ) for the $Q_{w}K_{j}^{\top}$.
+According to the layout requirement of this instruction, we find that $Q_{w}[8k+i]$ could share one quantization scale, and $K_{j}[8k+2i]$ and $K_{j}[8k+2i+1]$ could share one quantization scale.
+Such a quantization method is more fine-grained and with no additional overhead.
+This is because it assigns different GPU threads to distinct quantization groups based on the MMA instruction layout, with each thread performing dequantization only using a single quantization scale value.
+We show an example of per-thread quantization in Fig.
+4.
+The detailed formulation is shown in Equation A.5 and Fig.
+17 (Please see Appendix A.5 for more detail).
+Empirical results.
+As shown in Table 6 and Table 15, we compare the average and the worst accuracy of INT4 quantization at per-token, per-thread, per-block, and per-tensor granularity using real $Q,K,V$ across all layers of CogvideoX.
+Results indicate the accuracy of per-thread quantization is very close to per-token and outperforms a lot more than others.
+Algorithm 1 Implementation of SageAttention2.
+  Input: Matrices $Q(\text{FP16}),K(\text{FP16}),V(\text{FP16})\in\mathbb{R}^{N\times d}$, block size $b_{q},b_{kv}$, warp count $c_{w}$.
+  Preprocessing: $K=K-\mathrm{mean}(K)$, ${}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{V},\hat{V})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{V}}}(V)$.
+//per-channel.
+  Divide $Q$ to $T_{m}={N}/{b_{q}}$ blocks $\{Q_{i}\}$; divide $K$, and $V$ to $T_{n}={N}/{b_{kv}}$ blocks $\{K_{i}\}$, $\{V_{i}\}$;
+  for $i=1$ to $T_{m}$ do
+    ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\bar{q}_{i}=\mathrm{mean}(Q_{i}),~{}~{}(\delta_{Q},\hat{Q}_{i})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{Q}(Q_{i}-\bar{q}_{i})}}$ //per-thread ;
+    for j in [1, $T_{n}$do
+       ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{K},\hat{K}_{j})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{K}}}(K_{j})$ //per-thread,     $w=\text{range}(c_{w}),st=w*c_{w}$ ;
+       $S_{ij}[st:st+c_{w}]=\psi^{-1}_{\delta_{Q}\delta_{K}}(\mathrm{Matmul}(\hat{Q}_{i}[st:st+c_{w}],\hat{K}_{j}^{\top}))+\text{GEMV}(\bar{q}_{i},K_{j}^{\top})$ ; // Paralleled by $c_{w}$ warps.
+The $\psi^{-1}_{\delta_{Q}\delta_{K}}$ is illustrated in Fig.
+4.
+       $m_{ij}=\mathrm{max}(m_{i,j-1},\mathrm{rowmax}(S_{ij}))$, ${}\widetilde{P}_{ij}=\mathrm{exp}(S_{ij}-m_{ij})$, ${}l_{ij}=e^{m_{i,j-1}-m_{ij}}+\mathrm{rowsum}(\widetilde{P}_{ij})$ ;
+       $O_{ij}$(FP22) = $\mathrm{Matmul}((\widetilde{P}_{ij}*448)\text{.to(FP8.e4m3)},V_{j})$ ;
+       $O_{ij}$(FP32) = $\mathrm{diag}(e^{m_{i,j-1}-m_{ij}})^{-1}O_{i,j-1}(FP32)+O_{ij}$(FP22) ;
+    end for
+[…]
+  end for
+  return $O=\{O_{i}\}$
+3.3 FP8 quantization for $\tilde{P}V$
+We now turn to the MM $\tilde{P}V$, where $\tilde{P}_{ij}=\exp(S_{ij}-m_{ij})$ is the unnormalized quantity according to Eq.
+(1).
+The distribution of $\tilde{P}$ is unique and differs from other activations.
+First, we note that $S_{ij}-m_{ij}\leq 0$, so $P_{ij}\in[0,1]$ ($\leq$ and $\in$ apply element-wisely).
+We find that $\tilde{P}$ often consists of many small elements, but their sum is non-negligible (e.g., 5000 elements around $10^{-4}$).
+In this case, we must represent small elements accurately.
+INT quantization is unsuitable for this proposal since it puts the quantization points evenly within the numerical range.
+SageAttention (Zhang et al., 2025) choose to retain $\tilde{P}$ and $V$ in FP16, and accelerate the MM by decreasing the accumulator precision.
+However, this strategy is only effective on very few GPUs.
+We propose to quantize $\widetilde{P},V$ to FP8 with 4 exponent bits and 3 mantissa bits (E4M3).
+The numerical range of E4M3 is $[-448,+448]$.
+We quantize $P$ with a static scale: $\delta_{P}=\frac{1}{448}$ since the original $P$ elements is already in $[0,1]$.
+We quantize $V$ per-channel to address the channel-wise outliers shown in Fig.
+3.
+Empirical results in Table 7 and Table 16 show the average and the worst accuracy of different data types used for $\widetilde{P},V$ across all layers of CogvideoX.
+The accumulator is always 32-bit.
+We can see that the accuracy of E4M3 is very close to that of FP16 and superior to E5M2 and INT8.
+Most modern GPUs have tensor cores that support FP8 Matmul operations, which are twice as fast as those using FP16.
+3.4 FP32 MMA Buffer for FP22 Accumulator
+While FP8 quantization for $\tilde{P}V$ above is theoretically accurate in simulation, we observe that the actual CUDA implementation suffers a consistent accuracy degradation.
+After narrowing down the problem, we find that the accumulator for the mma(f32f8f8f32) instruction on the Ada and Hopper architecture is actually FP22, specifically with 1 sign bit, 8 exponent bits, and 13 mantissa bits.
+Specifically, for mma(f32f8f8f32) instruction $C=AB+D$, where $A,B$ are FP8 matrices and $C,D$ are FP32 matrices, we initialize the $A,B$ to zero and vary $D$ to test the data type of the accumulator.
+When $D$ is initialized with 1 sign bit, 8 exponent bits, and 13 mantissa bits, the value of $C$ matches the result of the high-precision mma(f16f16f32) instruction.
+However, when $D$ is initialized with more than 13 mantissa bits, the error of $C$ corresponds to the difference between the results of mma(f32f16f16f32) and mma(f32f8f8f32).
+Consequently, the matrix multiplication of $\widetilde{P}V$, quantized to FP8, incurs a certain degree of accuracy loss compared to using an FP32 accumulator.
+To mitigate this accuracy loss, we adopt a two-level accumulation strategy, which uses an FP32 buffer to accumulate the values of $\widetilde{P}_{ij}V_{j}$ in FP22.
+Specifically, we rewrite Eq.
+(1) as $R_{ij}=\widetilde{P}_{ij}V_{j},O_{ij}=\mathrm{diag}\left(\exp(m_{i,j-1}-m_{ij})\right)O_{i,j-1}+R_{ij}$.
+Here, two sets of accumulators $R_{ij}$ and $O_{ij}$ are maintained in the register.
+$R_{ij}$ is computed with the mma(f32f8f8f32) instruction, providing 22 effective bits, which is sufficient since we only accumulate over a small number of $b_{k}$ tokens (e.g., $b_{k}=64$).
+Then, $R_{ij}$ is accumulated to $O_{ij}$ in the high FP32 precision.
+Remark.
+The two-level accumulation strategy is also implemented in Cutlass (NVIDIA, 2023) and DeepSeekv3 (DeepSeek-AI et al., 2024) for computing activation-weight products in linear layers.
+To the best of our knowledge, we are the first to discover and investigate the effect of the FP22 accumulator and implement the two-level accumulation for attention.
+Optional smooth V technique.
+We also figure out another way to mitigate the accuracy loss due to the FP22 accumulator when $V$ possesses channel-wise biases: $\overrightarrow{V_{m}}=\mathrm{mean}(V,\text{axis}=0),V=V-\overrightarrow{V_{m}}$.
+Additionally, to maintain the correctness of the attention computation, it is only necessary to add $\overrightarrow{V_{m}}$ to the final calculation of $O$: $O=O+\overrightarrow{V_{m}}$.
+This is because the sum of each row of the $\widetilde{P}$ matrix equals 1, so $\widetilde{P}\overrightarrow{V_{m}}=\overrightarrow{V_{m}}$.
+Remark.
+For details on smoothing V, see Appendix A.3.
+This technique is optional and not employed in our main experiments, as it provides significant benefits only when $V$ exhibits channel-wise bias, which are absent in some models, such as Llama3.1 (see Fig.
+3).
+Figure 5: Speed comparison between SageAttention2 and baselines (RTX4090, headdim=128).
+Table 2: End-to-end metrics across text, image, and video generation models. ✗ indicates an inability to generate results for evaluation.
+Model Attention WikiText (Ppl.) $\downarrow$ Lambda (Acc.) $\uparrow$ MMLU (Acc.) $\uparrow$ Longbench $\uparrow$
+Llama3.1 Full-Precision 6.013 0.815 0.635 49.40
+HadmdAttn 7.872 0.762 0.500 44.07
+SmoothAttn 7.180 0.783 0.541 44.69
+SageAttention 6.017 0.812 0.634 49.55
+SageAttn2-4b 6.256 0.798 0.607 48.79
+SageAttn2-8b 6.019 0.811 0.634 49.59
+GLM4 Full-Precision 7.241 0.432 0.743 49.78
+HadmdAttn 7.989 0.435 0.669 45.97
+SmoothAttn 8.943 0.449 0.592 42.20
+SageAttention 7.243 0.433 0.744 49.79
+SageAttn2-4b 7.352 0.433 0.725 49.23
+SageAttn2-8b 7.242 0.432 0.745 49.60
+Model Attention CLIPSIM $\uparrow$ CLIP-T $\uparrow$ VQA-a $\uparrow$ VQA-t $\uparrow$ FScore $\uparrow$
+CogvideoX (1.5-5B) Full-Precision 0.1778 0.9979 70.231 70.928 2.507
+HadmdAttn 0.1576 0.9933 8.990 2.299
+SmoothAttn 0.1559 0.9950 8.812 2.277
+SageAttention
+FlashAttn3-fp8 0.1562 0.9918 6.531 2.181
+SageAttn2-4b 0.1721 0.9978 57.729 52.989 2.884
+SageAttn2-8b 0.1775 0.9980 69.492 74.415 2.487
+Hunyuan Video Full-Precision 0.1783 0.9995 82.516 75.934 0.604
+HadmdAttn 0.1727 0.9989 7.514 0.762 0.175
+SmoothAttn 0.1739 0.9988 6.987 0.609 0.148
+SageAttention 0.1786 0.9995 82.496 79.843 0.597
+FlashAttn3-fp8 0.1742 0.9941 4.433 1.460
+SageAttn2-4b 0.1751 0.9995 81.478 65.371 0.610
+SageAttn2-8b 0.1782 0.9996 81.786 75.354 0.586
+Mochi Full-Precision 0.1798 0.9986 45.549 65.416 1.266
+HadmdAttn 0.1733 0.9980 9.053 25.133 0.704
+SmoothAttn 0.1687 0.9978 3.383 3.480 0.241
+SageAttention 0.1800 0.9987 48.707 63.763 1.269
+FlashAttn3-fp8 0.1762 0.9982 14.964 13.711 0.457
+SageAttn2-4b 0.1783 0.9986 35.955 43.735 1.137
+SageAttn2-8b 0.1797 0.9986 46.760 64.901 1.255
+          Model           Attention          FID $\downarrow$          sFID $\downarrow$          CLIP $\uparrow$          IR $\uparrow$
+          Flux           Full-Precision          10.960          16.648          26.180          1.009
+          HadmdAttn          11.353          18.495          26.123          0.965
+          SmoothAttn          11.149          19.017          26.109          0.959
+          SageAttention          10.944          16.641          26.171          1.008
+          SageAttn2-4b          10.577          17.497          26.141          0.998
+          SageAttn2-8b          10.927          16.723          26.175          1.009
+                   Stable-Dif          fusion3.5           Full-Precision          14.105          15.646          25.505          0.902
+          HadmdAttn          14.259          15.909          25.513          0.886
+          SmoothAttn          14.161          15.649          25.510          0.887
+          SageAttention          14.140          15.678          25.503          0.902
+          SageAttn2-4b          14.097          15.397          25.487          0.895
+          SageAttn2-8b          14.106          15.647          25.499          0.901
+4 Experiment
+Main result.
+SageAttention2 is faster than FlashAttention2 and xformers by approximately 3x and 4.5x, respectively.
+Moreover, SageAttention2 matches the speed of FlashAttention3(fp8) on the Hopper GPUs and is much more accurate than FlashAttention3(fp8).
+Our approach maintains end-to-end metrics across language, image, and video generation models.
+4.1 Setup
+Models.
+We validate the effectiveness of SageAttention2 across a diverse set of representative models from language, image, and video generation.
+Specifically, we conduct experiments on ten models: Llama2 (7B) (Touvron et al., 2023), Llama3.1 (8B) (Dubey et al., 2024), and GLM4 (9B) (GLM et al., 2024) for text2text, CogvideoX (2B), CogvideoX (1.5-5B) (Yang et al., 2025), HunyuanVideo (Kong et al., 2024), and Mochi (Team, 2024) for text2video, Flux (schnell) (Black Forest Labs, 2023) and Stable-Diffusion3.5 (turbo) (Stability AI, 2023) for text2image, and TIMM (Wightman, 2019) for image classification.
+Datasets and metrics. For Details about the datasets and metrics we used, please refer to Appendix. A.6.
+Table 3: Two kernel implementations of SageAttention2.
+Kernel $\psi_{Q}(Q)$, $\psi_{K}(K)$ $\psi_{P}(\widetilde{P})$, $\psi_{V}(V)$
+SageAttn2-4b INT4 per-thread FP8 per-block and per-channel
+SageAttn2-8b INT8 per-thread FP8 per-block and per-channel
+Implemetation. We implement two attention kernels as shown in Table 3 using CUDA.
+Baselines.
+(1) SmoothAttn.
+Following Qserve (Lin et al., 2024), we apply smooth quant for $Q,K$ with smoothing factor $\alpha=0.5$.
+(2) HadmdAttn.
+Following Quarot (Ashkboos et al., 2024), we apply random Hadamard transformation for $Q,K$ before INT4 quantization.
+(3) SageAttention (Zhang et al., 2025), which uses smoothing $K$, INT8 per-block quantization for $Q,K$, and FP16 for $\widetilde{P},V$.
+(4) FlashAttn3(fp8), the FP8 version of FlashAttention3, which only runs on Hopper GPUs.
+Table 4: Average accuracy across all layers of CogvideoX using different smoothing methods.
+Method CosSim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
+None 80.04% 0.3906 0.2223
+HadmdAttn 79.77% 0.3782 0.2180
+SmoothAttn 90.21% 0.3383 0.1952
+Smooth K 98.07% 0.1493 0.0743
+Smooth Q 98.30% 0.1250 0.0712
+Smooth Q+K 99.46% 0.0648 0.0334
+Table 5: End-to-end metrics comparison, where $Q,K$ are quantized into INT4, while $\widetilde{P},V$ stay in full precision.
+Q, K Smooth (Q+K) Llama3.1 (Lambda) $\uparrow$ Llama3.1 (WikiText) $\downarrow$ CogVideoX (vqa-t) $\uparrow$
+Full-Precision - 81.5% 6.013 75.360
+INT4 Quantization 72.6% 11.698 24.670
+80.8% 6.219 75.147
 Table 6: Average accuracy across all layers of CogvideoX using different quantization granularities.
@@
 Per-tensor 97.15% 0.1800 0.0865
+Table 7: Average accuracy using different data types of $(\widetilde{P},V)$ across all layers of CogvideoX, where $(Q,K)$ are smoothed.
+$Q,K$ $\widetilde{P},V$ Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
+INT4 INT8 77.05% 0.5618 0.5044
+E5M2 99.20% 0.0905 0.0903
+E4M3 99.44% 0.0683 0.0347
+FP16 99.45% 0.0649 0.0335
+Table 8: End-to-end generation latency using SageAttention2 (The latency of Llama3.1 is the time to first token generation using different sequence lengths).
+Model GPU Original SageAttn 2-8b SageAttn 2-4b
+CogvideoX (2B) RTX4090 86 s 54 s 52 s
+CogvideoX (1.5-5B) RTX4090 1040 s 577 s 555 s
+HunyuanVideo L20 2221 s 1486 s 1435 s
+Mochi L20 2336 s 1316 s 1190 s
+Llama3.1 (48K token) RTX4090 9.2 s 5.7 s 5.6 s
+Llama3.1 (100K token) L20 39.9 s 25.4 s 23.2 s
+4.2 Speed and Accuracy of Kernels
+Kernel Speed.
+We compare the Speed of SageAttention2 against baselines using configurations with headdim=64 and headdim=128, both with and without Causal Mask (Vaswani, 2017).
+Specifically, Fig.
+5 shows the speed across varying sequence lengths on RTX4090, indicating that SageAttn2-4b and SageAttn2-8b are approximately 3x and 2.7x faster than FlashAttention2, and about 4.5x and 4x faster than xformers, respectively.
+Fig.
+91011121314, and 15 in Appendix A.2 show more kernel speed comparison on RTX4090, L20, H20, H100 GPUs.
+Accuracy.
+Table 4 and 17 show the average accuracy of different methods with INT4 $Q,K$ and FP8 $P,V$ across all layers of CogvideoX.
+The results indicate the accuracy of SageAttn2-4b is superior to other baselines.
+4.3 End-to-end Performance
+Metrics loss.
+We assessed the end-to-end metrics of various models using SageAttention2 compared to baselines.
+Detailed evaluation results are presented in Table 2.
+The results indicate that SageAttn2-4b outperforms all baselines and maintains most of the end-to-end accuracy across all models.
+Additionally, SageAttn2-8b incurs almost no metrics loss across various models.
+More experiment results on other models are shown in Appendix A.7.
+Visible image and video examples.
+Fig.6 and Fig.
+8 in Appendix A.1 show some visible comparision examples from HunyuanVideo, Mochi and CogvideoX (1.5-5B).
+We can observe that SageAttn2-8b does not introduce any visible differences compared to full-precision attention, whereas SageAttn2-4b has minor differences but is much better than the baselines.
+End-to-end speedup.
+We compared the original generation latency and the latency using SageAttention2 for models with long sequence lengths in Table 8, observing significant speedup effects.
+For instance, SageAttention2 achieved a 1.8x speedup in CogvideoX (1.5-5B) without any metrics loss (SageAttn2-8b).
+SageAttn2-4b further accelerated these models but with a little metrics loss.
+4.4 Ablation Study
+Overhead of techniques we proposed.
+As shown in Table 18, the overhead on kernel speed of per-thread quantization, smoothing Q, and two-level accumulation are 0.35%, 3.7%, and 0% compared to the attention kernel.
+Benefit of smoothing V. The experiment showing the benefit of smoothing $V$ is shown in Appendix. A.4.
+5 Conclusion
+We introduce SageAttention2, an efficient and accurate quantized attention.
+First, we propose to quantize matrixes $(Q,K)$ in a thread-level granularity and $(\widetilde{P},V)$ to FP8.
+Second, we propose a method to smooth $Q$, enhancing the accuracy of $QK^{\top}$.
+Third, we propose a two-level accumulation strategy to enhance the accuracy of FP8 $\widetilde{P}V$.
+SageAttention2 is faster than FlashAttention2 and xformers by approximately 3x and 4.5x, respectively.
+Moreover, SageAttention2 matches the speed of FlashAttention3(fp8) on the Hopper GPUs, but offers significantly higher accuracy.
+Extensive testing confirms that our approach maintains end-to-end metrics across language, image, and video generation models.
+Appendix A Appendix
+A.1 Visible Comparison Exmaples
+Figure 6: Comparison examples from HunyuanVideo, prompts are sampled from open-sora prompt sets.
+Figure 7: Comparison examples from CogvideoX (2B), prompts are sampled from open-sora prompt sets.
+Figure 8: A comparison example using SageAttn2-8b and FlashAttention3 on CogvideoX-1.5, Mochi and HunyuanVideo.
+A.2 Additional Kernel Speed Comparison
+Fig.
+9,  10,  11,  12,  13,  14, and  15 compare the speed of SageAttention2 against baselines using configurations with headdim=64 and headdim=128, both with and without Causal Mask (Vaswani, 2017), on RTX4090, L20, H100, and H20 GPUs.
+Table 9 summarizes the performance gain of different attention methods against baselines on various modern GPUs.
+Figure 9: Speed comparison between SageAttention2 and baselines (RTX4090, headdim=64).
+Figure 10: Speed comparison between SageAttention2 and baselines (L20, headdim=64).
+Figure 11: Speed comparison between SageAttention2 and baselines (L20, headdim=128).
+Figure 12: Speed comparison between SageAttention2 and baselines (H100, headdim=64).
+Figure 13: Speed comparison between SageAttention2 and baselines (H100, headdim=128).
+Figure 14: Speed comparison between SageAttention2 and baselines (H20, headdim=64).
+Figure 15: Speed comparison between SageAttention2 and baselines (H20, headdim=128).
+Table 9: Speedup of different attention methods on various GPUs.
+Method 3090 4090 A100 L40 L20 H100 H20
+FlashAttention2 1.00 1.00 1.00 1.00 1.00 1.00 1.00
+FlashAttention3 1.37 1.57
+FlashAttention3 (fp8) 2.63 3.06
+SageAttention1 1.97 1.96 1.37 1.45 1.24 1.53 1.52
+SageAttention2 2.93 2.60 2.46 2.61 3.12
+Figure 16: An example of dot product precison a row of $\widetilde{P}$ and a column of $V$ presented by FP22 data type.
+Table 10: An accuracy example on real tensors of CogvideoX model with or without smoothing $V$.
+Smooth V Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
+98.25% 0.1980 0.2387
+99.75% 0.0406 0.0773
+A.3 Smoothing V
+As shown in Fig.
+16, this strategy could enhance the accuracy of FP22 for values in $\widetilde{P}V$ for the following reasons: Each row of $\widetilde{P}$ spans a value range from 0 to 1, and each column of $V$ in some models consistently features channel-wise biases that are exclusively positive or negative, for instance, ranging between 8 and 9 in CogvideoX.
+Consequently, the values of $\widetilde{P}V$ could be quite large.
+However, the floating-point number representation range is not uniform—it is denser near zero.
+Therefore, by subtracting the mean $\overrightarrow{V_{m}}$ along the channel dimension from $V$, the values of $\widetilde{P}V$ will be closer to zero, resulting in higher representational precision (see Fig.
+16 for a visual demonstration).
+Additionally, to maintain the correctness of the attention computation, it is only necessary to add $\overrightarrow{V_{m}}$ to the final calculation of $O$: $O=O+\overrightarrow{V_{m}}$.
+This is because the sum of each row of the $\widetilde{P}$ matrix equals 1, so $\widetilde{P}\overrightarrow{V_{m}}=\overrightarrow{V_{m}}$.
+In other words, this method decomposes $V$ into two parts: $\overrightarrow{V_{m}}$ and $V$.
+For $V$, it centers the values of each column around zero, which leads to the dot product result between a row from the quantized $\widetilde{P}$ matrix and a column from the quantized $V$ matrix being closer to zero.
+This makes the representation of FP22 more accurate.
+Meanwhile, $\overrightarrow{V_{m}}$ is retained in FP16 and added to $O$ at the end without causing a loss of precision for the $\overrightarrow{V_{m}}$ part.
+A.4 Experiment of Smoothing V
+Table 10 shows the attention accuracy on real tensors sampled from CogvideoX with and without smoothing $V$.
+It demonstrates that smoothing V could improve the accuracy of SageAttention2 when quantizing $Q,K$ to INT4 and $\widetilde{P},V$ to FP8.
+A.5 Per-Thread Quantization Formulation
+Figure 17: Memory layout of INT4/INT8 tensor core for accumulator matrix $C$ and $D$ in $D=A*B+C$ among 32 threads (T0 $\sim$ T31) in a warp.
+$C$ and $D$ is of shape 16x8.
+Each thread only holds 4 out of the 128 elements.
+To further clarify the per-thread quantization, we first introduce the INT4 MMA instruction of Tensor Core, and then give the formulation of per-thread quantization.
+Tensor cores, first introduced in NVIDIA’s Volta architecture, are specialized units designed for efficient matrix-multiply-and-accumulate (MMA) operations.
+Their usage significantly enhances computational efficiency and performance in AI and high-performance computing (HPC) workloads.
+Tensor cores compute small tiles of MMA operations, specifically $D=A*B+C$ on a warp (32 contiguous threads) basis.
+Each thread in the warp holds a fragment of input matrices and will get a fragment of output matrix as a computation result.
+The INT4 mma.m16n8k64 tensor core operation computes the product of a $16\times 64$ INT4 matrix $A$ and a $64\times 8$ INT4 matrix $B$, both stored in registers.
+It accumulates the result into a $16\times 8$ INT32 matrix $C$, also stored in registers, and returns the final product matrix $D$, which has the same shape ($16\times 8$), data type (INT32), and storage location (registers).
+Each thread holds only $\frac{1}{32}$ of the input and output data.
+Fig.
+17 extracted from the PTX document (NVIDIA, ) shows the memory layout of matrix $C$ and $D$ among 32 threads in a warp.
+Each thread only holds 4 out of the 128 result elements.
+$i_{\delta q}=$ $\lfloor(n*8*c_{w}/b_{q})\rfloor$
+$q_{i}[i_{\delta q}]=$$\{8\times(n\%8)+\lfloor(n*\frac{c_{w}}{b_{q}})\rfloor*\frac{b_{q}}{c_{w}}\},~{}n\in[0,N]$
+$\delta_{Q}[i_{\delta q}]=$ $\frac{\max(\left|\,Q[q_{i}[i_{\delta q}]]\,\right|)}{7}$
+$\hat{Q}[q_{i}[i_{\delta q}]]=$ $\left\lceil\frac{Q[q_{i}[i_{\delta q}]]}{\delta_{Q}[i_{\delta q}]}\right\rfloor$
+$i_{\delta k}=$ $\lfloor(n*4/b_{k})\rfloor$ (3)
+$k_{n}[i_{\delta k}]=$$\{8\times(n\%8)+\lfloor n/b_{k}\rfloor*b_{k}\}\cup$
+$\{8\times(n\%8)+1+\lfloor n/b_{k}\rfloor*b_{k}\},~{}~{}n\in[0,N]$
+$\delta_{K}[i_{\delta k}]=$ $\frac{\max(\left|\,K[k_{n}[i_{\delta k}]]\,\right|)}{7}$
+$\hat{K}[k_{n}[i_{\delta k}]]=$ $\left\lceil\frac{K[k_{n}[i_{\delta k}]]}{\delta_{K}[i_{\delta k}]}\right\rfloor$
+In Eq. A.5, $c_{w}$ is the count of GPU Warps. $b_{q}$ and $b_{k}$ are the block size of $Q,K$. $n$ is the token index of $Q,K$.
+By ensuring results held by each thread share a common dequantization scale (belong to the same quantization group), we can avoid the overhead associated with per-token quantization.
+Leveraging this observation, we design per-thread quantization, as shown in Fig.
+4.
+For typical block size of $b_{q}=128$, $b_{k}=64$ and warp number $c_{w}=4$ (as used in FlashAttention2), each warp processes a tile of 32 query tokens and 64 key tokens.
+Query tokens $i,8+i,16+i,24+i$ ($i=0,1,\cdots,7$) can be made into one quantization group and key tokens $j,1+j,8+j,9+j,\cdots,56+j,57+j$ ($j=0,1,2,3$) can be made into one quantization group.
+This design aligns with the memory layout of output matrix $D$ of tensor core shown in Fig.
+17, ensuring that each thread only needs one $Q$ scale and one $K$ scale for dequantization.
+As a result, this approach creates 32 quantization groups for $Q$ (8 for each of the 4 warps) and 4 quantization groups for $K$ in a 128x64 block, providing 32×32$\times$32 × and 4×4$\times$4 × finer granularity compared to per-block quantization for query tokens and key tokens, respectively.
+Table 6 and Table 15 show the accuracy gains by using per-thread quantization.
+Per-thread quantization achieves accuracy that closely matches per-token quantization, without introducing any kernel speed degradation (see Fig.
+18).
+A.6 Datasets and Metrics in Experiments
+Datasets.
+Text-to-text models are evaluated on four zero-shot tasks: WikiText (Merity et al., 2022) to assess the model’s prediction confidence, LAMBADA (Paperno et al., 2016) evaluate contextual understanding, MMLU (Hendrycks et al., 2020) for measuring knowledge across various subjects, and Longbench (Bai et al., 2024) for comprehensive assessment of long context understanding capabilities.
+Text-to-video models are evaluated using the open-sora (Zheng et al., 2024) prompt sets.
+Text-to-image models are assessed on MJHQ-30K (Li et al., 2024).
+TIMM is evaluated on on three image datasets: ImageNet (Deng et al., 2009), ImageNet-Sketch (Sketch) (Wang et al., 2019), and ImageNet-Rendition (ImageNet-r) (Hendrycks et al., 2021).
+End-to-end metrics.
+For text-to-text models, we use perplexity (ppl.) (Jelinek et al., 1977) for WikiText, Accuracy (Acc.) for LAMBADA and MMLU, and Longbench score (Bai et al., 2024).
+For text-to-video models, following Zhao et al.
+(2025), we evaluate the quality of generated videos on five metrics: CLIPSIM and CLIP-Temp (CLIP-T) (Liu et al., 2024) to measure the text-video alignment; (VQA-a) and (VQA-t) to assess the video aesthetic and technical quality, respectively; and Flow-score (FScore) for temporal consistency (Wu et al., 2023).
+For text-to-image models, generated images are compared with the images in MJHQ-30K dataset in three aspects: FID (Heusel et al., 2017) and sFID (Salimans et al., 2016) for fidelity evaluation, Clipscore (CLIP) (Hessel et al., 2021) for text-image alignment, and ImageReward (IR) (Xu et al., 2023) for human preference.
+For TIMM, we use classification accuracy.
+Accuracy metrics.
+We use three metrics to assess the accuracy of quantized attention output $O^{\prime}$ compared to attention output in full-precision $O$: First, we flatten $O^{\prime}$ and $O$ into vectors in the shape of $1\times n$.
+Then, Cosine similarity: $CosSim=\sum OO^{\prime}/\smash{\sqrt{\sum O^{2}}}\smash{\sqrt{\sum O^{\prime 2}}}$, Relative L1 distance: $L1=\sum|O-O^{\prime}|/\sum|O|$, Root mean square error: $RMSE=\sqrt{(1/n)\sum(O-O^{\prime})^{2}}$.
+Table 11: End-to-end metrics on Llama2 (7B).
+Model Attention WikiText (Ppl.) $\downarrow$ Lambda (Acc.) $\uparrow$ MMLU (Acc.) $\uparrow$
+Llama2 Full-Precision 5.823 0.886 0.439
+HadmdAttn 6.771 0.860 0.360
+SmoothAttn 6.717 0.867 0.392
+SageAttention 5.824 0.887 0.439
+SageAttn2-4b 5.912 0.881 0.428
+SageAttn2-8b 5.828 0.886 0.438
+Table 12: End-to-end metrics on CogvideoX (2B).
+Model Attention CLIPSIM $\uparrow$ CLIP-T $\uparrow$ VQA-a $\uparrow$ VQA-t $\uparrow$ FScore $\uparrow$
+CogvideoX (2B) Full-Precision 0.1836 0.9975 77.605 75.360 3.006
+HadmdAttn 0.1742 0.9877 29.780 23.985 0.499
+SmoothAttn 0.1741 0.9870 41.703 47.043 0.624
+SageAttention 0.1833 0.9976 76.997 71.360 2.988
+SageAttn2-4b 0.1821 0.9973 77.368 74.906 2.603
+SageAttn2-8b 0.1829 0.9977 76.532 74.281 2.941
+Table 13: End-to-end metrics on an image classification model.
+Model Attention ImageNet (Acc.) $\uparrow$ Sketch (Acc.) $\uparrow$ ImageNet-r (Acc.) $\uparrow$
+TIMM Full-Precision 84.79% 45.32% 59.55%
+HadmdAttn 84.50% 44.89% 58.80%
+SmoothAttn 84.40% 44.68% 58.73%
+SageAttention 84.74% 45.38% 59.95%
+SageAttn2-4b 86.67% 45.24% 59.29%
+SageAttn2-8b 84.79% 45.39% 59.57%
+Table 14: Comparison with FlashAttention3(fp8) on Llama-3-262k (8B) on InfiniBench (Zhang et al., 2024) (H100 GPU).
+Attention Eng.Sum Eng.QA Eng.MC Code.Debug Math.Find Retr.PassKey Retr.Num Retr.KV Avg.
+Full-Precision 18.03 12.5 64.19 24.37 18.29 100.0 100.0 7.0 43.05
+FlashAttn3-fp8 19.03 11.73 55.90 22.59 22.57 100.0 100.0 0.4 41.53
+SageAttention2 18.17 12.46 64.19 25.63 17.43 100.0 100.0 6.6 43.06
+(a) Full Precision
+(b) FlashAttn3-fp8
+(c) SageAttention2
+Figure 18: Needle In A Haystack results on Llama-3-262k (8B).
+Figure 19: An example of quantized value distribution of $Q$ before and after smoothing $Q$.
 Table 15: Worst accuracy across all layers of CogvideoX using different quantization granularities.
@@
 Per-tensor 85.85% 0.4687 0.2261
-3.3 INT4 Per-thread Quantization
-SageAttention uses per-block quantization, which quantizes each block of $Q$ and $K$ for a GPU streaming processor.
-Such a quantization strategy could achieve an accuracy performance close to per-token quantization and avoid the overhead of the dot product of the quantization scale vectors $\delta_{Q}$ and $\delta_{K}$.
-However, quantizing $Q$ and $K$ to INT4 demands a more accurate quantization granularity.
-We propose per-thread quantization, a more precise and granular quantization approach than the per-block quantizer, also without the additional overhead of vectors dot product required for per-token quantization.
-Specifically, each block of $Q$, i.e., $Q_{i}$, in SageAttention will be split into $c_{w}$ segments and processed by $c_{w}$ GPU warps in a GPU streaming processor (SM).
-We call each segment of $Q_{i}$ as $Q_{w}$, and $k_{w}=K_{j}$ because a $K_{j}$ only corresponds to one warp.
-Then, each warp containing 32 threads uses the mma.m16n8k64 PTX instruction (NVIDIA, ) for the $Q_{w}K_{j}^{\top}$.
-According to the layout requirement of this instruction, we find that the $Q_{w}[8\times(n\%8)]$ could share one quantization scale, and $K_{j}[8\times(n\%8)]$ and $K_{j}[8\times(n\%8+1)]$ could share one quantization scale, where $n$ is the token index.
-Such a quantization method is more fine-grained and with no additional overhead.
-This is because it assigns different GPU threads to distinct quantization groups based on the MMA instruction layout, with each thread performing dequantization only using a single quantization scale value.
-The per-thread quantization process can be formulated as follows.
-$i_{\delta q}=$ $\lfloor(n*8*c_{w}/b_{q})\rfloor$
-$q_{i}[i_{\delta q}]=$$\{8\times(n\%8)+\lfloor(n*c_{w}/b_{q})\rfloor*b_{q}/c_{w}\},~{}~{}n\in[0,N]$
-$\delta_{Q}[i_{\delta q}]=$ $\frac{\max(\left|\,Q[q_{i}[i_{\delta q}]]\,\right|)}{7}$
-$\hat{Q}[q_{i}[i_{\delta q}]]=$ $\left\lceil\frac{Q[q_{i}[i_{\delta q}]]}{\delta_{Q}[i_{\delta q}]}\right\rfloor$ (5)
-$i_{\delta k}=$ $\lfloor(n*4/b_{k})\rfloor$
-$k_{n}[i_{\delta k}]=$$\{8\times(n\%8)+\lfloor n/b_{k}\rfloor*b_{k}\}\cup\{8\times(n\%8)+1+\lfloor n/b_{k}\rfloor*b_{k}\},~{}~{}n\in[0,N]$
-$\delta_{K}[i_{\delta k}]=$ $\frac{\max(\left|\,K[k_{n}[i_{\delta k}]]\,\right|)}{7}$
-$\hat{K}[k_{n}[i_{\delta k}]]=$ $\left\lceil\frac{K[k_{n}[i_{\delta k}]]}{\delta_{K}[i_{\delta k}]}\right\rfloor$ (6)
-We show an example of per-thread quantization in Figure 6.
-Algorithm 1 Implementation of SageAttention2.
-  Input: Matrices $Q(\text{FP16}),K(\text{FP16}),V(\text{FP16})\in\mathbb{R}^{N\times d}$, block size $b_{q},b_{kv}$, warp count $c_{w}$.
-  Preprocessing: $K=K-\mathrm{mean}(K)$, $\overrightarrow{Q}_{m}=\mathrm{mean}(Q)$, $Q=Q-\overrightarrow{Q}_{m}$, $\Delta S=\text{GEMV}(\overrightarrow{Q}_{m},K^{\top})$.
-  Quantization: ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{Q},\hat{Q})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{Q}}}(Q)$, ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{K},\hat{K})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{K}}}(K)$.
-//per-thread.
-${}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{V},\hat{V})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{V}}}(V)$.
-//per-channel.
-  Divide $\hat{Q}$ to $T_{m}={N}/{b_{q}}$ blocks $\{\hat{Q}_{i}\}$; divide $\hat{K}$, $\hat{V}$, and $\Delta S$ to $T_{n}={N}/{b_{kv}}$ blocks $\{\hat{K}_{i}\}$, $\{\hat{V}_{i}\}$, and $\{\Delta S_{i}\}$;
-  for $i=1$ to $T_{m}$ do
-     Load $\hat{Q}_{i}$ and $\delta_{Q}=\delta_{Q}[c_{w}*8i:c_{w}*8(i+1)]$ into an SM ;
-     for j in [1, $T_{n}$do
-        Load $\hat{K}_{j}$, $\hat{V}_{j}$, and $\delta_{K}=\delta_{K}[4j:4(j+1)]$ into the SM ;
-        $w=\text{range}(c_{w}),st=w*c_{w}$
-        $S_{i}^{j}[st:st+c_{w}]=\psi^{-1}_{\delta_{Q}\delta_{K}}(\mathrm{Matmul}(\hat{Q}_{i}[st:st+c_{w}],\hat{K}_{j}^{\top}))+\Delta S_{j}$; // Paralleled by $c_{w}$ warps.
-The dequantization $\psi^{-1}_{\delta_{Q}\delta_{K}}$ is illustrated in Figure 6.
-        $m_{i}^{j}=\mathrm{max}(m_{i}^{j-1},\mathrm{rowmax}(S_{i}^{j}))$, ${}\widetilde{P}_{i}^{j}=\mathrm{exp}(S_{i}^{j}-m_{i}^{j})$, ${}l_{i}^{j}=e^{m_{i}^{j-1}-m_{i}^{j}}+\mathrm{rowsum}(\widetilde{P}_{i}^{j})$ ;
-        $O_{i}^{j}$(FP22) = $\mathrm{Matmul}((\widetilde{P}_{i}^{j}*448)\text{.to(FP8.e4m3)},V_{j})$ ;
-        $O_{i}^{j}$(FP32) = $\mathrm{diag}(e^{m_{i}^{j-1}-m_{i}^{j}})^{-1}O_{i}^{j-1}(FP32)+O_{i}^{j}$(FP22) ;
-     end for
-     Load $\delta_{V}$ into an SM;
-[…]
-     Write $O_{i}$ ;
-  end for
-  $O=\{O_{i}\}$
-  return $O$
-Table 4: Error of the FP8 Matmul instruction of mma(f8f8f32).
-Precision of Accumulated Value E8M13 E8M23
-Error compared to FP32 0 ||||mma(f32.f16.f16.f32) - mma(f32.f8.f8.f32)||||
-3.4 FP32 MMA Buffer for FP22 Accumulator
-To avoid the drawback of SageAttention where FP16 Matmul with an FP16 accumulator for $PV$ is only effective on GPUs such as the RTX4090, we adopt the approach of quantizing $P$ and $V$ to FP8 to leverage the universal acceleration of FP8 Tensor cores.
-However, we discover that the accumulator for the mma(f32f8f8f32) instruction on the Ada architecture is actually FP22, specifically with 1 sign bit, 8 exponent bits, and 13 mantissa bits.
-Specifically, for mma(f32f8f8f32) instruction $C=AB+D$, where $A,B$ are tensors in FP8 data type and $C,D$ are tensors with FP32 data type, we initialize the $A,B$ to zero and vary $D$ to test the data type of the accumulator.
-As shown in Table 4, when $D$ is initialized with 1 sign bit, 8 exponent bits, and 13 mantissa bits, the value of $C$ matches the result of the mma(f16f16f32) instruction.
-However, when $D$ is initialized with more than 13 mantissa bits, the error of $C$ corresponds to the difference between the results of mma(f32f16f16f32) and mma(f32f8f8f32).
-Consequently, the matrix multiplication of $\widetilde{P}V$, quantized to FP8, incurs a certain degree of accuracy loss compared to using an FP32 accumulator.
-To mitigate this accuracy loss as much as possible, we propose to use an FP32 buffer to accumulate the values of $\widetilde{P}_{i}^{j}V_{j}$ in FP22.
-Specifically, as shown in Algorithm 1, after each Matmul between the tiled $\widetilde{P}$ and $V$, we store the Matmul result in the FP32 buffer.
-This approach prevents the FP22 accumulator from accumulating errors across the entire sequence, confining errors to the tiled range instead.
-Table 5: Average accuracy using different data types of $(\widetilde{P},V)$ across all layers of a CogvideoX model, where $(Q,K)$ are smoothed.
-$Q,K$ $\widetilde{P},V$ Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
-INT4 INT8 77.05% 0.5618 0.5044
-E5M2 99.20% 0.0905 0.0903
-E4M3 99.44% 0.0683 0.0347
-FP16 99.45% 0.0649 0.0335
 Table 16: Worst accuracy using different data types of $(\widetilde{P},V)$ across all layers of a CogvideoX model, where $(Q,K)$ are smoothed.
@@
 FP16 96.76% 0.1916 0.0775
-Table 7: Two kernel implementations of SageAttention2.
-      Kernel       $\psi_{Q}(Q)$, $\psi_{K}(K)$       $\psi_{P}(\widetilde{P})$, $\psi_{V}(V)$
-      SageAttn2-4b       INT4, per-thread       FP8, per-block;  FP8, per-channel
-      SageAttn2-8b       INT8, per-thread       FP8, per-block;  FP8, per-channel
-3.5 Quantization for Q, K, P, V
-Quantization for $Q,K$.
-We propose $\psi_{Q}(Q)$ and $\psi_{K}(K)$ in granularity of per-thread.
-This is first because per-channel quantization is not feasible since the scale factors of the inner axis of $QK^{\top}$ cannot used to do dequantization (Xiao et al., 2023a).
-Second, as shown in Table 3 and Table 3, we compare the average and worst-case accuracies of INT4 quantization at per-token, per-thread, per-block, and per-tensor granularity using real $Q,K,V$ across all layers of CogvideoX.
-Results indicate that the accuracy of per-thread quantization is very close to per-token and outperforms a lot more than per-block and per-tensor.
-Furthermore, per-thread quantization incurs less dequantization overhead than per-token, as discussed in Section 3.3.
-Quantization for $\widetilde{P},V$.
-We choose FP8, specifically the E4M3 data type, for $\psi_{P}(\widetilde{P})$ and $\psi_{V}(V)$ for two reasons.
-First, most GPUs have tensor cores that support FP8 Matmul operations, which are twice as fast as those using FP16.
-Second, Table 6 and Table 6 show the average and worst accuracy of different data types used for $\widetilde{P},V$ using real $Q,K,V$ across all layers of CogvideoX.
-We can see that the accuracy of using E4M3 is very close to that of using FP16 and superior to that of E5M2 and INT8.
-We propose to use $\psi_{P}(\widetilde{P})$ in per-block and $\psi_{V}(V)$ in per-channel for three reasons.
-First, per-channel quantization for $\widetilde{P}$ and per-token quantization for $V$ are not viable because dequantization requires scale factors of outer axis.
-Second, $\widetilde{P}=\mathrm{exp}(S_{i}-\mathrm{rowmax}(S_{i}))$, where $S_{i}$ is the Matmul result of a block of $Q$ and $K^{\top}$, the max value in each row of $\widetilde{P}$ is 1.
-Hence, we can assign a single static scale $s=\frac{1}{448}$ to a block $\widetilde{P}$.
-Third, per-channel quantization can address the channel-wised outlier of $V$.
-Accuracy metrics.
-We use three metrics to assess the accuracy of quantized attention output $O^{\prime}$ compared to attention output in full-precision $O$: First, we flatten $O^{\prime}$ and $O$ into vectors in the shape of $1\times n$.
-Then, Cosine similarity: $CosSim=\sum OO^{\prime}/\smash{\sqrt{\sum O^{2}}}\smash{\sqrt{\sum O^{\prime 2}}}$, Relative L1 distance: $L1=\sum|O-O^{\prime}|/\sum|O|$, Root mean square error: $RMSE=\sqrt{(1/n)\sum(O-O^{\prime})^{2}}$.
-Figure 7: An example of dot product precison a row of $\widetilde{P}$ and a column of $V$ presented by FP22 data type.
-3.6 An Optional Technique: Smooth V
-We also figure out another way to mitigate the accuracy loss due to the FP22 accumulator when $V$ possesses channel-wise biases:
-$\gamma(V)=V-\mathrm{mean}(V),$ $~{}~{}~{}~{}\mathrm{mean}(V)=\overrightarrow{V_{m}}$
-$\phi_{V}(V)=$ $\psi_{V}\circ\gamma,~{}\overrightarrow{V_{m}}$ (7)
-As shown in Figure 7, this strategy could enhance the accuracy of FP22 for values in $\widetilde{P}V$ for the following reasons: Each row of $\widetilde{P}$ spans a value range from 0 to 1, and each column of $V$ consistently features channel-wise biases that are exclusively positive or negative, for instance, ranging between 8 and 9.
-Consequently, the values of $\widetilde{P}V$ could be quite large.
-However, the floating-point number representation range is not uniform—it is denser near zero.
-Therefore, by subtracting the mean along the channel dimension from $V$, the values of $\widetilde{P}V$ will be closer to zero, resulting in higher representational precision.
-We believe such an explanation and strategy could address lots of issues [^1] https://github.com/triton-lang/triton/issues/4476, https://github.com/triton-lang/triton/issues/5065 reported by the community.
-Additionally, to maintain the correctness of the attention computation, it is only necessary to add $\overrightarrow{V_{m}}$ to the final calculation of $O$: $O=O+\overrightarrow{V_{m}}$.
-This is because the sum of each row of the $\widetilde{P}$ matrix equals 1, so $\widetilde{P}\overrightarrow{V_{m}}=\overrightarrow{V_{m}}$.
-In other words, this method decomposes $V$ into two parts: $\overrightarrow{V_{m}}$ and $V$.
-For $V$, it centers the values of each column around zero, which leads to the dot product result between a row from the quantized $\widetilde{P}$ matrix and a column from the quantized $V$ matrix being closer to zero.
-This makes the representation of FP22 more accurate.
-Meanwhile, $\overrightarrow{V_{m}}$ is retained in FP16 and added to $O$ at the end without causing a loss of precision for the $\overrightarrow{V_{m}}$ part.
-Remark.
-This technique is optional and not employed in our main experiments.
-This is because its benefits are marginal, only significantly enhancing attention accuracy primarily when $V$ exhibits channel-wise outliers, which are absent in some models, such as $V$ in Llama3.1 (see Figure 4).
-Figure 8: Speed comparison between SageAttention2 and baselines (RTX4090, headdim=64).
-Figure 9: Speed comparison between SageAttention2 and baselines (RTX4090, headdim=128).
-Figure 10: Speed comparison between SageAttention2 and baselines (L20, headdim=64).
-Figure 11: Speed comparison between SageAttention2 and baselines (L20, headdim=128).
-4 Experiment
-4.1 Setup
-Models.
-We validate the effectiveness of SageAttention2 across a diverse set of representative models from language, image, and video generation.
-Specifically, we conduct experiments on ten models: Llama2 (7B) (Touvron et al., 2023), Llama3.1 (8B) (Dubey et al., 2024), and GLM4 (9B) (GLM et al., 2024) for text2text, CogvideoX (2B), CogvideoX (1.5-5B) (Yang et al., 2024), HunyuanVideo (Weijie Kong & Jie Jiang, 2024), and Mochi (Team, 2024) for text2video, Flux (schnell) (Black Forest Labs, 2023) and Stable-Diffusion3.5 (turbo) (Stability AI, 2023) for text2image, and TIMM (Wightman, 2019) for image classification.
-Datasets.
-Text-to-text models are evaluated on four zero-shot tasks: WikiText (Merity et al., 2022) to assess the model’s prediction confidence, LAMBADA (Paperno et al., 2016) evaluate contextual understanding, MMLU (Hendrycks et al., 2020) for measuring knowledge across various subjects, and Longbench (Bai et al., 2023) for evaluating bilingual, multitask, and comprehensive assessment of long context understanding capabilities.
-Text-to-video models are evaluated using the open-sora (Zheng et al., 2024) prompt sets.
-Text-to-image models are assessed on MJHQ-30K (Li et al., 2024).
-TIMM is evaluated on on three image datasets: ImageNet (Deng et al., 2009), ImageNet-Sketch (Sketch) (Wang et al., 2019), and ImageNet-Rendition (ImageNet-r) (Hendrycks et al., 2021).
-Metrics.
-For text-to-text models, we use perplexity (ppl.) (Jelinek et al., 1977) for WikiText, Accuracy (Acc.) for LAMBADA and MMLU, and Longbench score (Bai et al., 2023).
-For text-to-video models, following Zhao et al.
-(2024), we evaluate the quality of generated videos on five metrics: CLIPSIM and CLIP-Temp (CLIP-T) (Liu et al., 2024) to measure the text-video alignment; (VQA-a) and (VQA-t) to assess the video aesthetic and technical quality, respectively; and Flow-score (FScore) for temporal consistency (Wu et al., 2023).
-For text-to-image models, generated images are compared with the images in MJHQ-30K dataset in three aspects: FID (Heusel et al., 2017) and sFID (Salimans et al., 2016) for fidelity evaluation, Clipscore (CLIP) (Hessel et al., 2021) for text-image alignment, and ImageReward (IR) (Xu et al., 2024) for human preference.
-For TIMM, we use classification accuracy.
-Implemetation details. We implement our attention kernels using CUDA and conduct experiments on Ubuntu 22.04 servers with RTX4090 and L20 GPUs.
-Baselines.
-(1) SmoothAttn.
-Following Qserve (Lin et al., 2024), we apply smooth quant for matrixes $Q,K$ with smoothing factor $\alpha=0.5$.
-(2) HadmdAttn.
-Following Quarot (Ashkboos et al., 2024), we apply random Hadamard transformation for matrixes $Q,K$ before INT4 quantization.
-(3) SageAttention (Zhang et al., 2024) uses smoothing $K$, INT8 per-block quantization for $Q,K$, and FP16 for $\widetilde{P},V$.
-Figure 12: Comparison examples from HunyuanVideo, prompts are sampled from open-sora prompt sets.
-Figure 13: Comparison examples from CogvideoX (2B), prompts are sampled from open-sora prompt sets.
-Table 8: Average accuracy across all layers of CogvideoX using different smooth methods.
-Method CosSim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
-None 80.04% 0.3906 0.2223
-HadmdAttn 79.77% 0.3782 0.2180
-SmoothAttn 90.21% 0.3383 0.1952
-Smooth K (Ours) 98.07% 0.1493 0.0743
-Smooth Q (Ours) 98.30% 0.1250 0.0712
-SageAttn2-4b 99.46% 0.0648 0.0334
 Table 17: Worst accuracy across all layers of CogvideoX using different smooth methods.
@@
 SageAttn2-4b 96.71% 0.1956 0.0779
-Table 10: End-to-end metrics across text and video generation models. - means the model does not support this dataset evaluation. ✗ indicates an inability to generate results for evaluation.
-Model Attention WikiText (Ppl.) $\downarrow$ Lambda (Acc.) $\uparrow$ MMLU (Acc.) $\uparrow$ Longbench $\uparrow$
-Llama2 Full-Precision 5.823 0.886 0.439 -
-HadmdAttn 6.771 0.860 0.360 -
-SmoothAttn 6.717 0.867 0.392 -
-SageAttn2-4b 5.912 0.881 0.428 -
-SageAttn2-8b 5.828 0.886 0.438 -
-SageAttention 5.824 0.887 0.439 -
-Llama3.1 Full-Precision 6.013 0.815 0.635 49.40
-HadmdAttn 7.872 0.762 0.500 44.07
-SmoothAttn 7.180 0.783 0.541 44.69
-SageAttn2-4b 6.256 0.798 0.607 48.79
-SageAttn2-8b 6.019 0.811 0.634 49.59
-SageAttention 6.017 0.812 0.634 49.55
-GLM4 Full-Precision 7.241 0.432 0.743 49.78
-HadmdAttn 7.989 0.435 0.669 45.97
-SmoothAttn 8.943 0.449 0.592 42.20
-SageAttn2-4b 7.352 0.433 0.725 49.23
-SageAttn2-8b 7.242 0.432 0.745 49.60
-SageAttention 7.243 0.433 0.744 49.79
-Model Attention CLIPSIM $\uparrow$ CLIP-T $\uparrow$ VQA-a $\uparrow$ VQA-t $\uparrow$ FScore $\uparrow$
-CogvideoX (2B) Full-Precision 0.1836 0.9975 77.605 75.360 3.006
-HadmdAttn 0.1742 0.9877 29.780 23.985 0.499
-SmoothAttn 0.1741 0.9870 41.703 47.043 0.624
-SageAttn2-4b 0.1821 0.9973 77.368 74.906 2.603
-SageAttn2-8b 0.1829 0.9977 76.532 74.281 2.941
-SageAttention 0.1833 0.9976 76.997 71.360 2.988
-CogvideoX (1.5-5B) Full-Precision 0.1778 0.9979 70.231 70.928 2.507
-HadmdAttn 0.1576 0.9933 8.990 2.299
-SmoothAttn 0.1559 0.9950 8.812 2.277
-SageAttn2-4b 0.1721 0.9978 57.729 52.989 2.884
-SageAttn2-8b 0.1775 0.9980 69.492 74.415 2.487
-SageAttention
-HunyuanVideo Full-Precision 0.1783 0.9995 82.516 75.934 0.604
-HadmdAttn 0.1727 0.9989 7.514 0.762 0.175
-SmoothAttn 0.1739 0.9988 6.987 0.609 0.148
-SageAttn2-4b 0.1751 0.9995 81.478 65.371 0.610
-SageAttn2-8b 0.1782 0.9996 81.786 75.354 0.586
-SageAttention 0.1786 0.9995 82.496 79.843 0.597
-Mochi Full-Precision 0.1798 0.9986 45.549 65.416 1.266
-HadmdAttn 0.1733 0.9980 9.053 25.133 0.704
-SmoothAttn 0.1687 0.9978 3.383 3.480 0.241
-SageAttn2-4b 0.1783 0.9986 35.955 43.735 1.137
-SageAttn2-8b 0.1797 0.9986 46.760 64.901 1.255
-SageAttention 0.1800 0.9987 48.707 63.763 1.269
-Table 11: Table 10 continued. End-to-end metrics across image generation and classification models.
-        Model         Attention        FID $\downarrow$        sFID $\downarrow$        CLIP $\uparrow$        IR $\uparrow$
-        Flux         Full-Precision        10.960        16.648        26.180        1.009
-        HadmdAttn        11.353        18.495        26.123        0.965
-        SmoothAttn        11.149        19.017        26.109        0.959
-        SageAttn2-4b        10.577        17.497        26.141        0.998
-        SageAttn2-8b        10.927        16.723        26.175        1.009
-        SageAttention        10.944        16.641        26.171        1.008
-               Stable-Dif        fusion3.5         Full-Precision        14.105        15.646        25.505        0.902
-        HadmdAttn        14.259        15.909        25.513        0.886
-        SmoothAttn        14.161        15.649        25.510        0.887
-        SageAttn2-4b        14.097        15.397        25.487        0.895
-        SageAttn2-8b        14.106        15.647        25.499        0.901
-        SageAttention        14.140        15.678        25.503        0.902
-Model Attention ImageNet (Acc.) $\uparrow$ Sketch (Acc.) $\uparrow$ ImageNet-r (Acc.) $\uparrow$
-TIMM Full-Precision 84.79% 45.32% 59.55%
-HadmdAttn 84.50% 44.89% 58.80%
-SmoothAttn 84.40% 44.68% 58.73%
-SageAttn2-4b 86.67% 45.24% 59.29%
-SageAttn2-8b 84.79% 45.39% 59.57%
-SageAttention 84.74% 45.38% 59.95%
-4.2 Speed and Accuracy of Kernels
-Speed.
-We conduct experiments to compare the Speed of SageAttention2 against baselines using configurations with headdim=64 and headdim=128, both with and without Causal Mask (Vaswani, 2017).
-Specifically, Figure 8 and Figure 9, show the speed of SageAttention2 and baselines across varying sequence lengths on RTX4090.
-These results indicate that SageAttn2-4b and SageAttn2-8b are approximately 3x and 2.7x faster than FlashAttention2, and about 4.5x and 4x faster than xformers, respectively.
-Accuracy.
-Table 9 and Table 9 show that the average and worst accuracy of different methods with INT4 $Q,K$ and FP8 $P,V$ across all layers of CogvideoX.
-The results indicate that the accuracy of SageAttn2-4b is superior to other baselines.
-Table 12: End-to-end generation latency using SageAttention2 (The latency of Llama3.1 is the time to first token generation using different sequence lengths).
-Model GPU Original SageAttn2-8b SageAttn2-4b
-CogvideoX (2B) RTX4090 86 s 54 s 52 s
-CogvideoX (1.5-5B) RTX4090 1040 s 577 s 555 s
-HunyuanVideo L20 2221 s 1486 s 1435 s
-Mochi L20 2336 s 1316 s 1190 s
-Llama3.1 (48K tokens) RTX4090 9.2 s 5.7 s 5.6 s
-Llama3.1 (100K tokens) L20 39.9 s 25.4 s 23.2 s
-4.3 End-to-end Performance
-Metrics loss.
-We assessed the end-to-end metrics of various models using SageAttention2 compared to using attention in full precision.
-Detailed evaluation results are presented in Table 10 for Llama2, Llama3.1, GLM4, CogvideoX (2B), CogvideoX (1.5-5B), HunyuanVideo, Mochi, Flux, Stable-Diffusion3.5, and TIMM, respectively.
-The results indicate that SageAttn2-4b outperforms all baselines and maintains most of the end-to-end accuracy across all models.
-Additionally, SageAttn2-8b incurs almost no metrics loss across various models.
-Visible examples.
-Figure12 and Figure 13 show some visible comparision examples from HunyuanVideo and CogvideoX (1.5-5B).
-We can observe that SageAttn2-8b does not introduce any visible differences compared to full-precision attention, whereas SageAttn2-4b has minor differences compared to full-precision attention but is much better than the baselines.
-End-to-end speedup.
-We compared the original generation latency and the latency using SageAttention2 for models with long sequence lengths in Table 12, observing significant speedup effects.
-For instance, SageAttention2 achieved a 1.8x speedup in CogvideoX (1.5-5B) without any metrics loss (SageAttn2-8b).
-SageAttn2-4b further accelerated these models but with a little metrics loss.
-Table 13: An accuracy example on real tensors of CogvideoX model with or without smoothing $V$.
-Smooth V Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
-98.25% 0.1980 0.2387
-99.75% 0.0406 0.0773
-4.4 Ablation Study
-Overhead of smoothing Q, K.
-Note that the overhead of smoothing $Q,K$ only contains $Q-\overrightarrow{Q_{m}}$, $K-\mathrm{mean}(K)$, and $\overrightarrow{Q_{m}}K^{\top}$.
-The speed overhead of smoothing $Q,K$ accounts for about 3.5% of the attention kernel.
-Benefit of smoothing V.
-Table 13 shows the attention accuracy on real tensors sampled from CogvideoX with and without smoothing $V$.
-It demonstrates that smoothing V could improve the accuracy of SageAttention2 when quantizing $Q,K$ to INT4 and $\widetilde{P},V$ to FP8.
-5 Conclusion and Future Work
-We introduce SageAttention2, an efficient and accurate quantization method for attention.
-First, we propose to quantize matrixes $(Q,K)$ to INT4 in a thread-level granularity and quantize matrixes $(\widetilde{P},V)$ to FP8.
-Second, we propose a method to smooth $Q$, enhancing the accuracy of INT4 $QK^{\top}$.
-Third, we propose to use an FP32 Matmul buffer for $PV$ to enhance the accuracy of FP8 $\widetilde{P}V$.
-Our implementation is faster than FlashAttention2 and xformers by approximately 3x and 5x on RTX4090, respectively.
-Extensive testing confirms that our approach maintains end-to-end metrics across various models, including language, image, and video generation.
-Future Work. We leave the implementation of FP8 MatMul with an FP16 accumulator for $\widetilde{P}V$ and SageAttention2 on the Hopper architecture for future work.
+Table 18: Overhead of per-thread quantization, smoothing Q, and two-level accumulation techniques measured on L20 GPU.
+      Method       TOPS
+      Attention (INT4 + FP8)       284
+      + Per-thread quantization       283
+      + Two-level accumulation       283
+      + Smoothing Q       273
+A.7 Additional Experiments and Analysis
+Additional Results. Table 11,  12 and  13 show results of SageAttention2 and other baselines on Llama2 (7B), CogvideoX (2B) and TIMM.
+Results of Super-Long Context.
+We further conduct experiments on super-long context using Llama-3-262k (8B) [^1] https://huggingface.co/gradientai/Llama-3-8B-Instruct-262k on InfiniBench (Zhang et al., 2024) and Needle-in-a-Haystack (NIAH) (Kamradt, 2023), with sequence lengths reaching up to 262k tokens on an H100 GPU.
+Results are shown in Table 14 and Fig 18.
+SageAttention2 maintains model performance even under super-long context, while FlashAttention3(fp8) suffers from end-to-end accuracy degradation.

### v7

@@
 Third, we propose a two-level accumulation strategy for $\widetilde{P}V$ to enhance the accuracy of FP8 $\widetilde{P}V$.
-The operations per second (OPS) of SageAttention2 surpass FlashAttention2 and xformers by about 3x and 4.5x on RTX4090, respectively.
+The operations per second (OPS) of SageAttention2 surpass FlashAttention2 and xformers by about 3x and 4.5x.
 Moreover, SageAttention2 matches the speed of FlashAttention3(fp8) on the Hopper GPUs, while delivering much higher accuracy.
-Comprehensive experiments confirm that our approach incurs negligible end-to-end metrics loss across diverse models—including those for language, image, and video generation.
-The code is available at https://github.com/thu-ml/SageAttention
+Comprehensive experiments confirm that our approach incurs negligible end-to-end metrics loss across diverse models, including those for language, image, and video generation.
+The code is available at https://github.com/thu-ml/SageAttention.
 Machine Learning, ICML
@@
 1 Introduction
-Due to the quadratic time complexity of attention, its efficient implementation becomes crucial as sequence lengths increase in real-world applications (Jiang et al., 2024).
-Several strategies have been developed to mitigate computational demands of attention —such as (1) Linear attention methods (Wang et al., 2020; Choromanski et al., 2021; Yu et al., 2022; Katharopoulos et al., 2020) that reduce complexity to $O(N)$ and (2) Sparse attention methods (Liu et al., 2021; Chu et al., 2021; Li et al., 2022; Xiao et al., 2024b, a; Chen et al., 2024; Jiang et al., 2024; Venkataramanan et al., 2024; Gao et al., 2024; Fu et al., 2024) that selectively process parts of the context — these methods are only suitable for a limited range of models and tasks.
-The widely used attention methods optimize attention by exploiting hardware capacities, such as FlashAttention (Dao et al., 2022), FlashAttention2 (Dao, 2024), FlashAttention3 (Shah et al., 2024), xformers (Lefaudeux et al., 2022), and SageAttention (Zhang et al., 2025).
+Due to the quadratic time complexity of attention, its efficient implementation becomes crucial as sequence lengths increase in real-world applications (Jiang et al., 2024; Zhang et al., 2025b).
+Several strategies have been developed to mitigate the computational demands of attention —such as (1) Linear attention methods (Wang et al., 2020; Choromanski et al., 2021; Yu et al., 2022; Katharopoulos et al., 2020) that reduce complexity to $O(N)$ and (2) Sparse attention methods (Liu et al., 2021; Chu et al., 2021; Li et al., 2022; Xiao et al., 2024b, a; Chen et al., 2024; Jiang et al., 2024; Venkataramanan et al., 2024; Gao et al., 2024; Fu et al., 2024; Zhang et al., 2025e; Xi et al., 2025; Yang et al., 2025a; Zhang et al., 2025f, h, i) that selectively process parts of the context — these methods are only suitable for a limited range of models and tasks.
+The widely used attention methods optimize attention by exploiting hardware capacities, such as FlashAttention V1, V2, V3 (Dao et al., 2022; Dao, 2024; Shah et al., 2024), xformers (Lefaudeux et al., 2022), and SageAttention (Zhang et al., 2025c, d, g, a).
 These works do not omit computations across the entire sequence and achieve impressive speed and accuracy performance across various tasks.
@@
 For the two matrix multiplication (Matmul) operations in attention: $QK^{\top}$ and $\widetilde{P}V$, SageAttention accelerates them by quantizing the $QK^{\top}$ to INT8 and uses FP16 Matmul with FP16 accumulators for $\widetilde{P}V$.
-Moreover, to keep the accuracy of attention, SageAttention proposes smoothing $K$ by eliminating its channel-wise outliers.
-SageAttention achieves 2 $\times$ speedup than FlashAttention2 and is the first quantized attention that incurs negligible end-to-end metrics loss across language, image, and video generation models.
+Moreover, to maintain attention accuracy, SageAttention proposes smoothing $K$ by eliminating its channel-wise outliers.
+SageAttention achieves 2 $\times$ speedup compared with FlashAttention2 and is the first quantized attention that incurs a negligible end-to-end metrics loss across language, image, and video generation models.
 However, SageAttention has two weaknesses.
@@
 For example, when only per-tensor quantizing $Q,K$ to INT4, the text-to-video model CogvideoX will generate a completely blurry video, and Llama3 only achieves a random-guessing-level accuracy of 25% on MMLU.
-After investigating deeply, we identified three primary challenges: (C1) The numerical range of INT4 is quite limited compared to FP16 or INT8, leading to significant quantization errors when $Q$ and $K$ have some abnormal values.
+After an in-depth investigation, we identified three primary challenges: (C1) The numerical range of INT4 is quite limited compared to FP16 or INT8, leading to significant quantization errors when $Q$ and $K$ have some abnormal values.
 (C2) We discover that the FP32 accumulator designed for FP8 matrix multiplication in the tensor core (mma.f32.f8.f8.f32) is actually FP22, specifically with 1 sign bit, 8 exponent bits, and 13 mantissa bits.
@@
 Simultaneously, to avoid the extra latency caused by per-token dequantization, where each GPU thread corresponds to multiple quantization scales, we propose an adequately precise quantization method that incurs no additional dequantization overhead.
-Specifically, we introduce a per-thread quantization approach based on the mapping between the GPU threads and memory layout of matrices as dictated by the PTX mma instructions.
+Specifically, we introduce a per-thread quantization approach based on the mapping between the GPU threads and the memory layout of matrices as dictated by the PTX mma instructions.
 This method groups tokens corresponding to the same thread for quantization and dequantization, ensuring that each thread is associated with a single quantization scale.
@@
 Then, we add $\overrightarrow{Q}_{m}K$ after the $QK^{\top}$ Matmul to ensure the correctness of the attention computation.
-To address (C2), the accuracy loss from the 22-bit accumulator for FP8 Matmul of $\widetilde{P}V$, we propose a two-level accumulation strategy that uses an FP32 buffer to accumulate the values from the 22-bit accumulator after each block Matmul of $\widetilde{P}V$.
+To address (C2), the accuracy loss caused by the 22-bit accumulator for FP8 Matmul of $\widetilde{P}V$, we propose a two-level accumulation strategy that uses an FP32 buffer to accumulate the values from the 22-bit accumulator after each block Matmul of $\widetilde{P}V$.
 This method confines the errors to the block range.
@@
 Performance.
-Importantly, we offer a high-performance implementation of SageAttention2 on RTX4090 and L20 GPUs.
+We offer a high-performance implementation of SageAttention2 on RTX4090 and L20 GPUs.
 This implementation achieves a peak performance of 481 TOPS on the RTX4090, outperforming FlashAttention2 and xformers by approximately 3x and 4.5x, respectively.
-Note that FlashAttention3 is tailored to and can only be used with the Nvidia Hopper architecture.
-Moreover, SageAttention2 matches the speed of FlashAttention3(fp8) on the Hopper GPUs, while delivering much better accuracy.
-For example, on CogVideoX-1.5 (Yang et al., 2025), Mochi (Team, 2024) and HunyuanVideo (Kong et al., 2024), our method does not compromise end-to-end accuracy, whereas videos generated using FlashAttention3(fp8) often suffer noticable degradation, as visualized in Fig.
-8.
-We extensively evaluate the end-to-end metrics of state-of-the-art text, image, and video generation models using SageAttention2.
+To support NVIDIA Hopper GPUs, which lack native INT4 tensor core support, we additionally provide SageAttention2-8b, a variant that quantizes $Q,K$ to INT8.
+FlashAttention3, in contrast, is tailored to and only compatible with the Hopper architecture.
+Moreover, SageAttention2-8b matches the speed of FlashAttention3(fp8) on Hopper GPUs, while delivering much better accuracy.
+For example, on popular video generation models, our method does not compromise end-to-end accuracy, whereas FlashAttention3(fp8) brings noticeable degradation, as visualized in Fig.
+7 and 9.
+We extensively evaluate the end-to-end metrics of state-of-the-art text, image, and video generation models.
 SageAttention2 can accelerate models in a plug-and-play way with negligible loss in end-to-end metrics.
@@
 The attention computation can be formulated as: $S=QK^{\top}/\sqrt{d},~P=\sigma(S),~O=PV$, where $\sigma(S)_{ij}=\exp(S_{ij})/\sum_{k}\exp(S_{ik})$.
-The matrices $Q$, $K$, and $V$ each has dimensionality $N\times d$, and $S$, $P$ are $N\times N$.
+The matrices $Q$, $K$, and $V$ each have dimensionality $N\times d$, and $S$, $P$ are $N\times N$.
 $d$ is typically small, e.g., 64 or 128, and $N$ can be thousands or millions.
@@
 $O_{ij}=$ $\mathrm{diag}\left(\exp(m_{i,j-1}-m_{ij})\right)O_{i,j-1}+\widetilde{P}_{ij}V_{j},$
-where $m_{ij}$ and $l_{ij}$ are $b_{q}$-dimenalional vectors, initialized with $\infty$- ∞ and 00 respectively.
-$\tilde{\sigma}(\cdot)$ is an online softmax operator: $m_{ij}=\max\{m_{i,j-1},\mathrm{rowmax}(S_{ij})\},~{}\widetilde{P}_{ij}=\exp(S_{ij}-m_{ij})$.
-Finally, the output is computed as $O_{i}=\mathrm{diag}(l_{i,n_{q}})^{-1}O_{i,n_{q}}$.
+[…]
 2.2 Quantization
 A matrix multiplication $C=AB$ can be accelerated with quantization as:
-$(\delta_{A},\hat{A})=\psi(A),~{}~{}(\delta_{B},\hat{B})=\psi(B),~{}~{}C=\psi^{-1}_{\delta_{A}\delta_{B}}(\hat{A}\hat{B})$
+[…]
 $\psi$ is a quantizer which converts a high-precision (e.g., FP32 or FP16) matrix $A$ to a low-precision format $\hat{A}$ (e.g., INT4 or FP8) with a scale $\delta_{A}$, and $\psi^{-1}$ is a dequantizer to convert back to high-precision.
 We should have $\psi^{-1}_{\delta_{A}}(\hat{A})\approx A$.
-The actual matrix multiplication $\hat{A}\hat{B}$ is carried out with low precision.
+The actual matrix multiplication $\hat{A}\hat{B}$ is performed with low precision.
 In modern GPUs, low-precision matrix multiplication is usually multiple times faster than higher-precision ones.
@@
 GPU MM Input MM Accumulator Speedup
-RTX4090 FP16 FP16 2x
+RTX4090
+FP16 FP16 2x
 FP8 FP32 2x
-L40, L20 H100 FP16 FP16 1x
+L40, L20
+H100
+FP16 FP16 1x
 FP8 FP32 2x
 2.3 SageAttention
-Based on the block tiling in FlashAttention (Dao et al., 2022), SageAttention (Zhang et al., 2025) quantizes $Q,K$ to INT8 in a per-block granularity, i.e., each $Q_{i},K_{i}$ has a separate scalar scale: $\delta_{Q_{i}}=\max(|{Q_{i}|})/127,\delta_{K_{j}}=\max(|{K_{j}|})/127$.
+Based on the block tiling in FlashAttention (Dao et al., 2022), SageAttention (Zhang et al., 2025c) quantizes $Q,K$ to INT8 in a per-block granularity, that is, each $Q_{i},K_{i}$ has a separate scalar scale: $\delta_{Q_{i}}=\max(|{Q_{i}|})/127,\delta_{K_{j}}=\max(|{K_{j}|})/127$.
 In this way, the product $S_{ij}$ in Eq.
 (1) can be approximated as $S_{ij}\approx\hat{Q}_{i}\hat{K}_{j}^{\top}\times(\delta_{Q_{i}}\delta_{K_{j}}/\sqrt{d})$.
-To maintain the accuracy, SageAttention proposes a preprocessing technique to by subtracting the token-wise mean from $K$.
+To maintain accuracy, SageAttention proposes a preprocessing technique by subtracting the token-wise mean from $K$.
 Additionally, SageAttention keeps $\widetilde{P}_{ij}$ and $V_{j}$ in FP16, but utilizes an FP16 accumulator (rather than FP32) for computing the product $\widetilde{P}_{ij}V_{j}$.
-Reducing the accumulator precision can accelerate the matrix multiplication (MM) on RTX4090 GPU.
+Reducing accumulator precision can accelerate matrix multiplication (MM) on the RTX4090 GPU.
 However, other GPUs, such as L20, L40, or H100, do not exhibit this behavior, as shown in Table 1.
+Figure 2: Typical examples of tensors’ data distribution in attention.
 3 SageAttention2
@@
 The workflow of SageAttention2 is shown in Fig.
-2.
+3.
 We quantize $Q,K$ to INT4 and $\tilde{P},V$ to FP8 to maximize the efficiency and propose several techniques, including $QK$-smoothing, per-thread quantization, and two-level accumulation to preserve the accuracy, which we shall discuss in subsequent subsections.
-Figure 2: Workflow of SageAttention2. 1 Smooth Q,K,V. 2 A GEMV to obtain $\Delta S$. 3 Per-thread quantize Q,K and per-channel quantize V. 4 Perform the SageAttention2 kernel. 5 Correct the output.
-Figure 3: Typical examples of tensors’ data distribution in attention.
+Figure 3: Workflow of SageAttention2.
+1 Smooth $Q,K,V$.
+2 A GEMV to obtain $\Delta S$.
+3 Per-thread quantize $Q,K$ and per-channel quantize $V$.
+4 Perform the SageAttention2 kernel.
+5 Correct the output.
+Figure 4: An example of per-thread quantization.
+The left figure shows the correspondence between the quantization scales and the tokens in each GPU warp.
+The right figure shows the correspondence between quantization tokens and GPU threads in a MMA.m16n8k64 instruction, showing that each GPU thread only corresponds to one quantization scale in $\delta_{Q}$ and $\delta_{K}$ in dequantization.
 3.1 Smooth $Q$
@@
 The numerical range of INT4 is notably restrictive.
-This affects quantization due to the presence of outliers (Lin et al., 2024).
+This affects quantization due to the presence of outliers (Lin et al., 2025).
 Given the INT4 range [-7, +7], any element will be quantized to zero if it is more than 14 times (0.5 vs 7) smaller than the largest element in the group.
-Since outliers are much larger than other elements, it is quite likely that most non-outlier elements are quantized to zero, resulting in significant accuracy degradation.
+Since outliers are much larger than other elements, it is likely that many non-outlier elements are quantized to zero, leading to substantial accuracy degradation.
 Therefore, to keep the quantization accurate, we need to keep the largest element small, making the magnitude of elements as uniform as possible.
 Such technique is called smoothing.
-Here, we propose a smoothing technique inspired by SageAttention (Zhang et al., 2025).
-SageAttention observed that $Q,K$ for all tokens are actually highly similar, with only small variations between different tokens (see Fig.
-3).
+Here, we propose a smoothing technique inspired by SageAttention (Zhang et al., 2025c).
+SageAttention observed that $Q,K$ for all tokens are actually highly similar, with only small variations between different tokens (Fig.
+2 shows the heatmap distributions of the $Q$, $K$, and $V$ randomly sampled from Llama3.1 (Dubey et al., 2024) and CogvideoX (Yang et al., 2025b)).
 We propose to smooth $K$ as SageAttention does and further smooth $Q$ by subtracting a common mean of each block:
 $\gamma(Q_{i})=Q_{i}-\bar{q}_{i},~~~\gamma(K_{j})=K_{j}-\bar{k},$ (2)
-where $\bar{q}_{i}=\mathrm{mean}(Q_{i}),\bar{k}=\mathrm{mean}(K)$ are $1\times D$ vectors, the mean is conducted along the token axis, and $\bar{q}_{i},\bar{k}$ are broadcasted to tokens in a block and a tensor for subtraction, respectively.
+where $\bar{q}_{i}=\mathrm{mean}(Q_{i}),\bar{k}=\mathrm{mean}(K)$ are $1\times D$ vectors, the mean is conducted along the token axis, and $\bar{q}_{i},\bar{k}$ are broadcasted to tokens in a block and a tensor for subtraction.
 With the decomposion, we have $S_{ij}=Q_{i}K_{j}^{\top}=(\bar{q}_{i}+\gamma(Q_{i}))(\bar{k}+\gamma(K_{j}))^{\top}=\bar{q}_{i}\bar{k}^{\top}+\bar{q}_{i}\gamma(K_{j})^{\top}+\gamma(Q_{i})\bar{k}^{\top}+\gamma(Q_{i})\gamma(K_{j})^{\top}=\gamma(Q_{i})\gamma(K_{j})^{\top}+\Delta{S_{ij}}+b$.
@@
 (2), apply quantization $(\delta_{Q_{i}},\hat{Q}_{i})=\psi_{Q}(\gamma(Q_{i})),(\delta_{K_{j}},\hat{K}_{j})=\psi_{K}(\gamma(K_{j}))$, and compute $\Delta{S_{ij}}=\bar{q}_{i}\gamma(K_{j})^{\top}$.
-The smoothing, quantization, and the GEMV (general matrix-vector multiplication) for computing $\Delta S$ can be fused into a single kernel, which scans the off-chip $Q$ and $K$ only once.
+Smoothing, quantization, and GEMV (general matrix-vector multiplication) for computing $\Delta S$ can be fused into a single kernel, which reads the off-chip $Q$ and $K$ only once.
 (2) attention: execute the low-precision GEMM, dequantize, and add back the vector $\Delta S$: $S_{ij}=\psi^{-1}_{\delta_{Q_{i}}\delta_{K_{j}}}(\hat{Q}_{i}\hat{K}_{j}^{\top})+\Delta S_{ij}$.
-These operations are all done on chip, and the dequantization and vector addition only adds a marginal overhead compared to the expensive mma operation for MM.
+These operations are all done on chip, and the dequantization and vector addition only add a marginal overhead compared to the expensive mma operation for MM.
 Importantly, $\gamma(Q_{i}),\gamma(K_{j})$ are quantized rather than $Q_{i},K_{j}$.
-Since the smoothed matrices are much smaller in magnitude and contain less outliers, the quantization accuracy can be significantly improved.
+Since the smoothed matrices are much smaller in magnitude and contain fewer outliers, the quantization accuracy can be significantly improved.
+A theoretical analysis of the benefit of smoothing is included in Appendix A.5.
 Remark.
 Classical techniques to improve the activation-weight MM, such as per-channel quantization, or SmoothQuant (Xiao et al., 2023) are not applicable here for the query-key MM in attention.
-Per-channel quantization is impracticable for $Q,K$ because quantization must be conducted along the outer axis (token dimension) of $QK^{\top}$ (Xiao et al., 2023).
+Per-channel quantization cannot be applied to $Q,K$ because the quantization must be conducted along the outer axis (token dimension) of $QK^{\top}$.
 On the other hand, both $Q$ and $K$ have significant outliers, so trading the quantization accuracy between them with SmoothQuant cannot work effectively, as shown in Sec.
-4.1.
-Here, we utilize the unique token-similarity pattern in attention to derive an accurate dedicated quantization method for $Q$ and $K$.
+4.
+Here, we utilize the unique token similarity pattern in attention to derive a dedicated quantization method for $Q$ and $K$.
 The previous work SageAttention only smooths $K$, so it is less accurate than our method.
@@
 We can find that with smoothing $Q$, the range of INT4 is utilized more uniformly and fully.
-Table 5 presents end-to-end metrics for different quantization methods with and without smoothing Q+K on Llama3.1 (Dubey et al., 2024) and CogvideoX (2b) (Yang et al., 2025).
+Table 5 presents end-to-end metrics for different quantization methods with and without smoothing Q+K on Llama3.1 and CogvideoX (2b).
 The results demonstrate that smoothing Q+K offers significant accuracy benefits.
-Also, Table 4 and 17 show that the order of effectiveness is smoothing Q+K $>$ smoothing Q $>$ smoothing K $>$ other baselines.
-Figure 4: An example of per-thread quantization.
-The left figure shows the correspondence between the quantization scales and the tokens in each GPU warp.
-The right figure shows the correspondence between quantization tokens and GPU threads in a MMA.m16n8k64 instruction, showing that each GPU thread only corresponds to one quantization scale in $\delta_{Q}$ and $\delta_{K}$ in dequantization.
+Also, Table 4 and 17 show that the order of effectiveness is smoothing Q+K >> smoothing Q >> smoothing K >> other baselines.
 3.2 INT4 Per-thread Quantization
-Orthogonal to smoothing, we can mitigate the problem of outliers by refining the quantization granularity so the number of affected elements by outliers becomes smaller.
-Although per-token offers a detailed level of granularity, it results in significant overhead during dequantization.
+Orthogonal to smoothing, we can mitigate the problem of outliers by refining the quantization granularity so that the number of affected elements by outliers becomes smaller.
+Although per-token quantization offers a detailed level of granularity, it results in significant overhead during dequantization.
 Specifically, each GPU thread in per-token quantization must handle multiple quantization scales, leading to a high latency of the dot product of the quantization scale vectors $\delta_{Q}$ and $\delta_{K}$.
@@
 However, quantizing $Q$ and $K$ to INT4 demands a finer quantization granularity.
-To address this, we propose per-thread quantization, a more precise and granular approach than the per-block quantizer, also without the additional overhead of vectors dot product between $\delta_{Q}$ and $\delta_{K}$.
+To address this, we propose per-thread quantization, a more precise and granular approach than the per-block quantizer, also without the additional overhead of the vector dot product between $\delta_{Q}$ and $\delta_{K}$.
 Specifically, each block of $Q$, i.e., $Q_{i}$, in SageAttention will be split into $c_{w}$ segments and processed by $c_{w}$ GPU warps in a GPU streaming processor (SM).
@@
 Then, each warp containing 32 threads uses the mma.m16n8k64 PTX instruction (NVIDIA, ) for the $Q_{w}K_{j}^{\top}$.
-According to the layout requirement of this instruction, we find that $Q_{w}[8k+i]$ could share one quantization scale, and $K_{j}[8k+2i]$ and $K_{j}[8k+2i+1]$ could share one quantization scale.
-Such a quantization method is more fine-grained and with no additional overhead.
+According to the layout requirement of this instruction, we find that $Q_{w}[8k+i]$ could share one quantization scale, and $K_{j}[8k+2i]$ together with $K_{j}[8k+2i+1]$ could share one quantization scale.
+Such a quantization method is more fine-grained with no additional overhead.
 This is because it assigns different GPU threads to distinct quantization groups based on the MMA instruction layout, with each thread performing dequantization only using a single quantization scale value.
@@
 Empirical results.
-As shown in Table 6 and Table 15, we compare the average and the worst accuracy of INT4 quantization at per-token, per-thread, per-block, and per-tensor granularity using real $Q,K,V$ across all layers of CogvideoX.
-Results indicate the accuracy of per-thread quantization is very close to per-token and outperforms a lot more than others.
+As shown in Table 6 and Table 15, we compare the average and worst accuracy of INT4 quantization at per-token, per-thread, per-block, and per-tensor granularity using real $Q,K,V$ across all layers of CogvideoX.
+Results indicate that the accuracy of per-thread quantization is very close to per-token and significantly outperforms other granularities.
+Moreover, Table 19 shows that per-thread quantization incurs almost no speed degradation, while per-token quantization introduces noticeable overhead due to the reduced hardware efficiency.
 Algorithm 1 Implementation of SageAttention2.
   Input: Matrices $Q(\text{FP16}),K(\text{FP16}),V(\text{FP16})\in\mathbb{R}^{N\times d}$, block size $b_{q},b_{kv}$, warp count $c_{w}$.
-  Preprocessing: $K=K-\mathrm{mean}(K)$, ${}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{V},\hat{V})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{V}}}(V)$.
+  Preprocessing: $K=K-\mathrm{mean}(K)$, ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{V},\hat{V})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{V}}}(V)$.
 //per-channel.
@@
   ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\bar{q}_{i}=\mathrm{mean}(Q_{i}),~~(\delta_{Q},\hat{Q}_{i})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{Q}(Q_{i}-\bar{q}_{i})}}$ //per-thread ;
-    for j in [1, $T_{n}$do
+  for j in [1, $T_{n}$] do
    ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{K},\hat{K}_{j})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{K}}}(K_{j})$ //per-thread,     $w=\text{range}(c_{w}),st=w*c_{w}$ ;
@@
 4.
-       $m_{ij}=\mathrm{max}(m_{i,j-1},\mathrm{rowmax}(S_{ij}))$, ${}\widetilde{P}_{ij}=\mathrm{exp}(S_{ij}-m_{ij})$, ${}l_{ij}=e^{m_{i,j-1}-m_{ij}}+\mathrm{rowsum}(\widetilde{P}_{ij})$ ;
+   $m_{ij}=\mathrm{max}(m_{i,j-1},\mathrm{rowmax}(S_{ij}))$, $\widetilde{P}_{ij}=\mathrm{exp}(S_{ij}-m_{ij})$, $l_{ij}=e^{m_{i,j-1}-m_{ij}}l_{i,j-1}+\mathrm{rowsum}(\widetilde{P}_{ij})$ ;
    $O_{ij}$(FP22) = $\mathrm{Matmul}((\widetilde{P}_{ij}*448)\text{.to(FP8.e4m3)},V_{j})$ ;
-       $O_{ij}$(FP32) = $\mathrm{diag}(e^{m_{i,j-1}-m_{ij}})^{-1}O_{i,j-1}(FP32)+O_{ij}$(FP22) ;
+   $O_{ij}$(FP32) = $\mathrm{diag}(e^{m_{i,j-1}-m_{ij}})O_{i,j-1}(FP32)+O_{ij}$(FP22) ;
   end for
-[…]
+  Load $\delta_{V}$ into an SM ;    $O_{i}=\mathrm{diag}(l_{i,T_{n}})^{-1}O_{i,T_{n}}\text{(FP32)}$ $448*\delta_{V}$ ;    Write $O_{i}$ ;
   end for
@@
 The distribution of $\tilde{P}$ is unique and differs from other activations.
-First, we note that $S_{ij}-m_{ij}\leq 0$, so $P_{ij}\in[0,1]$ ($\leq$ and $\in$ apply element-wisely).
+First, we note that $S_{ij}-m_{ij}\leq 0$, so $P_{ij}\in[0,1]$ ($\leq$ and $\in$ apply element-wise).
 We find that $\tilde{P}$ often consists of many small elements, but their sum is non-negligible (e.g., 5000 elements around $10^{-4}$).
 In this case, we must represent small elements accurately.
-INT quantization is unsuitable for this proposal since it puts the quantization points evenly within the numerical range.
-SageAttention (Zhang et al., 2025) choose to retain $\tilde{P}$ and $V$ in FP16, and accelerate the MM by decreasing the accumulator precision.
+INT quantization is unsuitable for this setting, since it distributes the quantization points evenly within the numerical range.
+SageAttention (Zhang et al., 2025c) choose to retain $\tilde{P}$ and $V$ in FP16, and accelerate the MM by decreasing the accumulator precision.
 However, this strategy is only effective on very few GPUs.
 We propose to quantize $\widetilde{P},V$ to FP8 with 4 exponent bits and 3 mantissa bits (E4M3).
-The numerical range of E4M3 is $[-448,+448]$.
-We quantize $P$ with a static scale: $\delta_{P}=\frac{1}{448}$ since the original $P$ elements is already in $[0,1]$.
+The numerical range of E4M3 is [448,+448][-448,+448].
+We quantize $P$ with a static scale: $\delta_{P}=\frac{1}{448}$ since the original $P$ elements are already in $[0,1]$.
 We quantize $V$ per-channel to address the channel-wise outliers shown in Fig.
-3.
-Empirical results in Table 7 and Table 16 show the average and the worst accuracy of different data types used for $\widetilde{P},V$ across all layers of CogvideoX.
+2.
+Empirical results in Table 7 and Table 16 show the average and worst accuracy of different data types used for $\widetilde{P},V$ across all layers of CogvideoX.
 The accumulator is always 32-bit.
@@
 Specifically, for mma(f32f8f8f32) instruction $C=AB+D$, where $A,B$ are FP8 matrices and $C,D$ are FP32 matrices, we initialize the $A,B$ to zero and vary $D$ to test the data type of the accumulator.
-When $D$ is initialized with 1 sign bit, 8 exponent bits, and 13 mantissa bits, the value of $C$ matches the result of the high-precision mma(f16f16f32) instruction.
-However, when $D$ is initialized with more than 13 mantissa bits, the error of $C$ corresponds to the difference between the results of mma(f32f16f16f32) and mma(f32f8f8f32).
-Consequently, the matrix multiplication of $\widetilde{P}V$, quantized to FP8, incurs a certain degree of accuracy loss compared to using an FP32 accumulator.
+When $D$ is initialized with 1 sign bit, 8 exponent bits, and 13 mantissa bits, the value of $C$ exactly matches $D$.
+However, when $D$ is initialized with more than 13 mantissa bits, the value of $C$ is equal to $D$ with its least significant 10 mantissa bits zeroed out (i.e., truncated).
+Consequently, matrix multiplication of $\widetilde{P}V$, quantized to FP8, incurs a certain degree of accuracy loss compared to using an FP32 accumulator.
 To mitigate this accuracy loss, we adopt a two-level accumulation strategy, which uses an FP32 buffer to accumulate the values of $\widetilde{P}_{ij}V_{j}$ in FP22.
@@
 Remark.
-The two-level accumulation strategy is also implemented in Cutlass (NVIDIA, 2023) and DeepSeekv3 (DeepSeek-AI et al., 2024) for computing activation-weight products in linear layers.
+The two-level accumulation strategy is also implemented in CUTLASS (NVIDIA, 2023) and DeepGemm (DeepSeek-AI et al., 2024) for computing weight-activation products in linear layers.
 To the best of our knowledge, we are the first to discover and investigate the effect of the FP22 accumulator and implement the two-level accumulation for attention.
@@
 We also figure out another way to mitigate the accuracy loss due to the FP22 accumulator when $V$ possesses channel-wise biases: $\overrightarrow{V_{m}}=\mathrm{mean}(V,\text{axis}=0),V=V-\overrightarrow{V_{m}}$.
-Additionally, to maintain the correctness of the attention computation, it is only necessary to add $\overrightarrow{V_{m}}$ to the final calculation of $O$: $O=O+\overrightarrow{V_{m}}$.
+Furthermore, to maintain the correctness of the attention computation, it is only necessary to add $\overrightarrow{V_{m}}$ to the final calculation of $O$: $O=O+\overrightarrow{V_{m}}$.
 This is because the sum of each row of the $\widetilde{P}$ matrix equals 1, so $\widetilde{P}\overrightarrow{V_{m}}=\overrightarrow{V_{m}}$.
@@
 This technique is optional and not employed in our main experiments, as it provides significant benefits only when $V$ exhibits channel-wise bias, which are absent in some models, such as Llama3.1 (see Fig.
-3).
+2).
 Figure 5: Speed comparison between SageAttention2 and baselines (RTX4090, headdim=128).
@@
      SageAttn2-8b     14.106     15.647     25.499     0.901
+Figure 6: Comparison examples from HunyuanVideo, prompts are sampled from open-sora prompt sets.
+Figure 7: A comparison example using SageAttn2-8b and FlashAttention3 on CogvideoX-1.5.
 4 Experiment
 Main result.
-SageAttention2 is faster than FlashAttention2 and xformers by approximately 3x and 4.5x, respectively.
-Moreover, SageAttention2 matches the speed of FlashAttention3(fp8) on the Hopper GPUs and is much more accurate than FlashAttention3(fp8).
-Our approach maintains end-to-end metrics across language, image, and video generation models.
+SageAttention2 is faster than FlashAttention2 and xformers by about 3x and 4.5x.
+Moreover, SageAttention2 matches the speed of FlashAttn3(fp8) on the Hopper GPUs and is much more accurate than FlashAttn3(fp8).
+SageAttention2 maintains end-to-end metrics across language, image, and video generation models.
 4.1 Setup
@@
 We validate the effectiveness of SageAttention2 across a diverse set of representative models from language, image, and video generation.
-Specifically, we conduct experiments on ten models: Llama2 (7B) (Touvron et al., 2023), Llama3.1 (8B) (Dubey et al., 2024), and GLM4 (9B) (GLM et al., 2024) for text2text, CogvideoX (2B), CogvideoX (1.5-5B) (Yang et al., 2025), HunyuanVideo (Kong et al., 2024), and Mochi (Team, 2024) for text2video, Flux (schnell) (Black Forest Labs, 2023) and Stable-Diffusion3.5 (turbo) (Stability AI, 2023) for text2image, and TIMM (Wightman, 2019) for image classification.
+Specifically, we conduct experiments on ten models: Llama2 (7B) (Touvron et al., 2023), Llama3.1 (8B) (Dubey et al., 2024), and GLM4 (9B) (GLM et al., 2024) for text2text, CogvideoX (2B), CogvideoX (1.5-5B) (Yang et al., 2025b), HunyuanVideo (Kong et al., 2024), and Mochi (Team, 2024) for text2video, Flux (schnell) (Black Forest Labs, 2023) and Stable-Diffusion3.5 (turbo) (Stability AI, 2023) for text2image, and TIMM (Wightman, 2019) for image classification.
 Datasets and metrics. For Details about the datasets and metrics we used, please refer to Appendix. A.7.
 Table 3: Two kernel implementations of SageAttention2.
-Kernel $\psi_{Q}(Q)$, $\psi_{K}(K)$ $\psi_{P}(\widetilde{P})$, $\psi_{V}(V)$
+[…]
 SageAttn2-4b INT4 per-thread FP8 per-block and per-channel
 SageAttn2-8b INT8 per-thread FP8 per-block and per-channel
-Implemetation. We implement two attention kernels as shown in Table 3 using CUDA.
+Implemetation.
+We implement two attention kernels as shown in Table 3 using CUDA.
+The 8-bit variant is adapted for NVIDIA Hopper GPUs, which lack native INT4 tensor core support, and incorporates all techniques described in Sec.
+3 except for smoothing Q.
 Baselines.
 (1) SmoothAttn.
-Following Qserve (Lin et al., 2024), we apply smooth quant for $Q,K$ with smoothing factor $\alpha=0.5$.
+Following Qserve (Lin et al., 2025), we apply smooth quant for $Q,K$ with smoothing factor $\alpha=0.5$.
 (2) HadmdAttn.
 Following Quarot (Ashkboos et al., 2024), we apply random Hadamard transformation for $Q,K$ before INT4 quantization.
-(3) SageAttention (Zhang et al., 2025), which uses smoothing $K$, INT8 per-block quantization for $Q,K$, and FP16 for $\widetilde{P},V$.
+(3) SageAttention (Zhang et al., 2025c), which uses smoothing $K$, INT8 per-block quantization for $Q,K$, and FP16 for $\widetilde{P},V$.
 (4) FlashAttn3(fp8), the FP8 version of FlashAttention3, which only runs on Hopper GPUs.
@@
 Full-Precision - 81.5% 6.013 75.360
-INT4 Quantization 72.6% 11.698 24.670
+INT4
+Quantization
+72.6% 11.698 24.670
 80.8% 6.219 75.147
@@
 Table 8: End-to-end generation latency using SageAttention2 (The latency of Llama3.1 is the time to first token generation using different sequence lengths).
-Model GPU Original SageAttn 2-8b SageAttn 2-4b
+Model GPU Original
+SageAttn
+2-8b
+SageAttn
+2-4b
 CogvideoX (2B) RTX4090 86 s 54 s 52 s
@@
 Kernel Speed.
-We compare the Speed of SageAttention2 against baselines using configurations with headdim=64 and headdim=128, both with and without Causal Mask (Vaswani, 2017).
+We compare the speed of SageAttention2 against baselines using headdim=64 and headdim=128, both with and without Causal Mask (Vaswani, 2017).
+Detailed setup can be found in Appendix A.8.
 Specifically, Fig.
@@
 Fig.
-91011121314, and 15 in Appendix A.2 show more kernel speed comparison on RTX4090, L20, H20, H100 GPUs.
+101112131415, and 16 in Appendix A.2 show more speed results on RTX4090, L20, H20, H100 GPUs.
 Accuracy.
@@
 Visible image and video examples.
-Fig.6 and Fig.
-8 in Appendix A.1 show some visible comparision examples from HunyuanVideo, Mochi and CogvideoX (1.5-5B).
+Fig.
+678, and 9 show some visible comparison examples from HunyuanVideo, Mochi and CogvideoX.
 We can observe that SageAttn2-8b does not introduce any visible differences compared to full-precision attention, whereas SageAttn2-4b has minor differences but is much better than the baselines.
@@
 We introduce SageAttention2, an efficient and accurate quantized attention.
-First, we propose to quantize matrixes $(Q,K)$ in a thread-level granularity and $(\widetilde{P},V)$ to FP8.
+First, we propose to quantize matrices $(Q,K)$ in a thread-level granularity and $(\widetilde{P},V)$ to FP8.
 Second, we propose a method to smooth $Q$, enhancing the accuracy of $QK^{\top}$.
@@
 Moreover, SageAttention2 matches the speed of FlashAttention3(fp8) on the Hopper GPUs, but offers significantly higher accuracy.
-Extensive testing confirms that our approach maintains end-to-end metrics across language, image, and video generation models.
+Extensive experiments confirm that our approach maintains end-to-end metrics across language, image, and video generation models.
+Acknowledgment
+This work was supported by the NSFC Projects (Nos. 92270001, 62376131). J.Z is also supported by the XPlorer Prize.
+Impact Statement
+This paper presents work that aims to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
 Appendix A Appendix
 A.1 Visible Comparison Exmaples
-Figure 6: Comparison examples from HunyuanVideo, prompts are sampled from open-sora prompt sets.
 Figure 8: Comparison examples from CogvideoX (2B), prompts are sampled from open-sora prompt sets.
-Figure 8: A comparison example using SageAttn2-8b and FlashAttention3 on CogvideoX-1.5, Mochi and HunyuanVideo.
+Figure 9: A comparison example using SageAttn2-8b and FlashAttention3 on Mochi and HunyuanVideo.
 A.2 Additional Kernel Speed Comparison
 Fig.
-9,  10,  11,  12,  13,  14, and  15 compare the speed of SageAttention2 against baselines using configurations with headdim=64 and headdim=128, both with and without Causal Mask (Vaswani, 2017), on RTX4090, L20, H100, and H20 GPUs.
+10,  11,  12,  13,  14,  15, and  16 compare the speed of SageAttention2 against baselines using headdim=64 and headdim=128, both with and without Causal Mask (Vaswani, 2017), on RTX4090, L20, H100, and H20 GPUs.
 Table 9 summarizes the performance gain of different attention methods against baselines on various modern GPUs.
@@
 As shown in Fig.
-16, this strategy could enhance the accuracy of FP22 for values in $\widetilde{P}V$ for the following reasons: Each row of $\widetilde{P}$ spans a value range from 0 to 1, and each column of $V$ in some models consistently features channel-wise biases that are exclusively positive or negative, for instance, ranging between 8 and 9 in CogvideoX.
+17, this strategy could enhance the accuracy of FP22 for values in $\widetilde{P}V$ for the following reasons: Each row of $\widetilde{P}$ spans a value range from 0 to 1, and each column of $V$ in some models consistently features channel-wise biases that are exclusively positive or negative, for instance, ranging between 8 and 9 in CogvideoX.
 Consequently, the values of $\widetilde{P}V$ could be quite large.
 However, the floating-point number representation range is not uniform—it is denser near zero.
-Therefore, by subtracting the mean $\overrightarrow{V_{m}}$ along the channel dimension from $V$, the values of $\widetilde{P}V$ will be closer to zero, resulting in higher representational precision (see Fig.
+Therefore, by subtracting the mean $\overrightarrow{V_{m}}$ along the channel dimension from $V$, the values of $\widetilde{P}V$ will be closer to zero, resulting in a higher representational precision (see Fig.
 17 for a visual demonstration).
@@
 This makes the representation of FP22 more accurate.
-Meanwhile, $\overrightarrow{V_{m}}$ is retained in FP16 and added to $O$ at the end without causing a loss of precision for the $\overrightarrow{V_{m}}$ part.
+Meanwhile, $\overrightarrow{V_{m}}$ is retained in FP16 and is added to $O$ at the end without causing a loss of precision for the $\overrightarrow{V_{m}}$ part.
 A.4 Experiment of Smoothing V
@@
 It demonstrates that smoothing $V$ could improve the accuracy of SageAttention2 when quantizing $Q,K$ to INT4 and $\widetilde{P},V$ to FP8.
-A.5 Per-Thread Quantization Formulation
-Figure 17: Memory layout of INT4/INT8 tensor core for accumulator matrix $C$ and $D$ in $D=A*B+C$ among 32 threads (T0 $\sim$ T31) in a warp.
+We find that smoothing V is generally effective for diffusion models (Zheng et al., 2023, 2024b, 2024a, 2025; Zhao et al., 2024, 2025a; Wang et al., 2024).
+A.5 Theoretical Analysis of Smoothing
+In this section, we analyze the benefit of smoothing from a theoretical perspective.
+Let $X\in\mathbb{R}^{N\times d}$ be $N$ activation tokens of dimension $d$.
+Following (Dettmers et al., 2023), we suppose that an activation token follows an Gaussian distribution $\mathcal{N}(\boldsymbol{\mu},\Sigma^{2})$, where $\boldsymbol{\mu}=(\mu_{1},\mu_{2},\ldots,\mu_{d})$ and $\Sigma^{2}$ is a diagonal matrix with $\Sigma^{2}=\mathrm{diag}(\sigma_{1}^{2},\sigma_{2}^{2},\ldots,\sigma_{d}^{2})$.
+Further, we suppose that different token $X_{i}$ is i.i.d.
+sampled from the same distribution.
+Suppose the absolute maximum value in a quantization group (Hu et al., 2025; Zhang et al., 2025j) is $M$, and the bit width is $b$, then there are $2^{b}$ quantization levels.
+Under the round-to-nearest strategy, the expected quantization error is $\frac{1}{2}\cdot\frac{2M}{2^{b}}$, which is proportional to the maximum absolute value in the quantization group.
+So a smaller absolute maximum value leads to smaller quantization error.
+After smoothing, we have:
+$Y_{ij}=X_{ij}-\frac{1}{N}\sum_{k=1}^{N}X_{kj}$ (3)
+and $Y_{ij}$ also follows a Gaussian distribution. The mean and variance of $Y_{ij}$ can be calculated as follows:
+$\mathbb{E}[Y_{ij}]$ $=\mathbb{E}[X_{ij}]-\frac{1}{N}\sum_{k=1}^{N}\mathbb{E}[X_{kj}]=\mu_{j}-\frac{1}{N}\sum_{k=1}^{N}\mu_{j}=0$ (4)
+$\mathrm{Var}[Y_{ij}]$ $=\mathrm{Var}[\frac{N-1}{N}X_{ij}]+\sum_{k=1,k\neq i}^{N}\mathrm{Var}[\frac{1}{N}X_{kj}]=\frac{(N-1)^{2}}{N^{2}}\sigma_{j}^{2}+(N-1)\frac{1}{N^{2}}\sigma_{j}^{2}=\frac{(N-1)}{N}\sigma_{j}^{2}$ (5)
+So $Y_{ij}$ has a smaller mean and variance compared to $X_{ij}$. Following the property of Gaussian distribution, we have:
+$P(|Y_{ij}|<\epsilon)>P(|X_{ij}|>\epsilon),\ \forall\epsilon>0$ (6)
+So the distribution of $X_{ij}$ is more concentrated toward 0 after smoothing. Then we know that
+$P(\mathrm{absmax}(Y_{i})<\epsilon)=\prod_{j=1}^{d}P(|Y_{ij}|<\epsilon)>\prod_{j=1}^{d}P(|X_{ij}|>\epsilon)=P(\mathrm{absmax}(X_{i})<\epsilon)\$(7)
+So this makes the distribution of absolute max value in a token more concentrated towards 0, leading to smaller quantization error.
+A.6 Per-Thread Quantization Formulation
+Figure 18: Memory layout of INT4/INT8 tensor core for accumulator matrix $C$ and $D$ in D=AB+CD=A*B+C among 32 threads (T0 $\sim$ T31) in a warp.
 $C$ and $D$ is of shape 16x8.
@@
 Their usage significantly enhances computational efficiency and performance in AI and high-performance computing (HPC) workloads.
-Tensor cores compute small tiles of MMA operations, specifically $D=A*B+C$ on a warp (32 contiguous threads) basis.
+Tensor cores compute small tiles of MMA operations, specifically D=AB+CD=A*B+C on a warp (32 contiguous threads) basis.
 Each thread in the warp holds a fragment of input matrices and will get a fragment of output matrix as a computation result.
@@
 $\hat{K}[k_{n}[i_{\delta k}]]=$ $\left\lceil\frac{K[k_{n}[i_{\delta k}]]}{\delta_{K}[i_{\delta k}]}\right\rfloor$
-In Eq. A.5, $c_{w}$ is the count of GPU Warps. $b_{q}$ and $b_{k}$ are the block size of $Q,K$. $n$ is the token index of $Q,K$.
 By ensuring results held by each thread share a common dequantization scale (belong to the same quantization group), we can avoid the overhead associated with per-token quantization.
-Leveraging this observation, we design per-thread quantization, as shown in Fig.
+Leveraging this observation, we design per-thread quantization as formulated in Eq.
+A.6, where $c_{w}$ is the count of GPU Warps, $b_{q}$ and $b_{k}$ are the block size of $Q,K$, and $n$ is the token index of $Q,K$.
+For typical block size of $b_{q}=128$, $b_{k}=64$ and warp number $c_{w}=4$ (as used in FlashAttention2), each warp processes a tile of 32 query tokens and 64 key tokens.
+Query tokens $i,8+i,16+i,24+i$ ($i=0,1,\cdots,7$) can be made into one quantization group and key tokens $j,1+j,8+j,9+j,\cdots,56+j,57+j$ ($j=0,1,2,3$) can be made into one quantization group, as visualized in Fig.
 4.
-For typical block size of $b_{q}=128$, $b_{k}=64$ and warp number $c_{w}=4$ (as used in FlashAttention2), each warp processes a tile of 32 query tokens and 64 key tokens.
-Query tokens $i,8+i,16+i,24+i$ ($i=0,1,\cdots,7$) can be made into one quantization group and key tokens $j,1+j,8+j,9+j,\cdots,56+j,57+j$ ($j=0,1,2,3$) can be made into one quantization group.
 This design aligns with the memory layout of output matrix $D$ of tensor core shown in Fig.
-17, ensuring that each thread only needs one $Q$ scale and one $K$ scale for dequantization.
-As a result, this approach creates 32 quantization groups for $Q$ (8 for each of the 4 warps) and 4 quantization groups for $K$ in a 128x64 block, providing 32×32$\times$32 × and 4×4$\times$4 × finer granularity compared to per-block quantization for query tokens and key tokens, respectively.
+18, ensuring that each thread only needs one $Q$ scale and one $K$ scale for dequantization.
+As a result, this approach creates 32 quantization groups for $Q$ (8 for each of the 4 warps) and 4 quantization groups for $K$ in a 128x64 block, providing $32\times$ and $4\times$ finer granularity compared to per-block quantization for query tokens and key tokens, respectively.
 Table 6 and Table 15 show the accuracy gains by using per-thread quantization.
-Per-thread quantization achieves accuracy that closely matches per-token quantization, without introducing any kernel speed degradation (see Fig.
-18).
-A.6 Datasets and Metrics in Experiments
+Per-thread quantization achieves accuracy that closely matches per-token quantization, without introducing any kernel speed degradation (see Table 19 and 19).
+A.7 Datasets and Metrics in Experiments
 Datasets.
-Text-to-text models are evaluated on four zero-shot tasks: WikiText (Merity et al., 2022) to assess the model’s prediction confidence, LAMBADA (Paperno et al., 2016) evaluate contextual understanding, MMLU (Hendrycks et al., 2020) for measuring knowledge across various subjects, and Longbench (Bai et al., 2024) for comprehensive assessment of long context understanding capabilities.
-Text-to-video models are evaluated using the open-sora (Zheng et al., 2024) prompt sets.
+Text-to-text models are evaluated on four zero-shot tasks: WikiText (Merity et al., 2022) to assess the model’s prediction confidence, LAMBADA (Paperno et al., 2016) evaluate contextual understanding, MMLU (Hendrycks et al., 2021b) for measuring knowledge across various subjects, and Longbench (Bai et al., 2024) for comprehensive assessment of long context understanding capabilities.
+Text-to-video models are evaluated using the open-sora (Zheng et al., 2024c) prompt sets.
 Text-to-image models are assessed on MJHQ-30K (Li et al., 2024).
-TIMM is evaluated on on three image datasets: ImageNet (Deng et al., 2009), ImageNet-Sketch (Sketch) (Wang et al., 2019), and ImageNet-Rendition (ImageNet-r) (Hendrycks et al., 2021).
+TIMM is evaluated on on three image datasets: ImageNet (Deng et al., 2009), ImageNet-Sketch (Sketch) (Wang et al., 2019), and ImageNet-Rendition (ImageNet-r) (Hendrycks et al., 2021a).
 End-to-end metrics.
@@
 For text-to-video models, following Zhao et al.
-(2025), we evaluate the quality of generated videos on five metrics: CLIPSIM and CLIP-Temp (CLIP-T) (Liu et al., 2024) to measure the text-video alignment; (VQA-a) and (VQA-t) to assess the video aesthetic and technical quality, respectively; and Flow-score (FScore) for temporal consistency (Wu et al., 2023).
+(2025b), we evaluate the quality of generated videos on five metrics: CLIPSIM and CLIP-Temp (CLIP-T) (Liu et al., 2024) to measure the text-video alignment; (VQA-a) and (VQA-t) to assess the video aesthetic and technical quality, respectively; and Flow-score (FScore) for temporal consistency (Wu et al., 2023).
 For text-to-image models, generated images are compared with the images in MJHQ-30K dataset in three aspects: FID (Heusel et al., 2017) and sFID (Salimans et al., 2016) for fidelity evaluation, Clipscore (CLIP) (Hessel et al., 2021) for text-image alignment, and ImageReward (IR) (Xu et al., 2023) for human preference.
@@
 Then, Cosine similarity: $CosSim=\sum OO^{\prime}/\smash{\sqrt{\sum O^{2}}}\smash{\sqrt{\sum O^{\prime 2}}}$, Relative L1 distance: $L1=\sum|O-O^{\prime}|/\sum|O|$, Root mean square error: $RMSE=\sqrt{(1/n)\sum(O-O^{\prime})^{2}}$.
+A.8 Kernel Benchmark Setup
+We benchmark kernel speed with a batch size of 4 and 32 attention heads across a variety of sequence lengths.
+Benchmarks are conducted using head dimensions of 64 and 128, both with and without Causal Mask (Vaswani, 2017).
+To generate input tensors for benchmarking, we follow standard practices adopted in prior works such as FlashAttention (Dao et al., 2022).
+For floating-point data types, inputs are drawn from a Gaussian distribution with mean 0 and standard deviation 1, while for integer data types, inputs are uniformly sampled within the representation range:[-128, 127] for INT8 and [-8, 7] for INT4.
 Table 11: End-to-end metrics on Llama2 (7B).
@@
 SmoothAttn 64.49% 0.9262 0.7204
-Smooth K (Ours) 90.86% 0.3565 0.1464
-Smooth Q (Ours) 93.10% 0.2989 0.2195
+Smooth K 90.86% 0.3565 0.1464
+Smooth Q 93.10% 0.2989 0.2195
 SageAttn2-4b 96.71% 0.1956 0.0779
@@
    + Smoothing Q    273
-A.7 Additional Experiments and Analysis
+Table 19: Comparison of different quantization granularities measured on L20 GPU, with $QK^{\top}$ in INT4 and $\widetilde{P}V$ in FP8.
+   Granularity    TOPS
+   Per-tensor    286
+   Per-block    284
+   Per-thread    283
+   Per-token    268
+A.9 Additional Experiments and Analysis
 Additional Results. Table 11,  12 and  13 show results of SageAttention2 and other baselines on Llama2 (7B), CogvideoX (2B) and TIMM.
@@
 We further conduct experiments on super-long context using Llama-3-262k (8B) [^1] https://huggingface.co/gradientai/Llama-3-8B-Instruct-262k on InfiniBench (Zhang et al., 2024) and Needle-in-a-Haystack (NIAH) (Kamradt, 2023), with sequence lengths reaching up to 262k tokens on an H100 GPU.
+Since Hopper GPUs lack native INT4 tensor core support, we use SageAttention2-8b for this evaluation.
+We compare it against FlashAttention3(fp8), ensuring both methods operate under the same bit width.
 Results are shown in Table 14 and Fig 19.
 SageAttention2 maintains model performance even under super-long context, while FlashAttention3(fp8) suffers from end-to-end accuracy degradation.
+Results of Audio Tasks.
+We evaluate Qwen2-Audio (7b) (Chu et al., 2024), a speech-to-text model, on the ASR task using the Librispeech (Panayotov et al., 2015) test split and measured its performance with the WER metric (Word Error Rate).
+As shown in Table 20, SageAttention2 consistently outperforms the baselines, highlighting its effectiveness in audio-related models and benchmarks.
+Table 20: End-to-end metrics on Qwen2-Audio (7B).
+Model Attention Test-Clean $\downarrow$ Test-Dev $\downarrow$
+Qwen2-Audio Full-Precision 1.74 4.01
+HadmdAttn 1.77 4.05
+SmoothAttn 1.76 4.01
+SageAttention 1.74 4.02
+SageAttn2-4b 1.73 3.99
+SageAttn2-8b 1.72 4.03